### v1

# A Practitioner’s Guide to Kolmogorov–Arnold Networks

Abstract
Kolmogorov-Arnold Networks (KANs) have recently emerged as a promising alternative to traditional Multilayer Perceptrons (MLPs), inspired by the Kolmogorov-Arnold representation theorem. Unlike MLPs, which use fixed activation functions on nodes, KANs employ learnable univariate basis functions on edges, offering enhanced expressivity and interpretability. This review provides a systematic and comprehensive overview of the rapidly expanding KAN landscape, moving beyond simple performance comparisons to offer a structured synthesis of theoretical foundations, architectural variants, and practical implementation strategies. By collecting and categorizing a vast array of open-source implementations, we map the vibrant ecosystem supporting KAN development. We begin by bridging the conceptual gap between KANs and MLPs, establishing their formal equivalence and highlighting the superior parameter efficiency of the KAN formulation. A central theme of our review is the critical role of the basis function; we survey a wide array of choices—including B-splines, Chebyshev and Jacobi polynomials, ReLU compositions, Gaussian RBFs, and Fourier series—and analyze their respective trade-offs in terms of smoothness, locality, and computational cost. We then categorize recent advancements into a clear roadmap, covering techniques for improving accuracy, efficiency, and regularization. Key topics include physics-informed loss design, adaptive sampling, domain decomposition, hybrid architectures, and specialized methods for handling discontinuities. Finally, we provide a practical “Choose-Your-KAN” guide to help practitioners select appropriate architectures, and we conclude by identifying current research gaps. We argue for a shift away from simplistic KAN-vs-MLP benchmarks toward a more methodical, basis-centric exploration of this rich architectural paradigm. The associated GitHub repository (https://github.com/AmirNoori68/kan-review) complements this paper and serves as a structured reference for ongoing KAN research.
1 Introduction
Multilayer perceptrons are universal approximators and remain a standard building block for regression, function approximation, and pattern recognition across scientific and engineering applications. A major extension of this paradigm is the physics-informed neural network (PINN) framework of Raissi et al. [1], which has rapidly grown and enabled large-scale studies.
Since then, numerous advances have broadened the scope of MLP-based PINNs. Fractional operators were addressed by fPINNs [2], while uncertainty quantification for forward and inverse problems was introduced in [3]. Extensions such as VPINNs [4], XPINNs for domain decomposition [5], and multi-fidelity PINNs [6] improved accuracy and scalability. The framework also inspired convergence analyses [7], NTK-based diagnostics [8], and practical toolkits like DeepXDE [9]. More recent contributions include multi-stage training with near machine-precision accuracy [10], adaptive residual weighting [11], and distributed solvers for extreme-scale PDEs [12, 13, 14, 15]. Collectively, these developments establish MLP–PINNs as a mature reference standard in scientific machine learning, many of which, as we show in Section 5, have directly inspired KAN development.
Despite their flexibility, MLPs face well-documented limitations. Fixed activation functions restrict adaptability [16]. Network behavior can be difficult to interpret [17] and to attribute causally [18]. Achieving high accuracy often entails large parameter counts, which can hinder efficient updates [19, 20]. Generalization robustness can degrade in challenging regimes [21, 22]. Optimization itself can be fragile or stiff, depending on the task and scaling [23]. MLPs also exhibit spectral bias: a tendency to learn low frequencies faster than high frequencies [24, 25]. The effect slows convergence and reduces accuracy for oscillatory or sharp-gradient solutions [26]. This motivates investigations into alternatives with explicit basis control.
Kolmogorov–Arnold Networks [27] introduce a distinct parameterization aimed at overcoming several limitations of MLPs. Instead of relying on fixed nonlinearities, KANs place learnable univariate functions on edges, inspired by the Kolmogorov representation theorem, which decomposes multivariate mappings into sums of one-dimensional transforms. The original implementation employs B-spline bases [27], but the framework is inherently modular: Chebyshev polynomials, Gaussian kernels, and other bases can be substituted to suit the problem structure. We regard this flexibility as an asset rather than a complication, since the choice of basis family directly governs smoothness, locality, and spectral behavior—thereby shaping both expressivity and interpretability. Consequently, the combination of adaptive basis learning with the freedom to select different basis families represent a major advantage of the KAN approach. Beyond basis design, KANs encompass a broad landscape of architectural extensions, optimization strategies, and theoretical developments—an overview of which is outlined in Table 1 to guide the reader through the remainder of this paper.
Table 1: Paper roadmap (clickable links to sections).
1 Introduction 2 MLP Basics 3 KAT and KAN 4 Bridging KANs and MLPs 5 How KANs Extend the MLPs Landscape 6 Basis Functions $\triangleright$ 6.1 B-spline $\triangleright$ 6.2 Chebyshev Polynomial $\triangleright$ 6.3 ReLU $\triangleright$ 6.4 Jacobi Polynomials $\triangleright$ 6.5 Gaussian RBF $\triangleright$ 6.6 Fourier $\triangleright$ 6.7 Wavelet $\triangleright$ 6.8 Finite-Basis $\triangleright$ 6.9 SincKAN 7 Accuracy Improvement $\triangleright$ 7.1 Physics Constraints & Loss Design $\triangleright$ 7.2 Adaptive Sampling & Grids $\triangleright$ 7.3 Domain Decomposition $\triangleright$ 7.4 Function Decomposition $\triangleright$ 7.5 Hybrid/Ensemble & Data $\triangleright$ 7.6 Sequence/Attention Hybrids $\triangleright$ 7.7 Discontinuities & Sharp Gradients $\triangleright$ 7.8 Optimization & Adaptive Training 8 Efficiency Improvement $\triangleright$ 8.1 Parallelism, GPU, and JAX Engineering $\triangleright$ 8.2 Matrix Optimization & Efficient Bases 9 Sparsity & Regularization 10 Scaling Laws & Convergence $\triangleright$ 10.1 Theoretical Approximation Rates $\triangleright$ 10.2 Spectral Bias $\triangleright$ 10.3 NTK-Based Convergence $\triangleright$ 10.4 Practical Trade-offs 11 Practical “Choose–Your–KAN” Guide 12 Current Gaps and Path Forward 13 Conclusion
When it comes to deciding whether MLP-PINNs or KAN-PIKANs are the better choice, the comparison is far from straightforward, with studies that emphasize fairness often reaching divergent or contradictory conclusions [28] [^1] https://github.com/yu-rp/KANbeFair, [29, 30]. A key reason is that KANs are not a monolithic architecture; their performance is critically influenced by the choice of basis function. For example, a comprehensive comparison by Farea and Celebi [31] [^2] https://github.com/afrah/pinn_learnable_activation found that the optimal basis (e.g., B-spline, Fourier, Gaussian) varies significantly with the PDE being solved, making a simple “KAN vs. MLP” verdict insufficient.
Despite these complexities, Table 2 summarizes reported outcomes across regression, function approximation, and PDE-solving tasks. The entries reflect prevailing tendencies in the literature rather than definitive results for every case. A consistent pattern is that KANs—particularly when equipped with specialized basis functions—tend to match or outperform vanilla MLPs and PINNs in terms of accuracy and convergence speed.
Table 2: Representative performance trends comparing KAN-based models with MLPs. Results are aggregated from the literature and reflect general outcomes. “Slower training” refers to higher per-iteration cost, though total wall-clock time can be different due to fewer epochs to convergence. All rows compare KANs to plain MLP/PINN baselines without advanced features.
Ref. Accuracy Convergence/Time(per-iter.) Basis Functions
Regression & Symbolic Representation
[32] BSRBF-KAN $\approx$ MLP Faster convergence; slower training B-spline + Gaussian
[33] MLP-KAN >> MLP B-Spline
[34] KAN << MLP Slower training; less generalizable B-Spline
[35] DE-KAN >> MLP Faster convergence B-spline
[36] KAN-ODE >> MLP-ODEs Faster convergence; slower training Gaussian
[37] KAN-Therm >> MLP Faster convergence B-spline
Function Approximation
[27] KAN $\geq$ MLP Faster convergence; slower training B-spline
[28] KAN >> MLP (symbolic) Slower training B-spline
[38] KAN $\approx$ MLP Faster training Free-knot B-spline
[39] KAN $\geq$ MLP (noisy) Faster convergence; slower training B-spline
[39] KAN $\leq$ MLP (non-smooth) Faster convergence; slower training B-spline
[40] PowerMLP >> MLP Slower training Power-ReLU
[41] SincKAN >> MLP Faster convergence; slower training Sinc
[42] ChebyKAN >> MLP Faster convergence; slower training Shifted Chebyshev
[43] QKAN >> MLP Faster convergence Quantum variational
PDE Solving
[44] KKAN >> PINN Faster convergence Various Basis
[45] PIKAN $\gg$ PINN Faster convergence; slower training B-spline
[…]
[46] DeepOKAN >> DeepONet Faster convergence; slower training Gaussian
[47] PIKAN >> PINN Faster convergence B-Spline
[48] KAN-MHA >> PINN Faster convergence; comparable time B-spline + Attention
[49] Res-KAN $\gg$ PINN Faster convergence; better generalization B-Spline + Residual
[50] HPKM-PINN >> PINN Faster convergence; slower training B-spline
[51] PI-KAN $\gg$ PINN Faster convergence Spline
[52] DPINN $\gg$ PINN Faster convergence; slower training B-spline + Fourier
[53] PIKAN $\approx$ PINN Slower training B-spline
[29] PIKAN $\approx$ PINN slower training Various Basis
[54] EPi-cKAN $\gg$ PINN Slower training; better generalization Chebyshev
[55] Scaled-cPIKAN $\gg$ PINN Faster convergence Chebyshev
[56] PIKAN $\gg$ PINN Faster convergence Chebyshev
[57] tanh-PIKAN >> PINN Chebyshev
[58] AL-PKAN >> PINN Faster convergence B-spline decoder
[59] KANtrol >> PINN Slower training B-spline
[60] PIKAN >> PINN Slower training B-spline
[61] KINN >> PINNs Slower training B-spline
[62] PIKAN $\approx$ PINNs Slower training B-spline
[63] PIKAN $\approx$ PINNs Slower training Fourier-based
[64] MR-PIKAN $\gg$ PINN Slower training Chebyshev
[…]
[65] J-PIKAN $\gg$ PINN Faster convergence; slower training Jacobi (orthogonal)
[66] Legend-KINN >> MLP, KAN faster convergence; slower training Legendre
A number of recent surveys have appeared on KANs, each highlighting different facets of this emerging architecture. Notable contributions include those by Andrade et al. [67], Basina et al. [68], Beatrize et al. [69], Faroughi et al. [70], Kilani et al. [71], Hou et al. [72], Dutta et al. [73], Essahraui et al. [74], and Somvanshi et al. [75], which together provide valuable entry points into the literature. Our review builds on these efforts by seeking to offer a more systematic and comprehensive perspective: rather than cataloguing studies in isolation, we integrate theoretical, architectural, optimization, and application viewpoints into a structured roadmap. By combining comparative analysis, methodological insights, and practical guidance (Sections 611), our goal is to provide a resource that complements existing surveys and serves readers aiming to both understand and apply KANs.
In parallel, KANs are rapidly expanding on open-access platforms, with diverse implementations and variants shared by the community. Curated repositories such as [76] [^3] https://github.com/mintisan/awesome-kan track the latest developments across different fields, making this growth more accessible to readers. Table 3 summarizes representative GitHub repositories that illustrate this fast-growing ecosystem in regression, function approximation, and PDE solving. These resources are also referenced in footnotes throughout this study when further details of specific methods are discussed.
Table 3: GitHub repositories related to KANs for regression, function approximation, or PDEs.
Repository (https://github.com/…) Description Ref.
KindXiaoming/pykan Official PyKAN for “KAN” and “KAN 2.0”. [77]
afrah/pinn_learnable_activation Compares various KAN bases vs. MLP on PDEs [31]
1ssb/torchkan Simplified PyTorch KAN with variants [78]
mintisan/awesome-kan Curated list of KAN resources, projects, and papers. [76]
sidhu2690/Deep-KAN Spline-KAN examples and a PyPI package. [79]
sidhu2690/RBF-KAN RBF-KAN examples [80]
yu-rp/KANbeFair Fair benchmarking of KANs vs MLPs. [28]
Blealtan/efficient-kan Efficient PyTorch implementation of KAN. [81]
srigas/jaxKAN JAX-based KAN package with grid extension support. [82]
ZiyaoLi/fast-kan FastKAN using RBFs. [83]
AthanasiosDelis/faster-kan Using SWitch Activation Function [84]
Indoxer/LKAN Implementations of KAN variations. [85]
pnnl/neuromancer (fbkans branch) Parametric constrained optimization. [86]
quiqi/relu_kan Minimal ReLU-KAN. [87]
OSU-STARLAB/MatrixKAN Matrix-parallelized KAN. [88]
Iri-sated/PowerMLP MLP-type network with KAN-level expressiveness. [40]
GistNoesis/FourierKAN Fourier-based KAN layer. [89]
GistNoesis/FusedFourierKAN Optimized FourierKAN with fused GPU kernels [90]
alirezaafzalaghaei/fKAN Fractional KAN using Jacobi functions. [91]
alirezaafzalaghaei/rKAN Rational KAN (Padé/Jacobi rational designs). [92]
M-Wolff/CVKAN Complex-valued KANs. [93]
DUCH714/SincKAN Sinc-based KAN with PINN applications. [41]
SynodicMonth/ChebyKAN Chebyshev polynomial-based KAN variant. [94]
Boris-73-TA/OrthogPolyKANs Orthogonal polynomial-based KAN implementations. [95]
kolmogorovArnoldFourierNetwork/kaf_act PyTorch activation combining with RFF. [96]
kolmogorovArnoldFourierNetwork/KAF Kolmogorov-Arnold Fourier Networks. [63]
kelvinhkcs/HRKAN Higher-order ReLU-KANs. [97]
yizheng-wang/KINN PIKAN for solid mechanics PDEs. [61]
Ali-Stanford/KAN_PointNet_CFD Jacobi-based network for CFD predictions. [62]
Jinfeng-Xu/FKAN-GCF FourierKAN-GCF for graph filtering. [98]
jdtoscano94/KKANs_PIML Kurkova-KANs combining MLP with basis functions. [44]
Zhangyanbo/MLP-KAN MLP-augmented KAN activations. [99]
Adamdad/kat Kolmogorov-Arnold Transformer. [100]
YihongDong/FAN Fourier Analysis Network (FAN). [101]
seydi1370/Basis_Functions Polynomial bases for KANs (comparative study). [102]
zavareh1/Wav-KAN Wav-KAN: wavelet-based KANs. [103]
Jim137/qkan Quantum variational activations and pruning tools. [43]
liouvill/KAN-Converge Additive & hybrid KANs for convergence-rate experiments [104]
hoangthangta/BSRBF_KAN Combines B-splines (BS) and radial basis functions (RBF). [32]
wmdataphys/Bayesian-HR-KAN Introduces Bayesian higher-order ReLU-KANs. [105]
zhang-zhuo001/Legend-KINN Legendre polynomial–based KAN. [66]
DiabAbu/DeepOKAN Deep Operator Network based on KAN. [46]
DENG-MIT/LeanKAN A memory-efficient Kolmogorov–Arnold Network. [106]
Organization (Table 1). Section 2 reviews MLP basics. Section 3 introduces the Kolmogorov–Arnold theorem (KAT) and formalizes KAN layers. Section 4 connects KANs to classical neural networks. Section 6 surveys basis families and computational trade-offs. Section 7 covers accuracy-improvement strategies, and Section 8 addresses efficiency via parallelism/GPU/JAX and matrix optimization. Section 9 treats sparsity, regularization, and Bayesian variants, while Section 10 discusses scaling laws and convergence. Section 11 presents the practical “Choose–Your–KAN” guide. We conclude with final remarks in Section 13.
2 MLP Basics
MLPs are feedforward neural networks composed of stacked affine layers followed by element-wise nonlinear activations. Given an input vector $\mathbf{x}\in[0,1]^{n}$, an MLP with $L$ layers produces an output $\hat{f}(\mathbf{x};\bm{\theta})\in\mathbb{R}$ via
$\mathbf{z}^{(0)}$ $=\mathbf{x},$ (1)
$\mathbf{z}^{(\ell)}$ $=\sigma\!\left(\mathbf{W}^{(\ell)}\mathbf{z}^{(\ell-1)}+\mathbf{b}^{(\ell)}\right),\quad\ell=1,\dots,L-1,$ (2)
$\hat{f}(\mathbf{x};\bm{\theta})$ $=\mathbf{W}^{(L)}\mathbf{z}^{(L-1)}+\mathbf{b}^{(L)},$ (3)
where
  • $\sigma(\cdot)$ is a nonlinear activation function (e.g., ReLU, $\tanh$),
  • $\mathbf{W}^{(\ell)}\in\mathbb{R}^{n_{\ell}\times n_{\ell-1}}$ and $\mathbf{b}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ are the weights and biases at layer $\ell$,
  • $\mathbf{z}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ is the hidden representation at layer $\ell$,
  • $n_{\ell}$ is the width (number of neurons) of layer $\ell$,
  • $\bm{\theta}=\{\mathbf{W}^{(\ell)},\mathbf{b}^{(\ell)}\}_{\ell=1}^{L}$ collects all trainable parameters.
3 KAT and KAN
Kolmogorov–Arnold Theorem (KAT). A landmark result by Kolmogorov [122, 123] established that every continuous multivariate function can be expressed as a superposition of continuous univariate functions. Specifically, for any $f:[0,1]^{n}\to\mathbb{R}$,
$f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\underbrace{\Phi_{q}\!\Bigg(\,\sum_{p=1}^{n}\phi_{q,p}(x_{p})\Bigg)}_{\text{outer nonlinear map}},$ (4)
where each inner map $\phi_{q,p}:[0,1]\to\mathbb{R}$ and each outer map $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ is continuous. A key feature is that the inner functions $\{\phi_{q,p}\}$ can be chosen universal, i.e., independent of the specific target $f$, while the outer functions $\{\Phi_{q}\}$ encode dependence on $f$ itself [124, 125].
In following, we use $C(\cdot)$ to denote the space of continuous functions. The class ${\rm Lip}^{\alpha}$ refers to Hölder/Lipschitz functions of order $\alpha\in(0,1]$: there exists $L>0$ such that
$|u(x)-u(y)|\leq L\,|x-y|^{\alpha}.$
The special case $\alpha=1$ corresponds to the standard Lipschitz condition. Fixed positive weights are written $\lambda_{p}$, shifts as $\eta_{q}$, offsets as $c_{q}$, and the notation $\nearrow$ indicates strictly increasing functions.
Simplified constructive variants. Since Kolmogorov’s original proof used highly irregular functions, later refinements proposed more structured forms:
  • Lorentz [126]. The representation can be simplified by using a single outer function $\Phi$ applied to weighted sums of smoother inner functions:
    $f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\Phi\!\Bigg(\sum_{p=1}^{n}\lambda_{p}\,\psi_{q}(x_{p})\Bigg),\quad\psi_{q}\in{\rm Lip}^{\alpha},\;\;\psi_{q}\nearrow.$
  • Sprecher [129]. The inner layer can be compressed further by reusing a single inner function $\phi$, shifted and offset differently across terms:
    $f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\Phi\!\Bigg(\sum_{p=1}^{n}\lambda_{p}\,\phi(x_{p}+\eta_{q})+q\Bigg).$ (5)
  • Ostrand [130]. Extended the two-layer superposition representation beyond cubes $0,1]^{n}$ to general compact metric spaces, broadening its applicability.
Note that the exact inner functions constructed in KAT are often pathological [131]. Modern refinements instead approximate both $\phi$ and $\Phi$ using smooth, learnable parameterizations—an idea that directly motivates KANs.
From KAT to KAN. KANs [27] translate the KAT blueprint into a trainable neural layer by learning the univariate maps and summing them:
$x^{(\ell+1)}_{q}=\sum_{p=1}^{P}\underbrace{\varphi^{(\ell)}_{q,p}\!\big(x^{(\ell)}_{p}\big)}_{\begin{subarray}{c}\text{edge-wise}\\ \text{1D transform}\end{subarray}},\quad q=1,\dots,Q,$ (6)
where $x^{(\ell)}_{p}$ is the $p$-th coordinate of the input at layer $\ell$, $P$ is the input dimension, and $Q$ is the output dimension. Compared with MLPs (which first form a linear combination and then apply a pointwise activation), KANs apply per-coordinate univariate transforms first and then aggregate additively.
Learning the 1D maps. A common parameterization is a basis expansion, e.g., with B-spline functions:
$\varphi^{(\ell)}_{q,p}(x)=\sum_{k=1}^{K}\underbrace{c^{(\ell)}_{q,p,k}}_{\begin{subarray}{c}\text{learnable}\\ \text{coefficient}\end{subarray}}\underbrace{B_{k}(x)}_{\begin{subarray}{c}\text{basis}\\ \text{function}\end{subarray}},$ (7)
where $\{B_{k}\}_{k=1}^{K}$ are spline basis functions, and their locality/smoothness yields compact and interpretable representations of $\varphi$. Other bases (Chebyshev, Jacobi, Fourier, radial basis functions, or ReLU powers) can be substituted directly; the structure in (6) remains unchanged, only the parameterization of the univariate maps $\phi$ or $\varphi$ differs.
Remark. Earlier depth-2, width-(2n+1)(2n{+}1) realizations of KAT were difficult to optimize in practice [132, 133, 134, 135, 136]; modern KANs address this via differentiable basis functions and end-to-end training [27].
4 Bridging KANs and MLPs
KANs and MLPs are both hierarchical function approximators, yet they differ fundamentally in where and how nonlinearities are applied. In a conventional MLP, inputs are first linearly mixed and then passed through a fixed activation function (mix $\rightarrow$ activate). In contrast, a KAN applies a (typically trainable) univariate transformation to each input coordinate before aggregation (activate $\rightarrow$ mix). This reversal of operations produces more localized, interpretable, and adaptable mappings.
This structure closely parallels Sprecher’s constructive version of the Kolmogorov superposition theorem (5), where inner univariate functions act independently on shifted coordinates:
$f(x_{1},\ldots,x_{n})=\sum_{q}\Phi\!\biggl(\smash{\underbrace{\sum_{p}\lambda_{p}\,\phi(x_{p}+\eta_{q})}_{\begin{subarray}{c}\lambda_{p}:\text{ weights}\\ \eta_{q}:\text{ shifts}\end{subarray}}}\;+\;\smash{\underbrace{c_{q}}_{\text{offset}}}\biggr),$
with positive weights $\lambda_{p}>0$, shifts $\eta_{q}$, and offsets $c_{q}$. The KAN formulation operationalizes this constructive principle within a neural architecture, turning theoretical decomposition into a learnable process.
Formal Equivalence. Wang et al. [45] establish a bidirectional correspondence between MLPs and KANs. Any MLP with $\mathrm{ReLU}^{k}$ activations,
$\sigma_{k}(x)=\max(0,x)^{k},$
can be represented by an equivalent KAN parameterization using B-spline bases of order $k$ and grid size G=2G{=}2. For an MLP approximating a target function $f:\mathbb{R}^{d}\!\to\!\mathbb{R}$,
[…]
there exists an equivalent KAN representation,
[…]
where each univariate function $\varphi_{ij}:\mathbb{R}\!\to\!\mathbb{R}$ is a learned B-spline that approximates $\sigma_{k}(w_{ij}x_{j}+b_{i})$. In essence, KANs replace the fixed nonlinearities of MLPs with learnable one-dimensional functions, enabling richer local adaptation without increasing architectural depth.
The reverse direction also holds under mild constraints. If a KAN layer employs polynomial-type univariate functions (e.g., B-splines of order $k$) and excludes non-polynomial activations such as SiLU, the model can be reformulated as an equivalent MLP with $\mathrm{ReLU}^{k}$ activations. In this conversion, a KAN of width $W$, grid size $G$, and depth $L$ induces an MLP whose effective width grows as $(G+2k+1)W^{2}$, resulting in a parameter complexity of $\mathcal{O}(G^{2}W^{4}L)$. By comparison, the original KAN requires only $\mathcal{O}(GW^{2}L)$ parameters—highlighting its superior parameter efficiency, especially for fine grids or high-order spline bases. This equivalence creates a bridge through which approximation and convergence results from MLP theory can be transferred directly to KANs (see Section 10).
Special Case: Piecewise-Linear Functions. Schoots et al. [137] demonstrate that KANs using piecewise-linear univariate functions are functionally identical to ReLU-based MLPs. Any piecewise-linear $\varphi:\mathbb{R}\to\mathbb{R}$ with $k$ breakpoints admits the expansion
$\varphi(x)=\underbrace{a_{0}}_{\text{bias}}+\underbrace{a_{1}x}_{\text{linear term}}+\sum_{j=1}^{k}\underbrace{\alpha_{j}}_{\text{coeff.}}\,\mathrm{ReLU}\!\bigl(x-b_{j}\bigr),$
where $a_{0},a_{1},\alpha_{j},b_{j}\in\mathbb{R}$. Thus, each KAN univariate map can be embedded into a compact ReLU subnetwork, reinforcing the view of KANs as structured and interpretable MLPs.
Actor et al. [38] and Gao et al. [138] reach similar conclusions: KANs can emulate MLP behavior while introducing task-adaptive nonlinearities—such as spline, Fourier, or Chebyshev bases—that improve inductive bias and generalization for structured data.
Summary. KANs and MLPs are expressively equivalent but structurally distinct. KANs achieve localized, interpretable representations by applying learnable univariate transformations before aggregation, whereas MLPs rely on fixed activations after linear mixing. This reversal leads to substantial parameter savings and smoother function representations, while maintaining theoretical consistency with the Kolmogorov–Arnold superposition principle. In practice, KANs can thus be viewed as structured, efficient, and interpretable extensions of MLPs.
5 How KANs Extend the MLP Landscape
KAN layers are increasingly adopted as drop-in replacements for traditional MLP blocks across convolutional, transformer, graph, and physics-informed models. They enhance expressivity, interpretability, and parameter efficiency—often with minimal architectural or training changes.
Vision and Representation Learning. In computer vision, Convolutional KANs embed spline-based activations directly within convolutional kernels, producing more expressive yet lightweight CNNs [139] [^4] https://github.com/AntonioTepsich/Convolutional-KANs. Residual KAN modules integrate into ResNet backbones to improve gradient flow and generalization [144] [^5] https://github.com/withray/residualKAN, while U-KAN extends U-Nets for image segmentation and diffusion models [146] [^6] https://github.com/CUHK-AIM-Group/U-KAN. Similarly, the KAN-Mixer adapts MLP-Mixer architectures for image classification [159] [^7] https://github.com/engichang1467/KAN-Mixer. Additional applications include remote sensing [148], hyperspectral imaging via Wav-KAN [147], medical image classification [149], and autoencoding tasks [145] [^8] https://github.com/SekiroRong/KAN-AutoEncoder.
Sequential and Temporal Modeling. For time-series and sequential tasks, KAN-AD employs Fourier expansions for efficient anomaly detection [150] [^9] https://github.com/issaccv/KAN-AD, while T-KAN and MT-KAN enhance forecasting under concept drift and multivariate dependencies [152]. Other extensions include satellite-traffic forecasting [151], recurrent temporal KANs (TKANs[153], and the transformer-based TKAT model [154] [^10] https://github.com/remigenet/TKAT, which integrates learnable univariate mappings within self-attention layers.
Graph and Structured Data. KAN layers have also been incorporated into graph learning frameworks, where replacing MLPs in GNN message-passing blocks improves numerical stability and feature smoothness. Implementations such as GraphKAN [155] [^11] https://github.com/WillHua127/GraphKAN-Graph-Kolmogorov-Arnold-Networks and KAN4Graph [156] [^12] https://github.com/yueliu1999/KAN4Graph achieve consistent accuracy gains. GKAN integrates spline-based kernels directly into graph convolutions [157], while general-purpose variants like S-KAN and S-ConvKAN enable task-dependent activation selection across architectures [30].
Physics-Informed and Operator Learning. In scientific computing, Physics-Informed KANs preserve the original PINN objective but replace fixed activations with learnable basis functions, improving locality and interpretability. This modularity enables direct architectural transfers—e.g., DeepONet $\rightarrow$ DeepOKAN, separable PINNs $\rightarrow$ separable PIKANs, and even NTK analyses in PINNs $\rightarrow$ corresponding analyses in PIKANs. Moreover, existing advances such as variational losses, residual reweighting, and adaptive sampling naturally carry over, bridging traditional PINN frameworks with modern kernel-based representations. Representative examples are summarized in Table 4.
Table 4: Representative advanced PINN baselines [108, 109] and corresponding KAN/PIKAN counterparts.
PINN contribution KAN/PIKAN counterpart
2017—[1] Foundational PINN for data-driven PDE solutions [27]
2018—[3] Uncertainty quantification for forward/inverse PINNs [105]
2019—[141] Neural ordinary differential equations (Neural ODEs) [36]
2019—[112] Neural operators for PDE solution maps (DeepONet) [46]
2020—[203] PINN for symbolic regression via recursive decomposition [204]
2020—[6] Multi-fidelity PINNs (low/high-fidelity correlation) [162]
2020—[115] Fourier-feature embeddings for multiscale structure [63]
2020—[211] Fourier Neural Operator (FNO) [210]
2021—[9] DeepXDE library (resampling strategies and benchmarks) [27]
2022—[8] NTK-based analysis explaining PINN training pathologies [138, 55]
2023—[121] Adaptive PINN (moving collocation points) [82]
2023—[184] Finite-basis PINNs with overlapping subdomains [86]
2023—[117] Separable PINN architectures [163]
2023—[140] Surrogate + PDE-constrained optimization [48]
2023—[142] Surrogate modeling in elasto-plasticity [54]
2025—[143] Multigrid and multi-resolution training strategy [64]
6 Basis Functions
This section reviews commonly used bases in KANs: B-splines, Chebyshev and Jacobi polynomials, ReLU compositions, Fourier series, Gaussian kernels, wavelets, finite-basis partitions, and Sinc functions.
We begin with an overview (Table 5) to orient the reader. Subsequent subsections expand on each family, beginning with B-splines.
Table 5: Comparison of Basis Functions in KANs
Name Support Equation Form Grid Basis/Activation Ref.
Required Type
B-spline Local $\sum_{n}c_{n}B_{n}(x)$ Yes B-spline [27]
Chebyshev Global $\sum_{k}c_{k}T_{k}(\tanh x)$ No Chebyshev + Tanh [169]
Stabilized Chebyshev Global $\tanh\!\Big(\sum_{k}c_{k}\,T_{k}(\tanh x)\Big)$ No Chebyshev + linear head [57]
Chebyshev (grid) Global $\sum_{k}c_{k}T_{k}\!\Big(\tfrac{1}{m}\sum_{i}\tanh(w_{i}x+b_{i})\Big)$ Yes Chebyshev + Tanh [44]
ReLU-KAN Local $\sum_{i}w_{i}R_{i}(x)$ Yes Squared ReLU [87]
HRKAN Local $\sum_{i}w_{i}[\mathrm{ReLU}(x)]^{m}$ Yes Polynomial ReLU [97]
Adaptive ReLU-KAN Local $\sum_{i}w_{i}v_{i}(x)$ Yes Adaptive ReLU [82]
fKAN (Jacobi) Global $\sum_{n}c_{n}P_{n}(x)$ No Jacobi [91]
rKAN (Padé/Jacobi) Global $\tfrac{\sum_{i}a_{i}P_{i}(x)}{\sum_{j}b_{j}P_{j}(x)}$ No Rational + Jacobi [92]
Jacobi-KAN Global $\sum_{i}c_{i}P_{i}(\tanh x)$ No Jacobi + Tanh [62]
FourierKAN Global $\sum_{k}a_{k}\cos(kx)+b_{k}\sin(kx)$ No Fourier [98]
KAF Global $\alpha\,\mathrm{GELU}(x)+\sum_{j}\beta_{j}\psi_{j}(x)$ No RFF + GELU [63]
Gaussian Local $\sum_{i}w_{i}\exp\!\Big(-\big(\tfrac{x-g_{i}}{\varepsilon}\big)^{2}\Big)$ Yes Gaussian RBF [83]
RSWAF-KAN Local $\sum_{i}w_{i}\left(s_{i}-\tanh^{2}\!\big(\tfrac{x-c_{i}}{h_{i}}\big)\right)$ Yes Reflectional Switch [84]
CVKAN Local $\sum_{u,v}w_{uv}\exp\!\big(-|z-g_{uv}|^{2}\big)$ Yes Complex Gaussian [93]
BSRBF-KAN Local $\sum_{i}a_{i}B_{i}(x)+\sum_{j}b_{j}\exp\!\big(-\tfrac{(x-g_{j})^{2}}{\varepsilon^{2}}\big)$ Yes B-spline + Gaussian [32]
Wav-KAN Local $\sum_{j,k}c_{j,k}\,\psi\!\big(\tfrac{x-u_{j,k}}{s_{j}}\big)$ No Wavelet [103]
FBKAN Local $\sum_{j}\omega_{j}(x)K_{j}(x)$ Yes PU + B-spline [86]
SincKAN Global $\sum_{i}c_{i}\,\mathrm{Sinc}\!\big(\tfrac{\pi}{h}(x-ih)\big)$ Yes Sinc [41]
Poly-KAN Global $\sum_{i}w_{i}P_{i}(x)$ No Polynomial [102]
6.1 B-spline
B-spline bases are among the most widely adopted in KANs due to their compact support, smoothness, local control, and piecewise–polynomial structure [38, 160, 88, 49, 53, 138, 39, 51, 52, 35, 161, 34, 162, 163, 59, 164, 165, 82, 60, 61, 58, 166, 137, 45, 167, 50, 168, 48, 86, 79, 179, 217, 210, 37, 212]. Their locality and flexibility make them expressive, numerically stable, and particularly well suited for interpretable representations. The original KAN formulation [27] and its accompanying software package further contributed to their early adoption and widespread use.
Each univariate KAN map $\varphi:\mathbb{R}\to\mathbb{R}$ is represented as a linear combination of fixed–knot B-spline basis functions:
$\varphi(x)\;=\;\sum_{n=0}^{N-1}c_{n}\,B_{n}^{(k)}(x),$ (8)
where $N$ is the number of basis functions, $B_{n}^{(k)}(x)$ denotes the $n$-th B-spline of polynomial degree $k$ defined on a knot vector $\mathbf{t}=(t_{0},\ldots,t_{N+k})$, and $c_{n}\in\mathbb{R}$ are trainable coefficients. The Cox–de Boor recursion yields $C^{k-1}$ continuity, and open–uniform knot choices give well–behaved boundary conditions.
To match boundary and interior regularity/support, the knot vector is often extended by $k$ points on each side. This padding allows the model to progressively capture finer details [45]. As illustrated in Figure 1, the number of active basis functions grows from GkG-k to $G+k$, where $G$ is the number of internal intervals in the original, non–extended grid.
Subfigures 1(a) and 1(b) visualize cubic (k=3k{=}3) B-spline bases on non–extended and extended grids, respectively. Without extension, boundary bases are truncated and lose symmetry; with extension, added knots outside the domain restore full polynomial support (gray bands). To compare expressivity, subfigures 1(c) and 1(d) synthesize univariate maps
$\varphi(x)\;=\;\sum_{n=0}^{N-1}c_{n}B_{n}^{(k)}(x),$
using the same coefficient vector across the two grids (truncated in the non–extended case). Differences in $\varphi$ thus arise solely from the basis itself: the extended grid yields richer boundary behavior and overall smoother maps, while the non–extended grid flattens near edges. This highlights the practical benefit of grid extension for boundary–sensitive tasks (e.g., PDEs).
(a) Non–extended cubic B-spline basis
(b) Extended cubic B-spline basis
(c) Learned map $\varphi(x)$ on non–extended grid
(d) Learned map $\varphi(x)$ on extended grid
Figure 1: Comparison of B-spline bases and synthesized univariate maps with/without grid extension. (a,b) Cubic (k=3k{=}3) bases on non–extended vs. extended grids. (c,d) Learned maps $\varphi(x)$ using identical random coefficients $c_{n}$. Gray regions mark padded boundary intervals.
In the original KAN [27] and KAN 2 [77] [^13] https://github.com/KindXiaoming/pykan, each univariate activation uses uniformly spaced cubic B-splines ($k=3$) together with a smooth residual term to aid gradients and enhance expressivity in flat regions:
$\tilde{\varphi}(x)\;=\;\sum_{n=0}^{N-1}c_{n}\,B_{n}^{(k)}(x)\;+\;\frac{x}{1+e^{-x}}.$
A global tanh is applied after each hidden layer (but not the output) to keep activations within the spline domain. KAN 2 further supports post–training grid refinement, which increases knot resolution without reinitializing parameters, thereby improving fidelity with minimal overhead. Further discussion on efficiency and accuracy improvements of B-spline KANs can be found in Sec. 7 and Sec. 8, respectively.
6.2 Chebyshev Polynomials
Chebyshev polynomials offer a clean and theoretically grounded alternative to B-splines for KAN univariate mappings—especially in PIKANs—due to their global orthogonality, excellent spectral approximation properties, and simple recursive structure suited for smooth or oscillatory targets.
Definition and Recurrence. Let $T_{k}(x)$ denote the Chebyshev polynomials of the first kind, defined recursively as
$T_{0}(x)=1,\qquad T_{1}(x)=x,\qquad T_{k}(x)=2x\,T_{k-1}(x)-T_{k-2}(x)\quad(k\geq 2),$ (9)
which form an orthogonal basis on [1,1][-1,1] with weight $w(x)=(1-x^{2})^{-1/2}$.
ChebyKAN Formulation. The ChebyKAN model [94] [^14] https://github.com/SynodicMonth/ChebyKAN represents each univariate map as
$\varphi^{(\ell)}_{q,p}(x)\;=\;\sum_{k=0}^{K}c^{(\ell)}_{q,p,k}\,T_{k}\!\big(\tilde{x}\big),\qquad\tilde{x}=\tanh(x)\ \ \text{(per-layer input normalization)},$
with trainable coefficients $c^{(\ell)}_{q,p,k}\in\mathbb{R}$. Each layer thus evaluates $T_{k}$ on normalized inputs $\tilde{x}\in[-1,1]$ for numerical stability, producing the output
[…]
consistent with the generic KAN layer formulation (6).
Efficient Evaluation. In practice, most implementations (e.g., [64]) compute $T_{k}$ via
$T_{k}(z)=\cos\!\big(k\,\arccos z\big),$
which vectorizes efficiently on GPUs. Alternatively, the recurrence relation (9) or Clenshaw’s algorithm provides equivalent, numerically stable computation—particularly advantageous for high-degree expansions where repeated $\arccos$ calls become costly. Mahmoud et al. [42] further employ shifted Chebyshev polynomials on $[0,1]$, defined as
$T_{n}(x)=\cos\!\big(n\,\arccos(2x-1)\big),$
to align the domain with the standard Kolmogorov–Arnold representation on non-symmetric intervals.
Normalization and Stability. Figure 2 illustrates how per-layer $\tanh$ normalization stabilizes Chebyshev activations. Panel (a) shows the standard basis $T_{k}(x)$ on [1,1][-1,1] for degree K=8K{=}8, where extrema occur at $T_{k}(\pm 1)=\pm 1$, forcing steep slopes near the boundaries. Panel (b) instead evaluates $T_{k}(\tanh x)$, compressing the effective range to $\tanh(\pm 1)\approx\pm 0.762$, which moderates endpoint slopes while preserving interior structure. This compression prevents saturation at $\pm 1$ and keeps features in a well-conditioned range—critical for deep stacks of Chebyshev layers.
The difference becomes clearer in Panel (c), comparing the deep composite maps
$\sum_{k=0}^{K}c_{k}\,T_{k}(\tanh x)\quad\text{and}\quad\sum_{k=0}^{K}c_{k}\,T_{k}(x)$
with identical coefficients $c_{k}$. Both are bounded, but by the chain rule,
$\frac{d}{dx}\!\left[\sum_{k=0}^{K}c_{k}\,T_{k}(\tanh x)\right]=\bigl(1-\tanh^{2}x\bigr)\,\sum_{k=0}^{K}c_{k}\,T_{k}^{\prime}\!\big(\tanh x\big),$
the $\tanh$-normalized version suppresses slope growth near $|x|\approx 1$, removing the synchronized oscillations visible in the unnormalized case. Even without training, this normalization yields smoother and better-conditioned responses.
(a) Chebyshev basis without $\tanh$ normalization
(b) Chebyshev basis with $\tanh$ normalization
(c) Deep Chebyshev KAN map: with vs. without $\tanh$ normalization
Figure 2: (a) Standard Chebyshev basis functions without per-layer $\tanh$ normalization. (b) Same basis with $\tanh$ normalization, showing compressed input range and moderated edge slopes. (c) Deep Chebyshev KAN map (K=8K{=}8) comparing $\sum c_{k}T_{k}(\tanh x)$ (blue) vs. $\sum c_{k}T_{k}(x)$ (orange); the normalized version exhibits smoother behavior and smaller endpoint gradients, indicating improved conditioning.
Grid-Averaged Variant. Following Toscano et al. [44], a practical enhancement introduces a small learnable grid before evaluating the Chebyshev expansion. For a univariate input $x$,
$g(x)=\frac{1}{m}\sum_{i=1}^{m}\tanh\!\big(w_{i}x+b_{i}\big),\qquad\varphi^{(\ell)}_{q,p}(x)=\sum_{k=0}^{K}c^{(\ell)}_{q,p,k}\,T_{k}\!\big(g(x)\big),$
where $m$ denotes the number of centers, $b_{i}$ are uniformly spaced biases within $\beta_{\min},\beta_{\max}]$ (e.g., 0.1-0.1 to $0.1$), and $w_{i}$ control local slope. The coefficients $c^{(\ell)}_{q,p,k}$ can follow the stable initialization proposed in [169]. This averaged-$\tanh$ grid constrains $g(x)\in(-1,1)$, introduces mild spatial warping, improves numerical conditioning, and—as reported in [44]—supports stable training even with higher polynomial degrees.
Spectral Properties and Efficiency. The spectral parameterization of Chebyshev-based KANs was first formalized in Sidharth et al. [169], while Guo et al [170] demonstrated superior parameter efficiency and generalization in data-scarce regimes. From a theoretical perspective, Faroughi and Mostajeran [56] showed that Chebyshev PIKANs (cPIKANs) yield better-conditioned NTKs with slower spectral decay, accelerating convergence for PDEs such as diffusion and Helmholtz equations.
Stabilization and Hybrid Designs. Yu et al. [41] confirmed these advantages for function approximation and PDE learning but also highlighted failure modes of raw polynomial stacks at high depth, motivating stabilizers such as domain normalization, nested nonlinearities, and contractive mappings. Building on these insights, Daryakenari et al. [57] proposed a stabilized Chebyshev stack by inserting an additional $\tanh$ between layers and replacing the Chebyshev head with a linear readout:
$x^{(\ell+1)}_{q}\;=\;\tanh\!\Bigg(\sum_{p=1}^{P}\sum_{k=0}^{K}c^{(\ell)}_{q,p,k}\,T_{k}\!\big(\tanh(x^{(\ell)}_{p})\big)\Bigg),\qquad\ell=0,\dots,L-1,$ (10)
with the network output
$\hat{f}(\mathbf{x})\;=\;\mathbf{W}^{\text{out}}\,\mathbf{x}^{(L)}+\mathbf{b}^{\text{out}},$ (11)
where $\mathbf{x}^{(L)}=[x^{(L)}_{1},\dots,x^{(L)}_{H_{L}}$$^{\top}$. The inter-layer $\tanh$ in (10) acts as a contraction, curbing gradient growth [41], while the linear head (11) isolates the final mapping from additional polynomial expansions, improving stability in inverse and PDE learning tasks [57].
Summary. Chebyshev-based KANs combine spectral approximation theory with practical stability mechanisms. Their global orthogonality and efficient recursion yield compact and interpretable representations, while nested nonlinearities and grid-averaged normalization ensure stable deep training. These features make them particularly well suited for scientific computing, operator learning, and inverse problems—complementing and often surpassing spline-based KANs within the broader KAN framework.
6.3 ReLU
ReLU–based KANs (ReLU-KANs) were introduced by Qiu et al. [87] [^15] https://github.com/quiqi/relu_kan as a hardware–efficient alternative to B-spline KANs. The key idea is to replace spline activations with compactly supported, bell–shaped functions constructed from ReLU compositions. This preserves the localized, compositional spirit of Kolmogorov–Arnold models while enabling fast, GPU–friendly primitives.
Local ReLU bases and univariate maps. In ReLU-KANs, each univariate map $\varphi:\mathbb{R}\to\mathbb{R}$ is a weighted sum of compact local bases:
$\varphi(x)\;=\;\sum_{i=0}^{G+k-1}w_{i}\,R_{i}(x),$
where $w_{i}\in\mathbb{R}$ are trainable weights, $G$ is the number of grid intervals, and $k$ controls overlap among neighboring bases. Each $R_{i}$ is supported on $s_{i},e_{i}]$ with uniformly spaced endpoints
$s_{i}\;=\;\frac{i-k}{G},\qquad e_{i}\;=\;s_{i}+\frac{k+1}{G},\quad i=0,\ldots,G+k-1.$
Inside its support, $R_{i}$ has a smooth bell shape (Figure 3):
$R_{i}(x)=\begin{cases}0,&x<s_{i},\\[2.0pt] \big((x-s_{i})(e_{i}-x)\big)^{2}\,\dfrac{16}{(e_{i}-s_{i})^{4}},&s_{i}\leq x\leq e_{i},\\[6.0pt] 0,&x>e_{i},\end{cases}$
which can be written equivalently using squared ReLUs:
$R_{i}(x)\;=\;\Big[\operatorname{ReLU}(e_{i}-x)\,\operatorname{ReLU}(x-s_{i})\Big]^{2}\cdot\frac{16}{(e_{i}-s_{i})^{4}},\qquad\operatorname{ReLU}(x)=\max(0,x).$
(a) Step-by-step construction of ReLU basis
(b) Complete set of normalized ReLU–KAN basis functions
Figure 3: (a) Step-by-step construction of the normalized ReLU-based local basis function from its constituent $\operatorname{ReLU}(e_{i}-x)$ and $\operatorname{ReLU}(x-s_{i})$ terms. (b) Complete set of normalized ReLU–KAN basis functions $R_{i}(x)$ with supports $s_{i},e_{i}] [$87].
Speed and a smoothness caveat. ReLU-KANs often train $5\text{--}20\times$ faster than spline variants in practice [87], but the squared–ReLU construction has limited smoothness, which can hinder PDE tasks requiring higher–order derivatives.
Higher–order ReLU KAN (HRKAN). So et al. [97] generalize the squared–ReLU basis by introducing local powers of ReLU. For a finite interval $s_{i},e_{i}]$, the basis function of order $m$ is
$v_{m,i}(x)\;=\;\Big[\operatorname{ReLU}(x-s_{i})\,\operatorname{ReLU}(e_{i}-x)\Big]^{\,m}\cdot\left(\frac{2}{\,e_{i}-s_{i}\,}\right)^{\!2m},$
where $\operatorname{ReLU}(x)=\max(0,x)$ and $m\in\mathbb{Z}_{+}$ controls the interior smoothness: $v_{m,i}\in C^{m-1}$. Larger $m$ produces lobes that are more peaked in the interior and decay more smoothly to zero at the boundaries, leading to higher continuity of derivatives and potentially better performance for PDEs that require smooth high-order derivatives.
Figure 4 compares individual basis functions $v_{m,i^{\star}}$ over $s_{i^{\star}},e_{i^{\star}}]$. Panel (a) shows the squared–ReLU case ($m=2$), which is $C^{1}$ but exhibits visible jumps in the second derivative at the boundaries. Panel (b) shows a higher–order ReLU with $m=4$, which is $C^{3}$ and decays smoothly to zero, eliminating derivative discontinuities. This single-basis view makes clear how increasing $m$ improves both interior smoothness and boundary regularity.
(a) Square–of–ReLU basis
(b) Higher–order ReLU basis
Figure 4: Symbolic basis on $s_{i^{\star}},e_{i^{\star}}]$: (a) $v_{2,i^{\star}}(x)$ with first and second derivatives; (b) $v_{4,i^{\star}}(x)$ with first and second derivatives. Observation: $v_{m,i}$ is globally $C^{m-1}$ (with the $m$-th derivative discontinuous at $s_{i^{\star}},e_{i^{\star}}$); hence $v_{4,i^{\star}}$ ($C^{3}$) offers smoother higher-order derivatives than $v_{2,i^{\star}}$ ($C^{1}$[97].
Figure 5 compares synthesized activations
$\varphi(x)\;=\;\sum_{i}c_{i}\,v_{m,i}(x),$
constructed from the two basis types using identical random coefficients ($G=3$, $k=1$) over support of $s_{i^{\star}},e_{i^{\star}}]=[-0.6,1]$. Because the coefficients are shared, differences between the dashed ($m=2$) and solid ($m=4$) curves arise solely from the change in $m$. The higher-order case produces narrower, more sharply peaked bumps with smoother interior derivatives, while the squared–ReLU case yields broader lobes with lower differentiability. Together with Figure 4, this demonstrates both the local and global effects of the order parameter $m$ in HRKAN.
Figure 5: Synthesized activation $\varphi(x)$ from ReLU-KAN bases using shared random coefficients. Dashed: squared–ReLU basis ($m=2$), Solid: higher–order ReLU basis ($m=4$). Higher order yields sharper, smoother lobes while preserving the same coefficient structure.
Both ReLU-KAN and HRKAN originally used fixed, uniformly spaced grids. To add adaptivity, Rigas et al. [82] [^16] https://github.com/srigas/jaxKAN define bases over a nonuniform, data–dependent grid $\mathcal{G}=\{x_{0},\ldots,x_{G}\}$. With neighborhood parameter $p$,
$s_{i}\;=\;\mathcal{G}[i]-\tfrac{1}{2}\big(\mathcal{G}[i+p]-\mathcal{G}[i-p]\big),\qquad e_{i}\;=\;2\mathcal{G}[i]-s_{i}.$
Adaptive widths and centers improve resolution in regions with singularities, steep gradients, or boundary layers; their jaxKAN implementation also uses resampling and loss reweighting for heterogeneous PDEs.
6.4 Jacobi Polynomials
Polynomial families constitute one of the most general and historically established classes of basis functions in KANs. Beyond splines and Chebyshev polynomials, a comprehensive benchmark by Seydi [102] [^17] https://github.com/seydi1370/Basis_Functions systematically evaluated eighteen distinct polynomial families as KAN activation functions, offering a unified comparison across orthogonal, recurrence-based, and rational constructions. Among all tested variants, the Gottlieb polynomial achieved the highest accuracy and stability metrics on the MNIST benchmark. The surveyed families can be grouped according to their mathematical origin:
  • Classical and General Orthogonal Polynomials: Charlier and Gottlieb (discrete orthogonal), Boas–Buck and Boubaker (generalized continuous families encompassing Hermite, Laguerre, and related forms).
  • Advanced Orthogonal Polynomials (Askey Scheme & Related): Askey–Wilson and Al–Salam–Carlitz ($q$-orthogonal series), Bannai–Ito (Racah generalization).
  • Recurrence-Based and Number-Theoretic Polynomials: Tribonacci, Tetranacci, Pentanacci, Hexanacci, Heptanacci, and Octanacci (generalized Fibonacci-type recurrences); Fermat, Vieta–Pell, and Narayana (number-theoretic families).
  • Rational Constructions: Padé approximants [92], explicitly adopted in the rational Jacobi network (rKAN).
Overall, Seydi’s benchmark revealed that KAN layers can flexibly host a wide variety of polynomial structures beyond splines, broadening the design space for architectures targeting spectral, combinatorial, or rational approximation behaviors. Among these, orthogonal and number-theoretic polynomials—particularly Gottlieb—exhibited the best numerical conditioning and convergence stability. These insights motivate the deeper exploration of Jacobi-type formulations discussed below, since Jacobi, Legendre, and Chebyshev families together form the classical orthogonal polynomial hierarchy within the Askey scheme.
Fractional Jacobi Basis (fKAN). To enhance smoothness, domain flexibility, and adaptivity, Aghaei introduced the Fractional KAN (fKAN) [91] [^18] https://github.com/alirezaafzalaghaei/fKAN, employing fractional-order Jacobi polynomials as trainable univariate maps:
$P_{n}^{(\alpha,\beta)}(z_{\gamma}),\qquad z_{\gamma}=\phi_{\gamma}(x)=2\,x^{\gamma}-1,\quad x\in[0,1],$
where $\alpha,\beta>-1$ are Jacobi exponents and $\gamma>0$ is a fractional warp parameter controlling the domain stretch. The basis remains orthogonal on the canonical interval [1,1][-1,1], and different $(\alpha,\beta)$ recover classical cases: Legendre for $(0,0)$, Chebyshev of the first kind for $(-\tfrac{1}{2},-\tfrac{1}{2})$, and Chebyshev of the second kind for $(\tfrac{1}{2},\tfrac{1}{2})$, as illustrated in Fig. 6.
Figure 6: Degree-$n=6$ polynomials on $z\in[-1,1]$: Chebyshev $T_{6}(z)$, Chebyshev $U_{6}(z)$, Legendre $P_{6}(z)$, and Jacobi $P_{6}^{(2.5,2.5)}(z)$. The Chebyshev–$T$ curve coincides with $T_{6}(z)$ in Fig. 2(a).
In fKAN, the univariate activation is expressed as
$\varphi(x)\;=\;P_{n}^{(\alpha,\beta)}\!\big(\,\phi_{\gamma}(\,x(r)\,)\,\big),$
where $r$ is the raw input, $x(r)\in[0,1]$ denotes a normalization map (linear, sigmoid, or $\tfrac{1}{2}(1+\tanh r)$), and $\phi_{\gamma}(x)=2x^{\gamma}-1$ applies fractional warping. The parameters $\alpha,\beta,\gamma$ are trainable, and positivity of $\alpha,\beta$ is enforced via ELU or sigmoid reparameterizations. This yields a globally smooth and tunable basis that performs effectively on regression and PDE benchmarks. Kashefi [62] implements fKAN using the $\tanh$-based normalization $x(r)=\tfrac{1}{2}(1+\tanh r)$ for stable input scaling.
Effect of Input Normalization. Implementations typically evaluate $P_{n}^{(\alpha,\beta)}(z_{\gamma})$ with $z_{\gamma}=2\,x(r)^{\gamma}-1$. Figure 7 illustrates how the normalization choice shapes the activation. Fixing $(\alpha,\beta)=(2.5,2.5)$, $n=2$, and $\gamma=0.1$, a linear mapping preserves dynamic range but produces steep edge slopes, whereas sigmoid and $\tfrac{1}{2}(1+\tanh r)$ compress the tails, improving conditioning. A smaller $\gamma<1$ increases the warp toward +1+1, further sharpening features near that end of the domain.
Figure 7: Degree-$n=2$ Jacobi polynomial with $(\alpha,\beta)=(2.5,2.5)$, $\gamma=0.1$: $\varphi(r)=P^{(\alpha,\beta)}_{2}\!\big(2\,x(r)^{\gamma}-1\big)$ under three normalizations $x(r)\in[0,1]$: linear $x=(r-r_{\min})/(r_{\max}-r_{\min})$ (dashed), sigmoid $x=\sigma(r)$, and $\tanh$-based $x=\tfrac{1}{2}(1+\tanh r)$. All curves are rescaled by $\varphi(r{=}1)$ for comparability.
Rational Extensions (rKAN). To further expand expressivity, Aghaei proposed the Rational KAN (rKAN) [92] [^19] https://github.com/alirezaafzalaghaei/rKAN, which generalizes fKAN by incorporating rational Jacobi compositions. Two representative variants are:
  • Padé–rKAN: a rational quotient of Jacobi expansions,
    $\varphi(x)=\frac{\sum_{i=0}^{p}w_{i}^{(P)}P_{i}^{(\alpha_{p},\beta_{p})}\!\big(\phi(\sigma(x))\big)}{\sum_{j=0}^{q}w_{j}^{(Q)}P_{j}^{(\alpha_{q},\beta_{q})}\!\big(\phi(\sigma(x))\big)},$
    where $w_{i}^{(P)}$ and $w_{j}^{(Q)}$ are trainable weights, and $\sigma(x)$ is a squashing function.
  • Jacobi–rKAN: a rational domain warping approach,
    $\varphi(x)=P_{n}^{(\alpha,\beta)}\!\big(\phi(x;\iota)\big),$
    with $\phi(x;\iota)$ representing a rational transform parameterized by $\iota>0$ (often implemented using SoftPlus for positivity).
Both fKAN and rKAN exploit Jacobi polynomials with adaptive shape and domain parameters: fKAN emphasizes fractional-order smoothness and controlled warping, while rKAN introduces rational expressivity and sharper nonlinear transitions.
Classical Fixed-Degree Jacobi KANs. Several works retain fixed polynomial degrees ($\gamma=1$) and use only $\tanh$ input compression into [1,1][-1,1]. Each univariate activation is then represented as
$\varphi^{(l)}_{ij}(x)=\sum_{k=0}^{K}c^{(l)}_{k,ij}\,P_{k}^{(\alpha,\beta)}\!\big(\tanh(x)\big),$
with trainable coefficients $c^{(l)}_{k,ij}$ and often fixed $(\alpha,\beta)$. Kashefi [62] [^20] https://github.com/Ali-Stanford/KAN_PointNet_CFD reported that low-degree expansions ($K=2$) with Chebyshev-like parameters $\alpha=\beta=-0.5$ achieve an optimal balance between accuracy and numerical stability, suppressing Runge oscillations in inverse PDEs with sparse boundary data. Shukla et al. [29] applied the same formulation within a PIKAN framework for high–Reynolds cavity flow, using $(\alpha,\beta)=(1,1)$ and degrees K=3K=388 depending on the regime. While higher degrees increase computational cost, their Jacobi–based PIKANs matched or exceeded conventional PINNs in accuracy.
Building upon this foundation, Xiong et al. [65] proposed the Jacobian Orthogonal Polynomial–based KAN (J–PIKAN), enforcing explicit orthogonality among polynomial components to ensure stable layerwise synthesis and interpretable coefficient updates. Leveraging the three-term recurrence of Jacobi polynomials, J–PIKAN efficiently evaluates higher-degree expansions and maintains numerical stability. Using fluid-dynamics benchmarks, their study compared B-spline, Fourier, Hermite, Legendre, Chebyshev, and Jacobi bases within a unified physics-informed framework, identifying Jacobi bases with moderate parameters ($\alpha=\beta\approx 2$) as offering the best compromise between accuracy, conditioning, and convergence rate.
In parallel, Zhang et al. [66] [^21] https://github.com/zhang-zhuo001/Legend-KINN explored the Legendre specialization $(\alpha,\beta)=(0,0)$ within Jacobi-based KANs, confirming its favorable conditioning and rapid convergence for physics-informed PDE solvers.
Summary. Jacobi-based KANs form a versatile and mathematically rich subclass within polynomial KANs. Classical (fixed-degree) variants emphasize numerical stability and geometric fidelity, while fractional (fKAN) and rational (rKAN) extensions introduce adaptive domain warping and enhanced expressivity. Together, they provide a cohesive Jacobi toolkit—bridging spectral approximation, rational modeling, and physics-informed learning within the broader KAN framework.
6.5 Gaussian (RBF)
As researchers sought to improve the efficiency and performance of the original B–spline–based KANs, the Gaussian radial basis function (RBF) emerged as a powerful and computationally elegant alternative. Its smooth, localized, and infinitely differentiable nature makes it an excellent choice for representing univariate activation functions along the edges of a KAN. Across recent studies [83, 46, 106, 84, 204, 80], Gaussian–based activations have appeared in several formulations—ranging from fixed–grid hybrids to fully learnable and reflectional variants—each striking a different balance between stability, adaptivity, and efficiency.
Foundational Gaussian Layer. The fundamental univariate Gaussian basis function is defined as
$\varphi(x;c,\varepsilon)=\exp\!\left[-\Big(\tfrac{x-c}{\varepsilon}\Big)^{2}\right],$ (12)
where $c$ is the center and $\varepsilon>0$ is the width (shape parameter). A general univariate activation is then constructed as a linear combination
$\phi(x)=\sum_{i=0}^{G-1}w_{i}\,\varphi(x;g_{i},\varepsilon),$ (13)
with trainable coefficients $w_{i}$ and equispaced grid centers $g_{i}\in[a,b]$. This formulation replaces the piecewise polynomial behavior of splines with globally smooth Gaussian features while retaining locality and analytical gradients.
Hybrid Gaussian–Residual Formulation (Fixed Grid).
Gaussian RBFs are a fast, smooth alternative to B–splines for KAN activations. The key observation is that cubic B–splines can be closely approximated by scaled Gaussians. Li et al. [83] [^22] https://github.com/ZiyaoLi/fast-kan proposed FastKAN, replacing the cubic B–spline $\mathcal{B}_{3}(x-c_{i})$ with
$\mathcal{B}_{3}(x-c_{i})\;\approx\;\lambda\,\exp\!\left[-\Big(\tfrac{x-c_{i}}{\sigma}\Big)^{2}\right],$ (14)
where $\lambda\in\mathbb{R}$ is a scaling constant and $\sigma\approx h$ matches the grid spacing. As shown in Figure 8, this preserves spline-like locality and smoothness while reducing computational cost.
Figure 8: Comparison of cubic B–spline basis functions $\mathcal{B}_{3}(x)$ (non-extended) and Gaussian RBFs $\exp[-((x-c_{i})/\sigma)^{2}]$ with matched grid spacing and width.
Figure 9: Learned univariate activation $\varphi(x)$ synthesized from Gaussian RBF and cubic B–spline bases (Figure 8) using identical random coefficients.
Figure 9 demonstrates that both bases yield nearly identical activation profiles with only minor differences near peaks. Each basis decays smoothly near the boundaries, producing flat tails where support vanishes—confirming Gaussian RBFs as faithful, efficient surrogates for B–splines in KAN layers.
Building upon this foundation, Li [83] introduced the FastKAN model, where Gaussian basis functions are combined with a simple residual activation such as SiLU (Swish) to form a hybrid composite activation:
$\Phi(x)=\underbrace{\sum_{i=0}^{G-1}w_{i}\,\exp\!\Big[-\Big(\tfrac{x-g_{i}}{\varepsilon}\Big)^{2}\Big]}_{\text{RBF component}}+\,w_{b}\,\sigma(x),$ (15)
where $\sigma(x)$ is a base activation and $w_{b}\in\mathbb{R}$ modulates its contribution. The Gaussian component captures local variations, while the residual stabilizes optimization and accounts for global trends. Here, both the grid $\{g_{i}\}$ and bandwidth $\varepsilon=\tfrac{b-a}{G-1}$ are fixed, ensuring efficient and stable evaluation. This configuration, later adopted in LeanKAN [106] and RBFKAN [80], represents a strong, stable baseline for Gaussian–based KANs.
Pure RBF Layers with Learnable Centers. A more flexible formulation introduces learnable Gaussian centers, as proposed by Abueidda et al. [46] [^23] https://github.com/DiabAbu/DeepOKAN. This allows the basis functions to adapt spatially during training. Each activation takes the form
$\varphi(x)=\exp\!\left[-\Big(\tfrac{x-g_{j}}{\beta}\Big)^{2}\right],$ (16)
where $g_{j}\in\mathbb{R}$ are trainable centers and $\beta>0$ controls the receptive field width. The corresponding layer transformation is
$x_{l+1}=W_{l}\,\varphi(x_{l};G_{l},\beta),$ (17)
where $W_{l}$ are learnable synthesis weights and $G_{l}=\{g_{j}\}$ denotes the center set. By optimizing both $W_{l}$ and $G_{l}$, the network gains geometric adaptivity—allowing basis functions to migrate toward regions of high functional complexity—while preserving the smoothness and differentiability of Gaussian kernels.
Reflectional Gaussian Approximations (RSWAF). To further reduce computational cost, several works have replaced the exponential kernel with an algebraically similar Reflectional Switch Activation Function (RSWAF) proposed by Delis [84] [^24] https://github.com/AthanasiosDelis/faster-kan and extended in Bühler et al. [204]. This activation approximates the Gaussian bell using the hyperbolic–secant squared function:
$\varphi_{\text{RSWAF}}(x;c,h)=1-\tanh^{2}\!\Big(\tfrac{x-c}{h}\Big)=\mathrm{sech}^{2}\!\Big(\tfrac{x-c}{h}\Big),$ (18)
which preserves reflectional symmetry and smooth decay. Both the centers $c_{i}$ and widths $h_{i}$ may be fixed or trainable, controlled by configuration flags (e.g., train_grid, train_inv_denominator) in implementations such as FasterKAN [84]. The full activation is often expressed as
$\phi(x)=\sum_{i}w_{i}\,s_{i}\,\bigg[1-\tanh^{2}\!\Big(\tfrac{x-c_{i}}{h_{i}}\Big)\bigg],$ (19)
where $w_{i}$ are learnable amplitudes, $s_{i}$ are reference scaling factors, and both $c_{i}$ and $h_{i}$ evolve during training. This formulation yields a localized, computationally efficient surrogate that mimics Gaussian curvature while offering explicit control over both shape and location.
Hybrid and Unified Gaussian Extensions. Koenig et al. [106] [^25] https://github.com/DENG-MIT/LeanKAN proposed a hybrid activation that mixes RBFs with a base nonlinearity:
$\phi(x)\;=\;\sum_{i=1}^{N}w_{i}\,\varphi(x-c_{i})\;+\;w_{b}\,b(x),$ (20)
where $\varphi(\cdot)$ is a Gaussian RBF, $b(x)$ is a base activation (e.g., Swish or linear), and $w_{b}$ modulates its contribution. Related implementations such as RBFKAN [80] [^26] https://github.com/sidhu2690/RBF-KAN also apply simple min–max normalization,
$x_{\text{norm}}\;=\;\frac{x-x_{\min}}{x_{\max}-x_{\min}},$ (21)
to stabilize input scaling.
The BSRBF–KAN model [32] [^27] https://github.com/hoangthangta/BSRBF_KAN unifies B–splines and Gaussian RBFs within a single layer:
$\mathrm{BSRBF}(x)\;=\;w_{b}\,b(x)\;+\;w_{s}\big(\mathrm{BS}(x)+\mathrm{RBF}(x)\big),$ (22)
where $\mathrm{BS}(x)$ and $\mathrm{RBF}(x)$ are evaluated at identical grid locations, and $w_{b},w_{s}\in\mathbb{R}$. This configuration inherits the local compactness of B–splines and the global smoothness of Gaussian functions.
Complex–Valued Gaussian Extensions. In the complex domain, Wolff et al. [93] [^28] https://github.com/M-Wolff/CVKAN introduced the Complex–Valued KAN (CVKAN), extending FastKAN with concepts from complex–valued neural networks (CVNNs) [205]. The residual pathway uses a complex SiLU activation:
$b_{\mathbb{C}}(z)=\mathrm{CSiLU}(z)=\mathrm{SiLU}(\mathrm{Re}(z))+i\,\mathrm{SiLU}(\mathrm{Im}(z)),$ (23)
and the Gaussian RBF is generalized to complex inputs [206]:
$\mathrm{RBFC}(z)=\sum_{i=1}^{N}w_{i}\,\varphi(|z-c_{i}|),\qquad c_{i}\in\mathbb{C},\;w_{i}\in\mathbb{C}.$ (24)
The resulting activation combines both terms with a bias term:
$\phi(z)=w_{s}\,\mathrm{RBFC}(z)+w_{b}\,b_{\mathbb{C}}(z)+\beta,\qquad\beta\in\mathbb{C}.$ (25)
More recently, Che et al. [207] extended this formulation by introducing a ModELU–based CVKAN, replacing the split–type CSiLU with a modulus–preserving residual and learnable RBF shape parameters, supported by a formal complex–valued Kolmogorov–Arnold theorem.
Practical caveat: the shape parameter. Gaussian RBF performance is highly sensitive to the width parameter ($\varepsilon$), which governs both smoothness and conditioning [173, 174, 175]. Even small changes can strongly impact accuracy and numerical stability [176, 177, 178]. From a kernel–theoretic perspective, this sensitivity reflects an uncertainty principle: larger shape (width) parameters produce flatter, globally correlated kernels that improve smoothness and accuracy but worsen conditioning, whereas smaller parameters yield well–conditioned, highly localized bases at the expense of approximation power [171, 172]. While heuristic tuning and learnable widths (as in FastKAN and DeepONet–RBF variants) offer partial remedies, robust and theoretically guided selection of this parameter remains an active research problem at the intersection of kernel methods and neural approximation.
Summary. Gaussian–based KANs form a versatile class of architectures bridging spline–based and kernel–theoretic perspectives. Starting from fixed equispaced Gaussian grids [83, 106], through adaptive RBF layers with learnable centers [46], to reflectional Gaussian and complex–valued extensions [84, 204, 93, 207], these models collectively demonstrate how Gaussian basis functions can serve as both efficient computational surrogates and expressive functional primitives within the KAN framework.
6.6 Fourier
Fourier features provide global, smooth, periodic expressivity that complements localized bases (e.g., B–splines, RBFs). Within the KAN family, Fourier–based KANs replace the univariate edge functions with harmonic expansions, retaining the KAN wiring while improving spectral resolution on tasks with strong periodic structure [98]. By contrast, Fourier networks such as FAN are not KANs: they embed cosine–sine transforms directly into each layer as a drop-in MLP replacement, targeting efficient periodicity modeling (often with reduced parameters/FLOPs) and stronger in-/out-of-domain generalization [101] [^29] https://github.com/YihongDong/FAN. We mention FAN here as a related spectral approach; our focus remains on Fourier features used inside KANs.
Standard FourierKAN. In the baseline FourierKAN architecture [89] [^30] https://github.com/GistNoesis/FourierKAN, each univariate mapping is expressed as a truncated Fourier series,
$\varphi(x)=\sum_{k=1}^{K}\big(a_{k}\cos(kx)+b_{k}\sin(kx)\big),$ (26)
with learnable coefficients $a_{k},b_{k}\in\mathbb{R}$ and cutoff frequency $K\in\mathbb{N}$. Cosines capture even–symmetric modes; sines capture odd–symmetric modes. The resulting $\varphi(x)$ is a flexible global harmonic mixture, well–suited to periodic or high–frequency structure. Guo et al. [127] adopt this classical truncated Fourier formulation, using fixed integer harmonics with trainable sine–cosine coefficients.
Random Fourier features. A limitation of (26) is the fixed set of frequencies $k=1,\ldots,K$. Zhang et al. [63] [^31] https://github.com/kolmogorovArnoldFourierNetwork/KAF propose the Kolmogorov–Arnold–Fourier Network (KAF), replacing fixed bases with a learnable spectral embedding via Random Fourier Features (RFF):
$\psi_{\mathrm{RFF}}(\mathbf{x})=\sqrt{\tfrac{2}{m}}\,\big[\,\cos(\mathbf{x}W+\mathbf{b}),\;\sin(\mathbf{x}W+\mathbf{b})\,\big]\in\mathbb{R}^{2m},$ (27)
where $W\in\mathbb{R}^{d_{\mathrm{in}}\times m}$ is a trainable frequency matrix and $\mathbf{b}\in\mathbb{R}^{m}$ is a learnable phase vector. KAF then blends this adaptive spectral block with a smooth baseline:
$\varphi(\mathbf{x})=\bm{\alpha}\cdot\mathrm{GELU}(\mathbf{x})\;+\;\bm{\beta}\cdot V\,\psi_{\mathrm{RFF}}(\mathbf{x}),$ (28)
with $\bm{\alpha},\bm{\beta}\in\mathbb{R}^{d_{\mathrm{in}}}$ and $V\in\mathbb{R}^{d_{\mathrm{in}}\times 2m}$. The GELU term captures low–frequency structure, while the RFF term adaptively models higher–frequency content throughout training. This continuous frequency control improves regression/classification where global and local patterns coexist [63].
Figure 10 compares a Gaussian (RBF) kernel slice with both random Fourier features (27) and a truncated Fourier expansion using the full cosine–sine basis. Settings match the code: one-dimensional grid $x\in[-3\varepsilon,3\varepsilon]$ with $\varepsilon=1.0$. The RFF map uses the unbiased scaling $1/\!\sqrt{m}$. With this setting, increasing $m$ makes the RFF curve visually coincide with the Gaussian slice—clearly showing how the two are connected in practice. The Fourier expansion (26), constructed from fixed integer harmonics $\{\cos(k\gamma x),\sin(k\gamma x)\}_{k=1}^{K}$, remains strictly periodic, so its kernel slice differs structurally from the Gaussian, in contrast to RFF where frequencies are sampled from a Gaussian distribution.
Figure 10: Gaussian (RBF) kernel slice compared with Random Fourier Features (RFF,  (27)) and a truncated Fourier expansion (26). RFF uses the unbiased scaling $1/\!\sqrt{m}$ with both $\cos$ and $\sin$ terms, so that the diagonal kernel value is close to 1. As $m$ increases, the RFF slice aligns with the Gaussian, while the Fourier expansion (fixed integer harmonics) retains its periodic structure.
In addition to classical Fourier-based designs, Jiang et al. [43] [^32] https://github.com/Jim137/qkan propose the quantum-inspired Kolmogorov–Arnold Network (QKAN), where each univariate map is realized by a data re-uploading variational quantum circuit (QVAF). These circuits naturally generate Fourier-like expansions with exponentially scalable frequency modes, offering parameter-efficient alternatives to grid-based Fourier KANs while retaining the structured interpretability of classical KANs.
Summary. Fourier–based KANs—via deterministic truncations or randomized spectral embeddings—inject global harmonic structure into KAN layers. Compared with local polynomial bases such as B–splines, these global bases are particularly effective for smooth periodicities or problems with broad spectral support.
6.7 Wavelet
Wavelet-based KANs (Wav-KANs) extend the KAN framework by using localized wavelet bases, enabling multiscale representation and spatial adaptivity—especially useful for heterogeneous or hierarchical data. The architecture was introduced by Bozorgasl et al. [103] [^33] https://github.com/zavareh1/Wav-KAN and draws on both the Continuous Wavelet Transform (CWT) and Discrete Wavelet Transform (DWT).
In wavelet analysis, a signal $g(t)\in L^{2}(\mathbb{R})$ is decomposed via scaled/translated copies of a mother wavelet $\psi(t)$. The CWT is defined [181, 182] as
$C(s,\tau)\;=\;\int_{-\infty}^{\infty}g(t)\,\frac{1}{\sqrt{s}}\,\psi\!\left(\frac{t-\tau}{s}\right)\,dt,$ (29)
where $s>0$ is the scale and $\tau\in\mathbb{R}$ is the translation. The discrete analog used in signal processing is the DWT:
$a_{j}(k)\;=\;\sum_{n}g(n)\,\phi_{j,k}(n),\qquad d_{j}(k)\;=\;\sum_{n}g(n)\,\psi_{j,k}(n),$ (30)
with $\phi_{j,k}$ and $\psi_{j,k}$ the scaling and wavelet functions at resolution level $j$ and position $k$.
Wavelet activations in KAN layers. Wav-KAN embeds analytic wavelet functions directly into univariate edge functions. Each activation is parameterized by scale $s$ and translation $u$:
$\varphi(x)\;=\;\psi\!\left(\frac{x-u}{s}\right),$
where $\psi(t)$ is chosen from canonical families:
$\psi_{\text{mex}}(t)$ $=\tfrac{2}{\sqrt{3}\,\pi^{1/4}}(1-t^{2})\,e^{-t^{2}/2},$ (31)
$\psi_{\text{mor}}(t)$ $=\cos(\omega_{0}t)\,e^{-t^{2}/2},\quad\omega_{0}=5,$ (32)
$\psi_{\text{dog}}(t)$ $=-t\,e^{-t^{2}/2}.$ (33)
To compare expressivity across wavelet families, we form synchronized linear combinations
$\varphi(t)\;=\;\sum_{j=1}^{M}c_{j}\,\psi(t-\mu_{j}),$
using the same coefficients $\{c_{j}\}$ and centers $\{\mu_{j}\}$ for each $\psi$. Figure 11 shows three canonical wavelets (Mexican Hat, Morlet, DoG) centered at $t=0$. Using these bases, Figure 12 presents the synthesized activations $\varphi(t)$ built with identical $\{c_{j}\}$ and $\{\mu_{j}\}$. The Morlet activation exhibits higher-frequency oscillations, whereas Mexican Hat and DoG yield smoother bell-shaped profiles. Since coefficients are shared, differences arise solely from the intrinsic wavelet shapes.
Figure 11: Three wavelet bases centered at $t=0$: Mexican Hat, Morlet (with $\omega_{0}=5$), and Derivative of Gaussian (DoG). Each has localized support and distinct oscillatory behavior.
Figure 12: Synthesized activations $\varphi(t)=\sum_{j=1}^{M}c_{j}\psi(t-\mu_{j})$ using synchronized coefficients and centers across wavelet families. Differences in frequency and amplitude follow from each wavelet’s intrinsic shape.
Learning $s$ and $u$ per neuron lets Wav-KAN adapt receptive fields and frequency resolution, aiding generalization in, e.g., image classification [103] where feature granularity varies spatially.
Patra et al. [164] integrate Wav-KAN into a physics-informed framework for coupled nonlinear PDEs, reporting faster convergence than spline-based Efficient-KAN at comparable accuracy under data-free collocation. In this setting, the wavelet activation is chosen as the sine–Gaussian
$\psi_{\text{sin}}(t)\;=\;\pi^{-\tfrac{1}{4}}\sin(\omega_{0}t)\,e^{-t^{2}/2},$
with $\omega_{0}$ controlling the center frequency.
In bioinformatics, Pratyush et al. [183] propose CaLMPhosKAN for phosphorylation-site prediction, employing a DoG wavelet. This enhances detection of residue-level signals in disordered protein regions.
For hyperspectral imaging, Seydi et al. [147] combine CWT and DWT within Wav-KAN to extract spatial–spectral features, outperforming spline-KANs and MLPs on Indian Pines and Salinas datasets.
Summary. Wavelet-based KANs provide an expressive, multiscale, and interpretable basis that improves convergence and robustness across domains—ranging from PDEs and bioinformatics to high-dimensional imaging—by uniting the locality of wavelets with the compositional structure of KANs.
6.8 Finite-Basis (FBKANs)
Finite-Basis KANs, introduced by Howard et al. [86] [^34] https://github.com/pnnl/neuromancer/tree/feature/fbkans/examples/KANs, [^35] https://github.com/pnnl/neuromancer/blob/feature/fbkans/src/neuromancer/modules, extend KANs with a domain-decomposition strategy inspired by finite-basis PINNs (FBPINNs) [184, 185, 186, 187]. The key idea is to represent a global mapping as a sum of local subnetworks, each modulated by a smooth partition-of-unity (PoU) function. This yields scalability, robustness, and improved generalization for multiscale and physics-informed problems.
Let $\Omega\subset\mathbb{R}$ be covered by $L$ overlapping subdomains $\{\Omega_{j}\}_{j=1}^{L}$. Assign a smooth PoU weight $\omega_{j}(x)$ to each $\Omega_{j}$ such that [188, 189, 190]
$\mathrm{supp}\big(\omega_{j}\big)=\Omega_{j},\qquad\sum_{j=1}^{L}\omega_{j}(x)\equiv 1\ \ \forall x\in\Omega.$
Cosine windows are used to build unnormalized weights $\hat{\omega}_{j}$, which are then normalized:
$\omega_{j}(x)\;=\;\frac{\hat{\omega}_{j}(x)}{\sum_{k=1}^{L}\hat{\omega}_{k}(x)},$ (34)
with
$\hat{\omega}_{j}(x)\;=\;\begin{cases}1,&L=1,\\[4.0pt] \Big[1+\cos\!\big(\pi\tfrac{x-\mu_{j}}{\sigma_{j}}\big)\Big]^{2},&L>1,\end{cases}$ (35)
and
$\mu_{j}\;=\;\frac{l(j-1)}{L-1},\qquad\sigma_{j}\;=\;\frac{\delta\,l}{2(L-1)},\qquad l\;=\;\max(x)-\min(x),$
where $\delta>1$ is the overlap ratio controlling the window width.
Each subdomain $\Omega_{j}$ hosts a local KAN, $K_{j}(x;\bm{\theta}_{j})$, trained only on its region. The global predictor is a smooth PoU-weighted sum of local outputs:
$f(x)\;\approx\;\sum_{j=1}^{L}\omega_{j}(x)\,K_{j}(x;\bm{\theta}_{j}),$ (36)
where $\bm{\theta}_{j}$ are the local parameters. Within each local KAN, a univariate edge function $\varphi_{j}$ is expanded in cubic B-splines:
$\varphi_{j}(x)\;=\;\sum_{n=0}^{N_{j}-1}c_{j,n}\,B_{n}^{(3)}(x),$
with trainable coefficients $c_{j,n}\in\mathbb{R}$. The local spline grid is initialized from the effective support of $\omega_{j}$:
$a_{j}\;=\;\min\{x\mid\omega_{j}(x)>\varepsilon\},\quad b_{j}\;=\;\max\{x\mid\omega_{j}(x)>\varepsilon\},\quad\varepsilon\;=\;10^{-4}.$
To illustrate functional flexibility, Figure 13 shows, on [1,1][-1,1], (A) the PoU weights $\{\omega_{j}\}_{j=1}^{4}$ with $\sum_{j}\omega_{j}\equiv 1$, and (B) the PoU–weighted decomposition of a single cubic B-spline basis $B_{i}^{(3)}(x)$. The colored curves are the localized pieces $\omega_{j}(x)\,B_{i}^{(3)}(x)$; their sum coincides with the original basis up to numerical precision. Only windows overlapping the spline support contribute nonzero pieces, so the PoU localizes the representation without changing the underlying spline span.
Figure 13: Two-panel illustration of PoU–weighted B-spline decomposition on [1,1][-1,1] ($L=4$, overlap $\delta=1.5$). (A) Cosine-based PoU windows $\{\omega_{j}\}$ summing to one. (B) Decomposition of a cubic B-spline basis $B_{i}^{(3)}(x)$ into localized pieces $\omega_{j}(x)B_{i}^{(3)}(x)$; their sum (black) matches the original basis, confirming $B_{i}^{(3)}(x)=\sum_{j=1}^{L}\omega_{j}(x)\,B_{i}^{(3)}(x)$.
Summary. FBKANs are practical because they let you model different parts of the domain with specialized sub-KANs, which can be trained independently (and in parallel) to cut memory use and wall time. They work well with physics-informed losses, handle noisy data robustly, and maintain smooth transitions between subdomains without requiring hard boundary constraints. Moreover, because sub-KANs operate on smaller supports, each local kernel matrix is better conditioned—an important advantage for large-scale PDEs.
6.9 SincKAN
To model singularities, sharp gradients, and high-frequency transitions, Yu et al. [41] [^36] https://github.com/DUCH714/SincKAN introduced the Sinc-based KAN (SincKAN). The building block is the globally supported Sinc kernel
$\mathrm{Sinc}(x)=\frac{\sin x}{x},\qquad\mathrm{Sinc}(0):=1,$
a classical tool for bandlimited approximation.
A basic SincKAN activation expands a univariate function using uniformly shifted Sinc atoms:
$\varphi(x)=\sum_{i=-N}^{N}c_{i}\,\mathrm{Sinc}\!\left(\frac{\pi}{h}\,(x-ih)\right),$ (37)
where $h>0$ is the step size, $\{c_{i}\}$ are trainable c oefficients, and $N$ controls the truncation degree (total number of atoms $2N+1$).
Practical multi–step-size form. To avoid choosing a single “optimal” $h$ and to improve flexibility on finite domains, SincKAN mixes several step sizes and applies a normalized coordinate transform $\gamma^{-1}$:
$\varphi_{\mathrm{multi}}(x)=\sum_{j=1}^{M}\sum_{i=-N}^{N}c_{i,j}\,\mathrm{Sinc}\!\left(\frac{\pi}{h_{j}}\big(\gamma^{-1}(x)-ih_{j}\big)\right).$ (38)
Here $\{h_{j}\}_{j=1}^{M}$ are (learned or preset) step sizes, $M$ is the number of scales, $c_{i,j}$ are trainable coefficients, and $\gamma^{-1}$ maps the working interval to $\mathbb{R}$ while normalizing scale. Setting M=1M{=}1 and $\gamma^{-1}(x){=}x$ recovers (37).
Figures 14 and 15 show the building blocks and a synthesized activation using a fixed truncation degree $(2N+1=11)$ and spacing $h=1.0$ over $x\in[-10,10]$. Figure 14 displays the truncated Sinc dictionary $\{\mathrm{Sinc}(\tfrac{\pi}{h}(x-kh))\}_{k=-N}^{N}$, whose elements are globally supported and oscillatory with decaying sidelobes. These atoms are linearly combined to form the learned univariate map $\varphi(x)$.
Figure 14: Truncated Sinc basis functions used in SincKAN. Each kernel is centered at $x=kh$ with spacing $h=1.0$; the total number of shifts is $2N+1=11$. This dictionary underlies the learned univariate map $\varphi(x)$.
Figure 15 plots synchronized learned activations
$\varphi(x)=\sum_{i=-5}^{5}c_{i}\,\mathrm{Sinc}\!\big(\tfrac{\pi}{h}(x-ih)\big)$
with the same random coefficients $\{c_{i}\}$, while varying only the step size $h\in\{0.8,1.0,1.4\}$. The step size $h$ sets the spacing of Sinc centers and therefore the effective bandwidth: smaller $h$ $\Rightarrow$ denser centers, higher-frequency capacity, more oscillations; larger $h$ $\Rightarrow$ smoother, lower-frequency behavior. Because the coefficients are synchronized, differences across curves reflect only the choice of $h$.
Figure 15: Synchronized learned activations with a pure (non-windowed) Sinc basis (degree $2N+1=11$). The same coefficients $\{c_{i}\}$ are used for all curves; only the step size $h$ changes ($h\in\{0.8,1.0,1.4\}$). Smaller $h$ yields higher-frequency, more oscillatory structure, whereas larger $h$ produces smoother responses. Vertical dotted lines mark the Sinc centers for $h=1.0$.
For numerical stability, SincKAN applies a coordinate normalization step that maps inputs to a bounded interval (typically [1,1][-1,1]) before evaluating the Sinc expansion. This preprocessing, consistent with other KAN variants, improves conditioning and convergence across layers.
7 Accuracy Improvement
We group works by how they improve accuracy, independent of task (regression, PINN, DeepONet), basis choice, or efficiency tricks. In most cases the same mechanism transfers across tasks.
To orient the reader, we begin with a compact summary of accuracy–improving mechanisms for KANs (Table 6). The table groups methods by mechanism and lists concise technique keywords with representative references. The remainder of this section expands on each group in turn, detailing when/why they help and practical choices and caveats.
Table 6: Compact map of accuracy-improving mechanisms for KANs. Cells use concise keywords; see Section 7 for context.
Mechanism Technique keywords Ref.
Physics-consistent & loss design
PIKAN Chebyshev recursion; RBA; EVM; Bayesian [29]
KAN-PINN Strongly nonlinear PDE residuals; actuator deflection [47]
KAN-MHA Navier–Stokes residuals + BC losses; attention-guided [48]
Res-KAN Residual physics + sparse regression; variable-coefficient PDEs [49]
KKAN Self-scaled residual reweighting; adaptive residual focus [44]
AL-PKAN Augmented Lagrangian; auto weight-balancing [58]
AIVT (cKAN) Velocity–vorticity loss; turbulence; residual weighting [194]
KANtrol Fractional / integro-differential operators [59]
Adaptive sampling & grids
[…]
Free-knot KAN Trainable knot positions; cumulative softmax ordering [38]
Grid extension Increase $G$ during training; optimizer state transition [82]
[…]
Multi-resolution PIKAN Scheduled coarse/fine sampling for Chebyshev-based PIKAN [64]
Domain decomposition
FBKAN (PoU) Overlapping subdomains; smooth PoU blend; local KANs [86]
Temporal subdomains Long-horizon split; improved NTK conditioning [56]
Function decomposition
MFKAN Freeze LF surrogate; linear+nonlinear HF heads; blend $\alpha$ [162]
SPIKAN Sum of products of 1D KAN factors; tensorized structure [163]
KAN-SR Recursive simplification; divide–and–conquer symbolic extraction [204]
Hybrid / Ensemble & data
MLP–KAN MLP experts for features; KAN for functional shaping [33]
HPKM-PINN Parallel MLP & KAN branches; learnable fusion [50]
KKAN Per-dim MLP features; explicit basis expansion [44]
Sequence / attention hybrids
FlashKAT (Transformer) Group-rational KAN blocks inside Transformer [166]
[…]
GINN-KAN Growing interpretable NN + KAN; PINN coupling [165]
KAN-ODE KAN as $\dot{u}$ in Neural ODE; adjoint training [36]
AAKAN-WGAN Adaptive-activation KAN + WGAN; data augmentation [168]
Discontinuities & sharp gradients
SincKAN Sinc interpolation; kinks; boundary layers [41]
rKAN Rational bases; asymptotics; jump capture [92]
DKAN $\tanh$ gate (jump) + spline background; shock localization [52]
KINN Singularities / stress concentrations; reduced spectral bias [61]
Two-phase PINN–KAN Hybrid optimizer; preserves saturation fronts [53]
Optimization & adaptive training
[…]
Bayesian Bayesian optimization of KAN hyperparameters; imbalanced data [161]
Bayesian PINN–KAN Variational inference; KL regularization; calibrated uncertainty [105]
7.1 Physics Constraints & Loss Design
Physics-informed learning, introduced by Raissi et al. [1], improves accuracy by embedding governing physical laws—PDE residuals and boundary/initial conditions—directly into the training objective.
A standard PINN-style loss combines three components:
[…]
where $\mathcal{L}_{\mathrm{data}}$ is the data-fidelity loss, $\mathcal{L}_{\mathrm{phys}}$ enforces PDE residuals, and $\mathcal{L}_{\mathrm{bc}}$ enforces boundary/initial conditions. The scalar weights $w_{u},w_{f},w_{b}>0$ determine the balance between data fidelity and physics consistency.
Building on this foundation, a wide range of studies have extended the basic PINN loss through specialized constraints and tailored formulations in cooperation with KAN architectures [27, 41, 47, 48, 49, 50, 44, 54, 55, 170, 165, 60, 97, 105, 51, 52, 53, 138, 208]. Among these, several works move beyond the conventional physics-based loss in (39) by embedding more advanced physics constraints and loss designs directly into the KAN architecture:
  • Zhang et al. [47]: KAN-PINN, embedding strongly nonlinear PDE residuals directly into the loss, enabling accurate prediction of IPMC actuator deflections with improvements over MLP-PINNs.
  • Yang et al. [48]: KAN-MHA, integrating Navier–Stokes residuals and boundary-condition losses into the training objective to guide convergence in airfoil flow prediction.
  • Guo et al. [49]: Res-KAN, combining residual-based physics losses with sparse regression terms to recover variable-coefficient PDEs under sparse and noisy observations.
  • Shukla et al. [29]: cPIKAN + RBA + EVM, strengthening physics-informed KANs by incorporating residual-based attention and entropy-viscosity stabilization, improving robustness in high-Reynolds-number fluid simulations.
  • Toscano et al. [44]: KKAN + ssRBA, introducing self-scaled residual-based attention that adaptively reweights PDE residual losses to preserve signal quality and enhance generalization.
  • Zhang et al. [58]: AL-PKAN, replacing fixed penalty weights with learnable Lagrange multipliers through an augmented Lagrangian loss, allowing dynamic balancing among multiple physics constraints.
  • Toscano et al. [194]: AIVT, reformulating the physics loss in velocity–vorticity form to improve reconstruction of turbulent convection fields from sparse experimental inputs.
  • Aghaei [59]: KANtrol, extending the loss to fractional and integro-differential physics operators via Gaussian quadrature and operational matrices, enabling accurate handling of optimal control PDEs.
  • Chen et al. [208]: WCN, introducing a weak-form physics loss from the Fokker–Planck equation with adaptive collocation sampling, enabling efficient recovery of drift and diffusion terms in SDEs from aggregate data.
  • Gong et al. [209]: Physics-inspired KAN, embedding Navier–Stokes coupling into a multi-field K-means clustering step that constrains sensor placement and guides data-fidelity loss in indoor flow reconstruction.
  • Guo et al. [217]: PKAN, embedding prior physics into initialization and mapping spline activations to symbolic operators, enabling compact analytic equation discovery.
  • Sen et al. [212]: KAN-ETS, enforcing Ehrenfest constraints [213] in the loss to preserve quantum dynamical laws in time-series modeling.
  • Wang et al. [214]: SPEL-KAN, encoding Hamiltonian dynamics and Lie group symmetries [215] with physics-driven sparsity constraints on system matrices.
  • Lou et al. [227]: DAE-KAN, a physics-informed KAN that directly solves high-index differential–algebraic equations without index reduction, using coupled KAN branches to preserve algebraic constraints and eliminate drift-off errors.
In all these approaches, loss engineering—choosing residual forms, designing constraint terms, and applying adaptive weighting strategies—serves as a powerful inductive bias. By aligning the optimization process with the underlying physics, these designs improve accuracy, stability, and extrapolation capabilities across a broad range of PDE-driven problems. Beyond modifying the loss, several works instead adapt the sampling or grid itself, yielding comparable accuracy gains through geometric rather than variational control.
7.2 Adaptive Sampling & Grids
Adaptive sampling and grid refinement focus computational effort where the solution is most difficult to capture (steep gradients, shocks, oscillations). One simple approach is multi-resolution training, where a model is trained on alternating coarse and fine point clouds to accelerate convergence and lower computational cost [64]. More advanced methods, long used in finite elements [195, 196], adapt the model’s internal grid or sampling based on runtime dynamics for complex geometries [61, 52]. In stochastic systems, Chen et al. [208] introduced adaptive collocation sampling, selecting Gaussian test functions directly from observed data distributions.
Geometric/grid refinement with KANs. Actor et al. [38] treat spline knots as a learnable grid, which can be refined during training:
  1. 1.
    Multilevel refinement. Training begins on a coarse knot grid and periodically switches to finer grids. Since spline spaces are nested (coarse $\subset$ fine), the coarse solution transfers exactly to the refined grid. This preserves training progress, adds capacity only where needed, and achieves lower error at the same FLOPs compared with single-level training.
  2. 2.
    Free-knot adaptation. Knot locations are made trainable on $[a,b]$, with ordering enforced by a cumulative softmax:
    $t_{i}\;=\;\underbrace{a}_{\text{left endpoint}}\;+\;\underbrace{(b-a)}_{\text{interval length}}\sum_{j=1}^{i}\underbrace{\mathrm{softmax}(s)_{j}}_{\text{ordered weights}},\qquad s\in\mathbb{R}^{n},$
    ensuring that endpoints remain fixed ($t_{0}=a$, $t_{n}=b$). This naturally places knots in regions of high variation or non-smoothness.
On both regression and physics-informed tasks, multilevel schedules and free-knot adaptation consistently reduced errors relative to fixed grids and MLP baselines [38].
Grid–adaptive PIKANs. Rigas et al. [82] extend this idea by adapting not only the spline grid but also the collocation points, so resolution follows the PDE residual:
  1. 1.
    Grid extension with smooth state transition. Start with few spline intervals ($G$) and periodically increase $G$. At each extension, an adaptive optimizer state transition preserves momentum and avoids loss spikes, while the finer grid captures unresolved structure (Fig. 16(a)).
  2. 2.
    Residual-based adaptive resampling (RAD). Compute residuals on a dense probe set $S$ and convert them into sampling probabilities: larger residuals $\Rightarrow$ higher sampling chance. A new collocation set is then drawn accordingly (Fig. 16(b)).
Since KAN basis functions are tied to the knot grid, jointly adapting the grid and collocation points reduces projection error and concentrates model capacity in residual hotspots.
(a) Loss vs. epochs
(b) Residual-based adaptive sampling
Figure 16: Adaptive PIKAN training after [82]. (a) Loss vs. epochs: adaptive state transition smooths grid updates. (b) RAD concentrates collocation points where PDE residuals are largest.
7.3 Domain decomposition
Splitting a difficult global map into easier local ones and blending them smoothly often improves accuracy. For example, temporal subdomaining can sharpen NTK conditioning and convergence [56]. Finite-basis KANs [86], as discussed in Sec. 6.8, achieve this by decomposing $\Omega$ (or $\Omega\times[0,T]$) into $L$ overlapping subdomains $\{\Omega_{j}\}_{j=1}^{L}$ with a smooth partition of unity (PoU) $\{\omega_{j}\}_{j=1}^{L}$ such that
$\underbrace{\mathrm{supp}(\omega_{j})}_{\text{support of local weight}}=\Omega_{j},\qquad\underbrace{\sum_{j=1}^{L}\omega_{j}(x)\equiv 1}_{\text{partition of unity}}.$
A schematic is shown in Fig. 17.
A convenient 1D choice uses normalized “cosine bumps.” With domain length $l$ and overlap ratio $\delta>1$:
$\hat{\omega}_{j}(x)=\bigl[1+\cos\!\big(\pi(x-\mu_{j})/\sigma_{j}\big)\bigr]^{2},\quad\mu_{j}=\underbrace{\tfrac{l(j-1)}{L-1}}_{\text{center of bump}},\quad\sigma_{j}=\underbrace{\tfrac{\delta\,l/2}{L-1}}_{\text{width/overlap}},$
$\omega_{j}(x)=\underbrace{\frac{\hat{\omega}_{j}(x)}{\sum_{k=1}^{L}\hat{\omega}_{k}(x)}}_{\text{normalized local weight}}.$
In $d$ dimensions, the construction extends by taking products across coordinates.
The global surrogate is then the PoU blend of local KANs:
$f_{\Theta}(x)=\sum_{j=1}^{L}\underbrace{\omega_{j}(x)}_{\text{local weight}}\,\underbrace{K_{j}(x;\bm{\theta}_{j})}_{\text{local KAN}},$
where each $K_{j}$ is trained only on $\Omega_{j}$, using its own knot grid induced by $\omega_{j}$ (basis/grid choices per Sec. 6.8).
A basic error estimate illustrates the benefit:
$\|u-f_{\Theta}\|\;\leq\;\sum_{j=1}^{L}\underbrace{\|\omega_{j}\|_{\infty}}_{\text{PoU weight}}\,\underbrace{\|u-K_{j}\|_{\Omega_{j}}}_{\text{local error}},$
showing that reducing each local error reduces the global error. Increasing $L$ or the tunable overlap $\delta$ concentrates resolution where it is most needed. Each local KAN operates on a smaller, better-conditioned subproblem, allowing parallel training with improved numerical stability. In practice, per-subdomain grids adapt to local scales and consistently yield lower relative $L_{2}$ errors than single-domain KANs at the same model capacity [86].
Figure 17: Schematic of FBKAN domain decomposition and PoU blending. Overlapping subdomains $\Omega_{j}$ host lightweight local KANs $K_{j}$, smoothly combined via PoU weights $\omega_{j}(x)$. This enables parallel training and per-subdomain knot/hyperparameter choices [86].
7.4 Function decomposition
[…]
$K_{H}(x)=\underbrace{(1-\alpha)\,K_{\ell}\big(x,K_{L}(x)\big)}_{\text{linear LF}\to\text{HF trend}}+\underbrace{\alpha\,K_{nl}\big(x,K_{L}(x)\big)}_{\text{nonlinear HF correction}}\!,$
where $\alpha\in[0,1]$ is a trainable parameter, $K_{\ell}$ is the linear block, $K_{nl}$ the nonlinear block, and $K_{L}$ the frozen LF surrogate.
The loss function balances three objectives:
[…]
Here, $\lambda_{\alpha}>0$ and $n\in\mathbb{N}$ control how strongly the model prefers the linear block, while $w>0$ applies weight-decay on the nonlinear block parameters $\Phi_{nl}$.
[…]
In a related line of work, Jacob et al. [163] propose a separable PIKAN (“SPIKAN”), where instead of one multivariate KAN, the solution is expressed as a sum of products of 1D KAN factors (one per coordinate), which are then combined to approximate the global field.
[…]
A complementary direction is given by Bühler et al. [204], who introduce KAN-SR, a symbolic regression framework that follows the recursive simplification strategies of Udrescu et al. [203]. Instead of fitting one large model, the method first checks whether the target function can be simplified by identifying separability (splitting variables into independent parts) or symmetry (invariance under transformations such as shifts or scalings). If such structure is found, the problem is broken down into smaller subproblems that are easier to solve. Each subproblem is then modeled with a compact KAN, and the results are combined to recover closed-form symbolic expressions. This divide–and–conquer strategy makes the learning process more interpretable and effective, especially in noisy or high-dimensional settings.
7.5 Hybrid/Ensemble & Data
A parallel stream of research improves accuracy not by new bases but by architectural integration—combining KANs with other neural modules. KANs have also been integrated with other neural components to form hybrid or ensemble models. These architectures combine the representation power of conventional neural networks with the interpretability and structured basis of KANs, aiming for higher accuracy, better conditioning, and improved use of data and physics priors.
Two representative examples highlight this direction:
  • He et al. [33] propose an MLP–KAN mixture-of-experts, where MLP blocks focus on representation learning while KANs specialize in function shaping. This division of labor improves flexibility without losing interpretability.
  • Xu et al. [50] introduce HPKM-PINN, a parallel hybrid in which KAN and MLP branches run side-by-side. Their outputs are fused through a learnable weight, yielding more accurate PDE solutions than either branch alone.
A more detailed hybrid design is KKAN, introduced by Toscano et al. [44]. This architecture integrates per-input feature learning with an explicit basis expansion. Each input coordinate $x_{p}$ is first mapped by a small MLP (the inner block) to one-dimensional features $\Psi_{p,q}(x_{p})$. A combination layer then sums across input dimensions to form $\xi_{q}=\sum_{p=1}^{d}\Psi_{p,q}(x_{p})$. For each $\xi_{q}$, an outer block applies a linear basis expansion
$G_{q}(\xi_{q})=\sum_{j=0}^{D}C_{q,j}\,b_{q,j}(\xi_{q}),$
where $b_{q,j}$ are simple basis functions (Chebyshev or Legendre polynomials, sine series [21], or RBFs). A final combination layer sums the outputs of all channels to produce the network output:
$\underbrace{f(x_{1},\dots,x_{d})}_{\text{output}}=\sum_{q=0}^{m}\underbrace{G_{q}\!\Big(\underbrace{\sum_{p=1}^{d}\Psi_{p,q}(x_{p})}_{\text{inner combination }(\xi_{q})}\Big)}_{\text{outer basis on }\xi_{q}}.$
In compact terms, KKAN follows the flow: inputs $\rightarrow$ inner MLPs $\rightarrow$ dimensionwise combination $\rightarrow$ basis expansion $\rightarrow$ final sum (see Fig. 19). The inner MLPs learn local, dimension-specific features, while the outer basis expansion provides a low-bias, well-conditioned representation. Toscano et al. [44] report that KKAN achieves lower test errors than both vanilla KANs and standard MLPs across tasks including function approximation, operator learning, and PDE solving.
Figure 19: KKAN architecture (after [44]). Each input coordinate $x_{p}$ These are summed across dimensions to form $\xi_{m}$, which are expanded via basis functions in the outer block. A final summation yields the output. Bases may include Chebyshev/Legendre polynomials, sine series, or RBFs. This “MLP inside + basis outside” layout mirrors the Kolmogorov–Arnold recipe.
7.6 Sequence/attention hybrids
[…]
  • Yang et al. [48]: KAN-MHA, which embeds KAN blocks within multi-head attention layers in a physics-informed framework.
  • Raffel et al. [166]: FlashKAT, introducing Group-Rational KAN modules inside Transformer layers for efficient multi-scale representation.
  • Zhang et al. [58]: AL-PKAN, a GRU–KAN encoder–decoder model for sequence-to-field mappings.
  • Ranasinghe et al. [165]: GINN-KAN, combining Growing Interpretable Neural Networks with KANs for interpretable PINNs.
  • Koenig et al. [36]: KAN-ODE, coupling KAN with neural ODEs for modeling continuous-time dynamics.
  • Shen et al. [168]: AAKAN-WGAN, pairing adaptive-activation KAN with a Wasserstein GAN to learn effectively from limited data.
  • Wei et al. [180]: Attention-KAN-PINN, integrating an attention module with a standard B-spline KAN and a physics-informed loss for lithium-ion battery SOH forecasting on the MIT dataset.
These hybrids follow a common recipe: the auxiliary module supplies temporal memory, global context, or synthetic data, while the KAN component provides a compact, interpretable functional mapping at the prediction stage. We highlight one representative case below.
7.7 Discontinuities and Sharp Gradients
Several KAN variants have been developed to better capture discontinuities, shocks, and steep gradients—settings where smooth spline or polynomial bases often fail. Representative approaches include:
  • Ye et al. [41]: SincKAN, employing Sinc interpolation to naturally handle kinks, boundary layers, and singularities, improving accuracy in regions with steep gradients.
  • Yang et al. [48]: KAN–MHA, combining KAN with multi-head attention to focus model capacity on high-variation flow zones (e.g., leading/trailing edges, stall onset), sharpening predictions near gradient hotspots.
  • Kalesh et al. [53]: Two-phase flow (Buckley–Leverett), a physics-informed KAN with hybrid optimization that captures sharp saturation fronts while preserving steepness during propagation.
  • Shukla et al. [29]: cPIKAN + entropy viscosity/RBA, stabilizing training near shocks and improving robustness against oscillations.
  • Toscano et al. [194]: AIVT, a cKAN-based framework embedding velocity–vorticity physics with residual weighting, enabling reconstruction of steep turbulent gradients from sparse measurements.
  • Toscano et al. [44]: KKAN + ssRBA, introducing self-scaled residual attention with robust convergence across both smooth and discontinuous benchmarks.
  • Mostajeran et al. [54]: EPi–cKAN, encoding yield physics into Chebyshev-KAN layers to capture abrupt stress–strain regime transitions in elasto-plasticity.
  • Wang et al. [61]: KINN, designed for singularities and stress concentrations in solid mechanics, mitigating spectral bias and better resolving sharp gradients.
  • Aghaei [92]: rKAN, employing rational basis functions to approximate asymptotics and near-singular behavior, improving fits across sharp transitions.
Taken together, these works target discontinuities by either localizing the basis (e.g., Sinc, rational functions), steering attention to high-variation zones, or adapting the physics/loss terms to stabilize training.
Edge-aware KANs. Lei et al. [52] push this idea further by embedding an explicit edge activation inside KAN layers, allowing the network to place sharp jumps where shocks occur while keeping the representation smooth elsewhere. Instead of a purely smooth basis, DKAN uses a combination of a sharp gate and a smooth background:
$\psi(x)=\underbrace{w_{t}\,\tanh\!\big(\alpha(x-\beta)\big)}_{\text{jump term}}\;+\;\underbrace{w_{s}\sum_{i}c_{i}\,S_{i}(x)}_{\text{smooth background}},$
where $\alpha,\beta,w_{t},w_{s}\in\mathbb{R}$ and $c_{i}\in\mathbb{R}$ are trainable, while $S_{i}(\cdot)$ denotes B-spline basis functions.
Intuitively, the $\tanh$ gate pinpoints the shock location ($\beta$) and steepness ($\alpha$), while the spline term fits the smooth regions on either side. Embedding this mechanism inside a PIKAN preserves the standard physics loss but makes steep fronts and discontinuities much easier to represent, typically with fewer parameters and more stable training. Together, these formulations reduce spectral bias and improve stability in non-smooth regimes—a long-standing weakness of both MLPs and spline-based KANs.
7.8 Optimization & Adaptive Training
Optimizing the training of KAN-based architectures is critical for accuracy and stability, especially in physics-informed or data-scarce problems. Two main directions have emerged: (i) optimizer selection and hybrid strategies, and (ii) Bayesian and probabilistic approaches.
[…]
Bayesian and probabilistic optimization. Probabilistic approaches enhance KANs by explicitly modeling uncertainty and automating hyperparameter selection. Lin [161] employed Bayesian optimization to tune hyperparameters of RBF–KANs for geohazard prediction, achieving higher accuracy on imbalanced and discrete-valued datasets. Giroux and Fanelli [105] [^37] https://github.com/wmdataphys/Bayesian-HR-KAN extended Higher–Order ReLU–KANs with Bayesian inference, treating both weights and basis parameters as random variables optimized via variational inference. A KL-regularization term anchors them to prior distributions, stabilizing training and mitigating overfitting, which improved generalization in stochastic PDEs such as Poisson and Helmholtz equations with random forcing. Hassan [128] further advanced this direction by introducing a probabilistic spline–based Bayesian–KAN, in which the spline coefficients $c_{n}$ in Eq. (8) are modeled as Gaussian random variables $c_{n}\!\sim\!\mathcal{N}(\mu_{n},\sigma_{n}^{2})$. This formulation enables uncertainty–aware and interpretable predictions across both inner and outer KAN mappings, yielding higher accuracy and better–calibrated uncertainty estimates than deterministic KANs and conventional neural networks on multiple medical benchmarks, thereby underscoring the role of Bayesian reasoning in enhancing trustworthiness and reliability. While optimizer choice governs convergence behavior, computational efficiency also depends on architectural and matrix-level design choices, discussed next.
8 Efficiency Improvement
We group efficiency strategies by how they reduce wall–clock training time or computational cost. The emphasis is on the method and its design, rather than accuracy gains. Table 7 compactly summarizes KAN efficiency strategies—parallelism/GPU/JAX engineering and matrix/basis/parameter optimizations.
8.1 Parallelism, GPU, and JAX engineering
Several works exploit parallelism and GPU acceleration to speed up KAN training. At the activation level, iterative spline evaluations have been replaced with non-iterative ReLU-power functions [40] or reformulated as matrix operations optimized for CUDA [87, 97], enabling direct GPU parallelization and large speedups. Model restructuring has also been explored: Zhang et al. [63] merged the dual-matrix structure with trainable Random Fourier Features to scale more efficiently in high dimensions, while Raffel et al. [166] redesigned the backward pass with custom GPU kernels to cut memory stalls and accelerate training. Basis evaluations were likewise accelerated by matrix-based B-spline computations on GPUs [88]. Parallel execution appears at the architectural level too, with multi-branch layouts such as KAN–MLP hybrids [50] and domain decomposition approaches [29, 86, 163] that allow independent subproblems to be distributed across GPUs. At the framework level, explicit JAX-based implementations [57, 82] improved automatic differentiation and GPU-aware memory management, achieving substantial speedups in training.
8.2 Matrix optimization and parameter-efficient bases
Another stream of work reduces KAN complexity by optimizing matrix operations or employing lighter basis functions. Replacing computationally heavy B-splines with faster alternatives has been widely studied: ReLU-power formulations [40, 87, 97], orthogonal polynomials such as Chebyshev, Jacobi, and Legendre [29, 170, 54, 55, 61], and compact bases like RBFs [161, 36] or wavelets [164] all reduce runtime and parameter count while retaining approximation strength.
Efficiency also improves through matrix structuring and parameter compression. Zhang et al. [63] fused KAN’s dual matrices with Random Fourier Features for high-dimensional compression, while Guo et al. [49] applied sparsity-inducing regularization. Actor et al. [38] introduced hierarchical channel-wise refinement with trainable spline knots, and Li et al. [35] applied Differential Evolution to optimize layer connectivity. Together, these methods cut redundant parameters, streamline computations, and make KANs more scalable for large-scale problems.
Lee et al. [210] show that mixing spectral (for derivatives) and spatial (for position-dependent coefficients) representations keeps the core matvecs near-diagonal. Practically, this avoids dense mode mixing, so fewer Fourier modes and smaller GEMMs are needed on variable-coefficient PDEs. The result is lower memory traffic and shorter wall-clock time than pure-spectral operator layers, without changing the overall training loop.
Table 7: Compact map of efficiency-improving strategies for KANs. Cells use concise keywords; see Section 8 for context.
Mechanism Technique keywords Ref.
Parallelism, GPU, and JAX engineering
ReLU-power activations Replace iterative spline eval; pointwise $\mathrm{ReLU}^{m}$; CUDA-friendly [40]
[…]
Matrix B-splines on GPU Fused matrix B-spline evaluation; reduced launch overhead; high throughput [88]
Dual-matrix merge + trainable RFF Operator fusion; compressed dual matrices; RFF for high-dim scaling [63]
Custom GPU backward (FlashKAT) Fused/custom kernels; fewer memory stalls; faster attention-style blocks [166]
[…]
Domain decomposition parallelism Independent subproblems; multi-GPU distribution; PoU / separable factors [29, 86, 163]
JAX/XLA implementations jit/vmap/pmap; XLA fusion; GPU-aware memory mgmt [57, 82]
Matrix optimization & parameter-efficient bases
ReLU-power vs. B-splines Fewer params; no de Boor; vectorized piecewise polynomials [40, 87, 97]
Orthogonal polynomials Chebyshev/Jacobi/Legendre; cheap recurrences; stable eval [29, 170, 54, 55, 61]
Compact bases (RBF) Local Gaussians; small supports; fast inner products [161, 36]
Wavelets Multi-resolution; sparse coefficients; fast transforms [164]
Dual-matrix + RFF compression Merge/fuse matrices; low-rank Fourier features; reduced memory traffic [63]
Sparsity regularization $\ell_{1}$/group sparsity on channels/edges; pruning; lighter layers [49]
Hierarchical channel-wise refinement Progressive channels; trainable knots; parameter sharing [38]
DE-based connectivity (DEKAN) Differential Evolution search of layer wiring; compact topologies [35]
Dual-domain sparsity Mix spectral (derivatives) + spatial (coefficients); near-diagonal ops; fewer retained modes [210]
9 Sparsity & Regularization
Regularization improves the generalization and stability of KANs, especially in noisy or ill-posed settings. Current strategies fall into sparsity penalties, weight decay, task-specific regularizers, and implicit/dropout-style methods.
L1 sparsity with entropy balancing. Liu et al. [27] propose an $\ell_{1}$ penalty on each univariate activation $\varphi_{ij}$, averaged over inputs, and aggregated across a layer:
$\|\varphi_{ij}\|_{1}\coloneqq\tfrac{1}{N_{p}}\sum_{s=1}^{N_{p}}\big|\varphi_{ij}(x^{(s)})\big|,\qquad\|\Phi\|_{1}\coloneqq\sum_{i,j}\|\varphi_{ij}\|_{1},$ (40)
where $N_{p}$ is the number of probe samples. An entropy term prevents a single edge from dominating:
$S(\Phi)\coloneqq-\sum_{i,j}\frac{\|\varphi_{ij}\|_{1}}{\|\Phi\|_{1}}\log\!\Bigg(\frac{\|\varphi_{ij}\|_{1}}{\|\Phi\|_{1}}\Bigg).$ (41)
The full objective is
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{pred}}+\lambda\!\left(\mu_{1}\sum_{\ell}\|\Phi_{\ell}\|_{1}+\mu_{2}\sum_{\ell}S(\Phi_{\ell})\right),$ (42)
with task-dependent coefficients $\lambda,\mu_{1},\mu_{2}$. EfficientKAN [81] [^38] https://github.com/Blealtan/efficient-kan and DeepKAN [79] [^39] https://github.com/sidhu2690/Deep-KAN simplify this by applying the $\ell_{1}$ penalty directly to weights. The same $\ell_{1}+$ entropy recipe is used in symbolic regression [45] and PDE learning [49], with Guo et al. adding a smoothness penalty to suppress noise. For Neural ODEs, Koenig et al. [36] apply a layerwise $\ell_{1}$ penalty followed by pruning:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{fit}}+\lambda_{\mathrm{L1}}\|\theta\|_{1},$ (43)
removing redundant parameters without accuracy loss.
Bühler et al. [204] extend these ideas in their symbolic regression framework (KAN-SR). Because each KAN-SR unit operates on the full input space, the model tends to be overspecified. To enforce compactness and interpretability, they introduce a composite regularizer combining three terms: (i) a magnitude penalty that reduces unnecessary input usage, (ii) row- and column-wise entropy losses that encourage each subunit to focus on a small subset of variables, and (iii) an $\ell_{1}$ penalty on base linear weights. This balances accuracy with sparse structure, aligning learned representations with the compact operator forms common in symbolic regression tasks.
Beyond deterministic penalties, Zou and Yan [216] propose Probabilistic KANs (PKAN) with Gaussian process priors, and a sparse variant (SSKAN) using spike–and–slab priors. This Bayesian approach enforces sparsity while also providing uncertainty quantification, offering a principled alternative to $\ell_{1}$ pruning.
L2 weight decay and extensions. Shen et al. [168] introduce AAKAN with a base loss including an $\ell_{2}$ penalty
$\mathcal{L}_{\mathrm{base}}=\mathcal{L}_{\mathrm{pred}}+\lambda_{1}\mathcal{L}_{\mathrm{L2}},$ (44)
augmented by temporal smoothing and mutual-information terms
$\mathcal{L}_{\mathrm{AAKAN}}=\mathcal{L}_{\mathrm{base}}+\lambda_{2}\mathcal{L}_{\mathrm{time}}+\lambda_{3}\mathcal{L}_{\mathrm{MI}},$ (45)
which promote stable dynamics and decorrelated features. In physics-informed studies, small $\ell_{2}$ weight decay is also used (e.g., $\lambda_{\ell_{2}}=10^{-5}$ in DeepOKAN and thermal PINNs [29, 194]) to reduce overfitting while keeping optimization smooth.
Implicit and dropout-style regularizers. Some approaches achieve regularization without explicit penalties. Daryakenari et al. [57] observe that nested activations such as $\tanh(\tanh(x))$ implicitly bound outputs to [0.76, 0.76][-0.76,\,0.76], smoothing gradients and stabilizing Chebyshev-based training. Altarabichi et al. [197] propose DropKAN, which randomly masks post-activation outputs instead of inputs, injecting noise directly into the spline evaluation. This yields more stable regularization than standard dropout in KAN layers.
Table 8: Compact map of sparsity and regularization strategies for KANs. Cells use concise keywords; see Section 9 for context.
Mechanism/Name Technique keywords Ref.
$\ell_{1}$ sparsity with entropy balancing
Layerwise $\ell_{1}$ + entropy Edge-activation $\ell_{1}$ (probe-avg); layer aggregation; entropy balance $S(\Phi)$ [27]
EfficientKAN Apply $\ell_{1}$ directly to weights; simplified sparsity; implementation-friendly [81]
Symbolic regression Same $\ell_{1}$+entropy recipe for sparse symbolic discovery [45]
PDE KAN $\ell_{1}$ + smoothness penalty to suppress noise; variable-coefficient PDEs [49]
Pruning Layerwise $\ell_{1}$; post-training pruning; parameter removal w/o accuracy loss [36]
KAN-SR Magnitude + entropy at subunit level; compact operator usage; $\ell_{1}$ on base weights [204]
$\ell_{2}$ weight decay and extensions
AAKAN Base loss with $\ell_{2}$; + temporal smoothing $\mathcal{L}_{\mathrm{time}}$; + MI regularizer $\mathcal{L}_{\mathrm{MI}}$ [168]
Weight decay Small $\ell_{2}$ (e.g., $10^{-5}$) in DeepOKAN / thermal PINNs; smoother optimization [29, 194]
Implicit and dropout-style regularizers
Nested activations Implicit bounding via $\tanh(\tanh(\cdot))$; outputs in [0.76, $\approx[-0.76$,\,0.76]; gradient smoothing [57]
DropKAN Post-activation masking (after spline eval), not input; injected noise [197]
10 Convergence and Scaling Laws
Classical universal approximation theorems ensure that neural networks can approximate any continuous function $f:[0,1]^{n}\!\to\!\mathbb{R}$ to arbitrary precision, but they say little about how fast the error decreases with depth, width, or training time [160]. Recent works have filled this gap for KANs by establishing Sobolev-space approximation rates, analyzing gradient-flow convergence via Neural Tangent Kernels (NTK) [198, 8], and documenting empirical power laws observed across optimization trajectories.
10.1 Theoretical Approximation Rates and Empirical Scaling Laws
Sobolev rates and superconvergence. Wang et al. [45] established formal approximation bounds showing that deep KANs can achieve faster convergence than classical neural or kernel models. Consider a smooth target function $f:\Omega\subset\mathbb{R}^{d}\to\mathbb{R}$ belonging to a Sobolev space $W^{s}(\Omega)$ with smoothness order $s>0$. They proved that there exists a KAN $g$ with finite width, depth $L$, and spline order $k$ such that the approximation error satisfies
$\|f-g\|_{L^{p}(\Omega)}\leq C\,L^{-2s/d},$
where $C$ is a constant independent of depth. This result implies a form of superconvergence: the error decreases twice as fast as in standard nonlinear approximation [199, 200]. In terms of the total number of trainable parameters $P$, the error scales as $\mathcal{O}(P^{-2s/d})$, whereas conventional networks or kernel methods typically achieve only $\mathcal{O}(P^{-s/d})$. These rates hold under certain smoothness and regularity assumptions detailed in [45].
Empirical scaling laws. Liu et al. [27] further observed that, in practice, the root-mean-square error (RMSE) $\ell$ of KANs follows a power-law decay with respect to model size $N$:
$\ell\propto N^{-\alpha},$
where the exponent $\alpha>0$ depends on the specific task and data complexity. Although $\alpha$ may vary, KANs consistently achieve lower error than MLPs for the same parameter count, indicating more efficient use of model capacity. Notably, increasing the resolution of the basis grid (number of spline segments $G$) improves accuracy more effectively than merely widening the network, in agreement with the theoretical depth-dependent rate predicted by [45]. Together, these empirical scaling laws and theoretical Sobolev rates present a consistent picture of KANs as hierarchically efficient approximators capable of rapid convergence across smooth function classes.
10.2 Spectral Bias and Frequency Learning Behavior
Let $\widehat{f}(\bm{\xi})$ denote the Fourier transform of $f$. A model exhibits spectral bias if, during gradient-based training, low-frequency components (small $|\bm{\xi}|$) are fit earlier/faster than high-frequency components (large $|\bm{\xi}|$[220, 221, 224]. In NTK terms, frequency content aligns with eigenfunctions of $\mathbf{K}$ on the sampled domain; flatter NTK spectra (larger $\lambda_{\min}$ and slower eigenvalue decay) correspond to more uniform learning across frequencies.
Wang [45] provides both theoretical and empirical evidence for this reduced bias. Through eigenvalue analysis of the KAN Hessian, they show that shallow KANs possess a well-conditioned optimization landscape, allowing gradient descent to advance all frequency components at nearly uniform rates. Their experiments demonstrate that KANs can learn oscillatory waveforms within a few iterations, while MLPs converge slowly toward high-frequency components. They conclude that this property gives KANs a distinct advantage in scientific computing tasks requiring the resolution of fine spatial or temporal details. Furthermore, grid extension (refining spline knots) expands the range of representable frequencies and accelerates oscillatory-mode learning [45], consistent with NTK diagnostics in [31] showing similar spectral widening when the basis richness increases.
Farea [31] independently confirmed this observation in the physics-informed setting. Using NTK analysis across multiple PDE types, they found that KANs and learnable-basis PINNs (e.g., Fourier- or Jacobi-based) maintain flatter NTK eigenvalue spectra, indicating a weaker spectral bias and enhanced high-frequency representation [225, 226]. However, their results revealed an important trade-off: broader eigenvalue spectra correlate with larger Hessian eigenvalues (reflecting higher curvature and poorer conditioning of the optimization landscape), and consequently, greater training instability. Crucially, Farea showed that minimizing spectral bias alone does not guarantee higher accuracy, as stability and expressivity must be co-optimized [223]. Similarly, Wang [45] noted that excessive grid refinement can amplify sensitivity to noise and overfitting, whereas modest grid coarsening can reintroduce mild bias and improve generalization.
Taken together, both studies firmly establish that KANs are inherently less prone to spectral bias than MLPs or standard PINNs, enabling faster and more complete learning of high-frequency solution features [45, 31]. Yet, this benefit is not without cost. The same mechanism that reduces spectral bias also amplifies sensitivity in the loss landscape (a kernel–theoretic analogue of the uncertainty principle [171, 172]), making optimization more fragile. Hence, while KANs offer superior representational capacity for oscillatory and high-frequency PDEs, their success ultimately depends on carefully matching the basis to the underlying physics—locally supported spline bases for boundary layers and localized oscillations, versus global Fourier-like bases for broad spectral coverage—together with stability-aware training schedules that manage curvature and convergence sharpness [45, 31, 223, 225, 226].
10.3 NTK-Based Convergence and Physics-Informed Extensions
NTK convergence under gradient flow. Gao et al. [138] showed that the training dynamics of a two-layer KAN with learnable basis functions can be characterized by a gradient flow equation in the Neural Tangent Kernel (NTK) framework. For a standard squared loss, the residual vector $s(t)$ evolves according to
$\frac{d}{dt}s(t)=-\,G\,s(t),$
where $G=D^{\top}D$ is the NTK Gram matrix, with $D$ the Jacobian of network outputs with respect to parameters. If $G$ is positive definite ($G\succ 0$), the loss decreases exponentially:
$\mathcal{L}(t)\;\leq\;\Big(1-\eta\,\tfrac{\lambda_{\min}}{2}\Big)^{t}\,\mathcal{L}(0),$
where $\lambda_{\min}$ denotes the smallest eigenvalue of $G$ and $\eta$ the learning rate. This establishes a rigorous convergence guarantee under suitable step-size conditions and extends to stochastic gradient descent in expectation [138].
Physics-informed settings and spectral conditioning. When the objective includes PDE residual terms, a similar exponential decay law holds for the physics-informed loss [138, 56]:
$\mathcal{L}_{\mathrm{PDE}}(t)\;\leq\;\Big(1-\eta\,\tfrac{\lambda_{\min}}{2}\Big)^{t}\,\mathcal{L}_{\mathrm{PDE}}(0).$
In this case, however, the conditioning of the NTK matrix—and thus the convergence rate governed by $\lambda_{\min}$—is strongly problem-dependent. Stiff differential operators can significantly reduce $\lambda_{\min}$, slowing convergence unless the model is widened, the basis functions are better preconditioned, or domain decomposition is applied [56]. Faroughi et al. demonstrated that Chebyshev-based KANs (cKANs) and their physics-informed variants (cPIKANs) achieve faster, more stable convergence when their NTK spectra are well structured, particularly on diffusion, Helmholtz, Allen–Cahn, and beam problems. They further showed that temporal decomposition and hybrid first–second order optimization strategies improve NTK conditioning and accelerate training, in agreement with curriculum-based and multi-stage training results in [55, 54].
10.4 Practical Trade-offs and Convergence Guarantees
Under the NTK framework, convergence speed depends critically on the conditioning of the Gram matrix $G$, which in turn is influenced by the choice of basis function, spectral bias, and PDE complexity. KANs’ structured basis expansions often yield larger $\lambda_{\min}$ and smoother optimization trajectories than MLPs, resulting in faster error decay and improved training efficiency [138, 45]. However, the same spectral properties that enhance expressivity can also increase curvature and sensitivity in the loss landscape, leading to an inherent trade-off between representation power and stability [31]. This balance must be tuned through grid resolution, basis selection, and learning-rate scheduling. In physics-informed settings, Chebyshev-based KANs (cKANs) and their variants (cPIKANs) demonstrate that well-conditioned NTK spectra correlate with rapid and stable convergence, particularly when combined with temporal decomposition and hybrid optimization strategies [56, 55, 54]. At the same time, computational cost can escalate in high-dimensional or recursively refined models [34]. Overall, NTK-based convergence guarantees [138], Sobolev approximation rates [45], and empirical scaling laws [27] together form a consistent theoretical foundation for KANs. Yet, their practical success ultimately hinges on managing the expressivity–stability trade-off—achieving sufficient spectral reach without compromising conditioning or convergence reliability [45, 31, 56].
11 Practical “Choose–Your–KAN” Guide
KANs form a flexible toolbox. The right choice depends on (i) your target function (smooth/oscillatory, discontinuous, multi-scale, periodic), (ii) constraints (time, memory, precision), and (iii) goals (accuracy, interpretability). Below is a concise, step-by-step recipe followed by two compact tables you can use as a checklist. The guidance synthesizes Sections 610.
Step 1 — Start from a robust default. If you are new to KANs or tackling standard regression/classification, begin with a cubic B-spline KAN on a uniform grid. It offers a stable balance of expressivity, locality, and conditioning, with mature implementations [27, 77]. For boundary-sensitive tasks (e.g., PDEs), enable grid extension and consider post-training grid refinement to sharpen boundary fidelity (Sec. 6.1; Fig. 1; [77]).
Step 2 — Match the basis to function structure. Pick the basis by anticipated behavior (Table 5):
  • Globally smooth/oscillatory: Chebyshev (ChebyKAN) for strong spectral approximation; use per-layer $\tanh$ to keep arguments in [1,1][-1,1] and improve conditioning [169, 94, 41, 56].
  • Tunable smoothness/warping or boundary control: Jacobi (fKAN, Jacobi-KAN) with fractional/domain parameters for adaptable spectra [91, 62].
  • Discontinuities, kinks, boundary layers: SincKAN (bandlimited atoms) [41], rational rKAN (Padé/Jacobi) [92], or DKAN (explicit $\tanh$ jump + smooth background) [52].
  • Periodic targets: FourierKAN or KAF (learnable RFF) [98, 63].
  • Multi-scale/local bursts: Wavelets (Wav-KAN) with scale/shift parameters [103, 164].
Step 3 — Choose for efficiency if wall-clock matters. If speed is critical, swap splines for GPU-friendly bases and parallel layouts:
  • Fast local bases: ReLU-KAN/HRKAN replace de Boor evaluation by $\mathrm{ReLU}^{m}$ compositions (often $5\!-\!20\times$ faster); raise $m$ when higher derivatives are needed [87, 97]. Gaussian RBFs (FastKAN) are also light and smooth [83].
  • Parallelism and compilers: JAX jit/XLA, fused CUDA kernels, and multi-branch/domain-decomposition designs bring large speedups [82, 57, 86].
  • Parameter economy: Prefer orthogonal polynomials or RBFs when they reduce FLOPs/params without harming accuracy; apply sparsity/entropy regularizers to prune [29, 49, 27].
Step 4 — Set grids/normalization to stabilize training. Keep basis arguments in well-conditioned ranges: $\tanh$ for Chebyshev/Jacobi; min-max or sigmoid/tanh normalization for RBF/wavelet; extended or free-knot grids for splines when boundaries or hotspots matter [38, 82].
Step 5 — Use physics and sampling where it counts. For PINNs/PIKANs, compose
$\mathcal{L}=w_{u}\mathcal{L}_{\mathrm{data}}+w_{f}\mathcal{L}_{\mathrm{phys}}+w_{b}\mathcal{L}_{\mathrm{bc}}$
[…]
Step 6 — Use proven optimizer schedules. Warm up with Adam/RAdam, then refine with (L-)BFGS; enable mixed precision and gradient clipping if needed [57, 55, 39, 56]. This schedule routinely improves convergence on stiff PDEs and reduces training time.
Step 7 — Scale capacity deliberately. KAN error typically follows $\ell\propto P^{-\alpha}$ with task-dependent $\alpha>0$; increasing basis resolution (e.g., spline $G$) often yields larger gains than merely widening layers (Sec. 10; [27, 45]). Favor targeted capacity: local refinement (free knots), subdomain PoU (FBKAN), or basis enrichment rather than uniform growth [38, 86].
The next two tables operationalize these steps.
A. Decision table: pick a KAN by problem traits from Table 9
Table 9: Heuristics for choosing KAN variants and key knobs from regression to PDEs.
Problem trait Recommended KAN/basis Why & key knobs
Smooth, periodic fields FourierKAN or ChebyKAN [98, 29, 63] Low-bias spectral bases; start K=1632K{=}16\!-\!32, modest $G$. Enforce periodic BCs in $\mathcal{L}_{\mathrm{bc}}$ [1].
Nonperiodic but globally smooth Cubic B-spline KAN (uniform grid) [27] Stable on bounded domains; start G=1632G{=}16\!-\!32; enable grid extension near boundaries [77].
Piecewise smooth, shocks, discontinuities SincKAN / rKAN / DKAN [41, 92, 52] Localize jumps (Sinc/rational) or add $\tanh$ gate for steep fronts; pair with RAD sampling [82]. Free knots recommended [38].
[…]
Complex geometry or heterogeneous domains Finite-Basis KAN (FBKAN) with PoU $\{\omega_{j}\}$ [86] Train local $K_{j}$ per $\Omega_{j}$; choose per-subdomain $G$, basis, and losses; parallelizable.
Stiff/high-order PDE operators ChebyKAN / Jacobi-KAN / HRKAN [29, 59, 97] Better conditioning for derivatives; consider residual-based attention or augmented Lagrangian [44, 58].
Operator learning with limited data DeepONet-style KANs (Gaussian/Cheby) [46, 29] Branch/trunk with localized bases; optionally include physics residuals in $\mathcal{L}$.
Time series / latent dynamics KAN-ODE (KAN as $\dot{u}$[36] Regularize with solver rollouts; shallow KAN layers with residual connections.
[…]
Symbolic discovery / sparsity need KAN with $\ell_{1}+$entropy on edges [27, 45] Promote parsimonious maps; prune after training; keeps formulas compact.
Resource-constrained inference RBF-KAN with fixed centers [46] Shallow, small $K$, quantization-friendly via localized Gaussians.
B. Training playbook: defaults and when to deviate from Table 10
Table 10: Compact training playbook with sensible defaults.
Component Default recipe Deviate when
Loss $\mathcal{L}=w_{u}\mathcal{L}_{\mathrm{data}}+w_{f}\mathcal{L}_{\mathrm{phys}}+w_{b}\mathcal{L}_{\mathrm{bc}}$ Start $w_{u}:w_{f}:w_{b}=1:1:1$ [1]; switch to augmented Lagrangian [58] or self-scaled RBA [44] if constraints conflict. Stiff operators/BCs dominate $\Rightarrow$ use learnable multipliers [58]; add entropy-viscosity for shocks [193, 192].
Sampling / collocation Uniform + RAD: $p(x)\propto r(x)^{\gamma},\ \gamma\in[1,2]$; refresh each epoch; dense probe set $S$ [82]. Residuals cluster $\Rightarrow$ cap per-cell samples; stratify near boundaries [82].
Grid / knots $(G,$ free knots)) Multilevel $G$ (e.g., $8\to 16\to 32$) with state-preserving optimizer transition [82]; enable free knots after warmup [38]. Loss spikes at grid change $\Rightarrow$ shorten step, carry optimizer moments [82]; increase PoU overlap in FBKAN [86].
Optimizer schedule Adam/RAdam warmup $\to$ (L-)BFGS refine; mixed precision OK [57, 55, 39, 56]. Ill-conditioning persists $\Rightarrow$ reduce LR, add clipping; switch earlier to second-order [55, 56].
Regularization Small $\ell_{2}$ ($\lambda_{\ell_{2}}\sim 10^{-6}\!-\!10^{-4}$[29, 194]; optional $\ell_{1}$ on edge activations ++ entropy [27]. Noisy labels / discovery $\Rightarrow$ stronger $\ell_{1}$, prune [27, 45]; temporal models $\Rightarrow$ add $\mathcal{L}_{\mathrm{time}}$ [168].
[…]
Early stopping / metrics Track relative $L_{2}$, residual norms, BC violation; keep best-on-physics checkpoint [1]. Operator learning $\Rightarrow$ also test out-of-grid generalization on unseen queries [46].
12 Current Gaps and Path Forward
The rapid proliferation of KANs, marked by an explosion of publications and open-source implementations (Tables 2 and 3), signals a field rich with opportunity yet characterized by fragmentation. While empirical studies have demonstrated KANs’ potential across regression, operator learning, and PDE solving, their comparative evaluations remain inconsistent. The field now stands at a crossroads: rather than engaging in superficial “KAN vs. MLP” contests, the next phase must focus on building a rigorous, principled science of KAN design, training, and theory.
The Limits of Current Comparative Frameworks
Most current KAN literature follows one of two dominant paradigms, both of which—though valuable—have clear limitations:
(1) Superficial “KAN vs. MLP” Benchmarks.
Many studies pit a single KAN variant against a baseline MLP with matched parameter counts. However, this comparison is often ill-defined: KANs and MLPs differ in the functional roles of parameters, the implicit regularization introduced by basis structures, and the stability constraints imposed by normalization or grid design. Consequently, results vary widely, from strong KAN outperformance to near equivalence [28, 29, 30]. Without a shared theoretical grounding, such results offer limited insight into why or when a KAN excels.
(2) Problem-Dependent “Basis-vs.-Basis” Races.
Another thread compares various bases—splines, Chebyshev polynomials, Gaussians—across selected problems [31]. While these studies show that performance is problem-dependent, they risk obscuring deeper principles. Each basis family spans a vast design space of normalization schemes, grid choices, and regularization methods. A single benchmark cannot meaningfully represent such diversity.
In sum, existing comparisons—whether architecture-based or basis-based—provide no predictive framework for design. A KAN is not a monolithic model; it is a framework whose performance depends on principled design choices rooted in numerical analysis, approximation theory, and optimization geometry.
Case Study: The Many Faces of a Chebyshev KAN
The “Chebyshev KAN” exemplifies both the pitfalls of naïve design and the power of principled refinement. A straightforward Chebyshev expansion often exhibits unstable training and inconsistent results. However, as multiple studies demonstrate, careful modifications dramatically improve stability and expressivity:
  • Domain stabilization: Applying tanh normalization at the input and between layers constrains arguments to [1,1][-1,1], where Chebyshev polynomials are numerically stable (Section 6.2).
  • Architectural hybrids: For inverse and PDE problems, replacing the final Chebyshev expansion with a linear head reduces overfitting and improves conditioning [57].
  • Grid-informed adaptivity: Introducing a small, learnable tanh grid before expansion, as in [44], provides local adaptivity and stabilizes training.
  • Principled optimization: Hybrid Adam–LBFGS schedules and modest weight decay help navigate sharp spectral loss landscapes [56, 55].
Once these refinements are incorporated, Chebyshev-based KANs become powerful and interpretable approximators. Similar trends hold for other bases: Gaussian KANs improve via adaptive centers [46]; B-spline KANs through free-knot adaptation [77, 38]; and finite-basis KANs through domain decomposition [86].
A Multi-Pillar Framework for the Next Generation of KAN Research
We propose a unified, multi-stage research agenda structured around six complementary pillars. Together, they chart a roadmap from empirical engineering to a comprehensive theoretical science of KANs.
Pillar 1: Building a Robust Component Library.
The immediate task is to consolidate best practices into a library of well-engineered, stable KAN layers. Each major basis family (Chebyshev, spline, Gaussian, Fourier, wavelet) requires a canonical, reproducible implementation with optimized normalization, initialization, and regularization. This engineering groundwork establishes reliable building blocks—such as StableChebyKANLayer or AdaptiveSplineKANLayer—for systematic study and fair benchmarking.
Pillar 2: Developing a Science of Basis Selection.
Basis choice is the defining inductive bias of a KAN, yet current guidelines are purely heuristic. A formal basis selection theory should link the mathematical properties of the target problem to the optimal basis type. Key open questions include:
  • How do spectral properties of a basis (global vs. local support) align with the Fourier spectrum of the data or PDE operator?
  • Can we prove that Chebyshev bases improve conditioning of the NTK for elliptic PDEs?
  • How can we quantify the trade-off between approximation power and implicit regularization (e.g., smooth B-splines vs. oscillatory polynomials)?
This theoretical bridge would replace trial-and-error tuning with principled, problem-dependent design.
Pillar 3: Toward a Comprehensive Optimization Theory Beyond the NTK Regime.
Existing convergence guarantees for KANs mirror those of MLPs under the Neural Tangent Kernel (NTK) framework [138, 56]. While powerful, NTK theory relies on assumptions that do not hold in practice—namely infinite width and “lazy training” near initialization. Real KANs are finite and undergo significant parameter movement, particularly in their learnable basis parameters (e.g., knot locations, RBF centers). A new optimization theory must therefore:
  1. 1.
    Describe the geometry of the loss landscape for finite-width KANs.
  2. 2.
    Characterize conditions that prevent convergence to poor local minima.
  3. 3.
    Analyze how adaptive basis parameters reshape this landscape during feature learning.
Such a theory would bridge the gap between NTK-based asymptotics and the empirical reality of nonlinear, feature-adaptive training.
Pillar 4: Understanding Generalization and Regularization.
KANs blur the boundary between architecture and regularization: the choice of basis implicitly constrains the function class. A low-degree polynomial basis naturally suppresses high-frequency noise, while a fine B-spline grid enables high-resolution fitting. Future work must quantify how this implicit bias interacts with explicit regularizers (e.g., L1, L2, entropy). Developing generalization bounds that depend jointly on the number of parameters and basis complexity (e.g., polynomial degree or grid resolution) would provide a rigorous explanation for KANs’ empirically strong generalization relative to MLPs of comparable size.
Pillar 5: A Formal Theory of Composition and Hierarchy.
While the Kolmogorov–Arnold theorem describes a two-layer representation, modern KANs are deep. We lack a theory describing how functional properties evolve under composition. Open questions include:
  • What is the functional meaning of “depth” in a KAN?
  • How do layers with heterogeneous bases (e.g., Fourier followed by spline) interact?
  • Under what conditions does composition enhance expressivity without destabilizing optimization?
A compositional function theory for deep KANs would formalize how information is processed in function space, providing principles for designing heterogeneous and hierarchical architectures.
Pillar 6: A Theory of Interpretability and Identifiability.
KANs are prized for interpretability: their univariate activations often resemble symbolic expressions like $\sin(x)$ or $x^{2}$. Yet this interpretability is empirical, not theoretical. Two different sets of learned functions can yield the same global mapping, raising issues of non-uniqueness and identifiability. A rigorous interpretability theory should:
  • Define when a learned function is a true, unique representation of the underlying law.
  • Provide conditions for stability—ensuring that small data perturbations do not radically alter learned symbolic forms.
  • Establish identifiability criteria guaranteeing that interpretable decompositions correspond to genuine structures.
This would ground symbolic interpretability in mathematical guarantees rather than anecdotal observation.
Outlook
The path forward for KANs is not a race to find a single dominant architecture. It is a structured progression: from building reliable components, to uncovering theoretical principles, to integrating them into interpretable, adaptive, and compositional systems. By developing robust engineering standards, principled basis theory, finite-width optimization frameworks, and formal interpretability guarantees, the KAN community can move beyond benchmarking toward a predictive, unified science of functional learning.
13 Conclusion
Kolmogorov–Arnold Networks have emerged as a compelling architectural paradigm, shifting the locus of learnable parameters from linear weights to univariate basis functions on the network’s edges. This review has synthesized the rapidly expanding KAN landscape, with a focus on scientific machine learning applications such as function approximation, regression, and PDE solving. Our analysis shows that KANs, while not a universal replacement for MLPs, offer a powerful and highly configurable alternative. The central theme of our findings is that the question is not whether KANs are better than MLPs, but rather which KAN is best suited for a given problem.
Across the literature, a consistent pattern emerges: well-chosen KAN variants frequently outperform vanilla MLP and PINN baselines in accuracy and convergence speed, though often at the cost of higher computational overhead per epoch. The key to unlocking this performance lies in navigating the vast design space that KANs open up. We have systematically organized this space, covering the diverse families of basis functions—from the locality of B-splines and Gaussians to the global spectral properties of Chebyshev and Fourier series—and detailing how choices in normalization, grid structure, and initialization critically impact stability and expressivity.
Furthermore, we have structured the ecosystem of techniques for enhancing KAN performance. For accuracy, strategies such as physics-informed losses, adaptive sampling, domain decomposition, and hybrid architectures have proven effective. For efficiency, innovations in GPU-aware engineering, matrix optimizations, and the use of computationally cheaper bases like ReLU-powers are mitigating KANs’ primary limitation. These advancements, coupled with principled regularization and a growing body of theoretical results on approximation rates and convergence, signal a maturing field.
The practical “Choose–Your–KAN” guide presented in Section 11 operationalizes these insights, providing a decision-making framework that maps problem characteristics to specific architectural and training choices. As we have argued, the true strength of the KAN framework lies not in a single canonical implementation, but in its modularity. By providing the tools to select, combine, and adapt basis functions, KANs empower researchers to inject valuable inductive biases tailored to the problem at hand, bridging the gap between versatile deep learning models and classical approximation theory. The journey ahead will involve deeper, basis-specific studies and standardized benchmarking, but it is clear that KANs have already secured their place as a vital and promising branch of neural network design.
14 Acknowledgments
This work was supported by the General Research Fund (GRF No. 12301824, 12300922) of Hong Kong Research Grant Council.