### v4

# A Practitioner’s Guide to Kolmogorov–Arnold Networks

Abstract
Kolmogorov–Arnold Networks (KANs), whose design is inspired—rather than dictated— by the Kolmogorov superposition theorem, have emerged as a structured alternative to MLPs. This review provides a systematic and comprehensive overview of the rapidly expanding KAN literature. The review is organized around three core themes: (i) clarifying the relationships between KANs and Kolmogorov superposition theory (KST), MLPs, and classical kernel methods; (ii) analyzing basis functions as a central design axis; and (iii) summarizing recent advances in accuracy, efficiency, regularization, and convergence.
Finally, we provide a practical “Choose–Your–KAN” guide and outline open research challenges and future directions. The accompanying GitHub repository (https://github.com/AmirNoori68/kan-review) serves as a structured reference for ongoing KAN research.
The following compact, clickable overview summarizes the main topics and organizational structure covered in this review.
Keywords: Kolmogorov–Arnold Networks; Kolmogorov superposition theorem; basis functions; kernel methods; neural network architectures; physics-informed learning; function approximation.
1 Introduction
Within the modern machine learning landscape, multilayer perceptrons (MLPs) remain a foundational architecture for function approximation and data-driven modeling. Recently, Kolmogorov–Arnold Networks (KANs), introduced by Liu et al. [1, 2], have emerged as a promising alternative that rethinks the role of nonlinearities in neural networks. While MLPs rely on fixed activation functions applied at nodes, KANs instead place learnable univariate functions along network edges, a design loosely inspired by the Kolmogorov–Arnold representation theorem. This structural shift has been argued to improve interpretability and adaptability, although empirical studies suggest that its advantages over well-designed MLPs depend strongly on the problem setting and architectural choices. This emphasis on interpretability has also extended to symbolic regression and equation discovery, where KAN-based formulations aim to produce sparser, more structured, and more human-readable representations [3, 4]. To place these developments in context, we first review the role of MLPs and their extensions in scientific machine learning before introducing the KAN framework, whose conceptual origins arise from a different theoretical tradition.
MLPs. Multilayer perceptrons are universal approximators and remain a standard building block for regression, function approximation, and pattern recognition across scientific and engineering applications. A major extension of this paradigm is the physics-informed neural network (PINN) framework of Raissi et al. [5], which has rapidly grown and enabled large-scale studies.
Since then, numerous advances have broadened the scope of MLP-based PINNs. Fractional operators were addressed by fPINNs [6], while uncertainty quantification for forward and inverse problems was introduced in [7]. Extensions such as VPINNs [8], XPINNs for domain decomposition [9], and multi-fidelity PINNs [10] improved accuracy and scalability. The framework also inspired convergence analyses [11], NTK-based diagnostics [12], and practical toolkits like DeepXDE [13]. More recent contributions include multi-stage training with near machine-precision accuracy [14], adaptive residual weighting [15], and distributed solvers for extreme-scale PDEs [16, 17, 18, 19]. Collectively, these developments establish MLP–PINNs as a mature reference standard in scientific machine learning. At the same time, KANs arise from a distinct line of thought rooted in functional representations inspired by the Kolmogorov superposition theorem, rather than as a direct modification of the MLP paradigm.
Despite their flexibility, MLPs face well-documented limitations. Fixed activation functions restrict adaptability [20]. Network behavior can be difficult to interpret [21] and to attribute causally [22]. Achieving high accuracy often entails large parameter counts, which can hinder efficient updates [23, 25, 24]. Generalization robustness can degrade in challenging regimes [277, 26]. Optimization itself can be fragile or stiff, depending on the task and scaling [27]. MLPs also exhibit spectral bias: a tendency to learn low frequencies faster than high frequencies [28, 29]. The effect slows convergence and reduces accuracy for oscillatory or sharp-gradient solutions [30]. This motivates investigations into alternatives with explicit basis control.
KANs. Kolmogorov–Arnold Networks in their modern multilayer form were introduced by Liu et al. [1] as a structured alternative to standard MLPs, in which fixed pointwise activations are replaced by learnable univariate functions defined on network edges and composed across layers. This design is motivated by the Kolmogorov representation theorem, but departs from its classical construction by learning the inner functions directly from data.
The original KAN architecture employs B-spline bases; however, the framework is inherently modular, allowing alternative basis families—such as Chebyshev polynomials, Gaussian kernels, Fourier features, and others—to be substituted without changing the overall architecture. This modularity is fundamental, as the choice of basis directly governs smoothness, locality, spectral content, and numerical stability, thereby influencing expressivity, interpretability, and optimization behavior. As a result, adaptive basis learning combined with interchangeable basis families constitutes a defining feature of KANs, alongside a growing collection of architectural variants, optimization strategies, and theoretical analyses.
In parallel with these methodological developments, KAN implementations have rapidly expanded across open-access platforms, with diverse variants shared by the research community. Curated resources such as [80] [^1] https://github.com/mintisan/awesome-kan provide a centralized overview of recent developments across regression, function approximation, and PDE-solving tasks. Table LABEL:githubs summarizes representative GitHub repositories illustrating this fast-growing ecosystem, which are also referenced throughout this review when discussing specific methods and design choices.
MLPs vs. KANs. When it comes to deciding whether MLP-PINNs or KAN-PIKANs are the better choice, the comparison is far from straightforward, with studies that emphasize fairness often reaching divergent or contradictory conclusions [41] [^2] https://github.com/yu-rp/KANbeFair, [42, 43]. A key reason is that KANs are not a monolithic architecture; their performance is critically influenced by the choice of basis function. For example, a comprehensive comparison by Farea and Celebi [44] [^3] https://github.com/afrah/pinn_learnable_activation found that the optimal basis (e.g., B-spline, Fourier, Gaussian) varies significantly with the PDE being solved, making a simple “KAN vs. MLP” verdict insufficient.
Despite these complexities, Table LABEL:KAN_vs_MLP summarizes reported outcomes across regression, function approximation, and PDE-solving tasks. The entries reflect prevailing tendencies in the literature rather than definitive results for every case. A consistent pattern is that KANs—particularly when equipped with specialized basis functions—tend to match or outperform vanilla MLPs and PINNs in terms of accuracy and convergence speed.
Related Surveys. A number of recent surveys have appeared on KANs, each highlighting different facets of this emerging architecture. Notable contributions include those by Andrade et al. [32], Basina et al. [33], Beatrize et al. [34], Faroughi et al. [35], Kilani et al. [36], Hou et al. [37], Dutta et al. [38], Essahraui et al. [39], and Somvanshi et al. [40], which together provide valuable entry points into the literature. Our review builds on these efforts by seeking to offer a more systematic and comprehensive perspective: rather than cataloguing studies in isolation, we integrate theoretical, architectural, optimization, and application viewpoints into a structured roadmap. By combining comparative analysis, methodological insights, and practical guidance, our goal is to provide a resource that complements existing surveys and serves readers aiming to both understand and apply KANs.
Contributions. This review makes the following contributions:
  • it systematically connects KAN architectures with the Kolmogorov superposition theorem;
  • it comprehensively bridges KANs with classical kernel methods;
  • it compiles and compares reported accuracy, convergence behavior, and computational cost between KANs and MLPs across tasks;
  • it provides a comprehensive and unified overview of basis functions used in KANs;
  • it collects and organizes a wide range of open-source KAN implementations;
  • it offers a structured and integrative perspective on KANs by unifying theoretical foundations, architectural designs, optimization strategies, and applications into a coherent roadmap for both understanding and practice.
Organization and Key Insights. This review is structured to build a comprehensive understanding of Kolmogorov-Arnold Networks, moving from foundational theory to practical application. Section 2 establishes that KANs are not exact realizations of the Kolmogorov Superposition Theorem, but rather data-driven approximations inspired by its compositional structure, using learned smooth functions and modern deep architectures. Section 3 demystifies their behavior by showing that shallow, one-dimensional KANs are mathematically equivalent to classical kernel methods, differing only in how their coefficients are optimized. Section 4 demonstrates the formal equivalence between KANs and MLPs, highlighting that KANs achieve superior parameter efficiency and different optimization dynamics by relocating nonlinearities from nodes to edges.
Building on this, Section 5 presents the choice of basis function as a central design axis, providing a guide for selecting bases (e.g., splines, polynomials, Fourier) to match the properties of the target function (smooth, periodic, or discontinuous). Sections 6 and 7 then catalog practical techniques—such as physics-informed constraints, adaptive grids, decomposition strategies, and parallel implementations—that systematically enhance KAN accuracy and efficiency. Section 8 covers methods like sparsity, pruning, and Bayesian formulations that improve interpretability and generalization.
Finally, Section 9 explains the theoretical advantages of KANs, showing they achieve faster convergence rates, suffer less from spectral bias, and benefit from better-conditioned NTK dynamics. These insights are distilled into a practical ”Choose-Your-KAN” guide in Section 10. We conclude by identifying open research challenges in Section 11 and summarizing our findings in Section 12.
Table 1: Representative performance trends comparing KAN-based models with MLPs. Results are aggregated from the literature and indicate qualitative trends. Reported accuracy and convergence metrics differ across studies (e.g., relative $L_{2}$, RMSE). “Slower training” refers to higher per-iteration cost; total time may vary with epochs. All rows compare KANs with plain MLP/PINN baselines.
Ref. Accuracy Convergence / Time (per iter.) Basis Functions
Regression & Symbolic Representation
[45] BSRBF-KAN $\approx$ MLP Faster convergence; slower training B-spline + Gaussian
[46] MLP-KAN >> MLP B-Spline
[47] KAN << MLP Slower training; less generalizable B-Spline
[48] DE-KAN >> MLP Faster convergence B-spline
[49] KAN-ODE >> MLP-ODEs Faster convergence; slower training Gaussian
[50] KAN-Therm >> MLP Faster convergence B-spline
Function Approximation
[1] KAN $\geq$ MLP Faster convergence; slower training B-spline
[41] KAN >> MLP (symbolic) Slower training B-spline
[51] KAN $\approx$ MLP Faster training Free-knot B-spline
[52] KAN $\geq$ MLP (noisy) Faster convergence; slower training B-spline
[52] KAN $\leq$ MLP (non-smooth) Faster convergence; slower training B-spline
[53] PowerMLP >> MLP Slower training Power-ReLU
[54] SincKAN >> MLP Faster convergence; slower training Sinc
[55] ChebyKAN >> MLP Faster convergence; slower training Shifted Chebyshev
[56] QKAN >> MLP Faster convergence Quantum variational
PDE Solving
[57] KKAN >> PINN Faster convergence Various Basis
[58] PIKAN $\gg$ PINN Faster convergence; slower training B-spline
[…]
[59] DeepOKAN >> DeepONet Faster convergence; slower training Gaussian
[60] PIKAN >> PINN Faster convergence B-Spline
[61] KAN-MHA >> PINN Faster convergence; comparable time B-spline + Attention
[62] Res-KAN $\gg$ PINN Faster convergence; better generalization B-Spline + Residual
[63] HPKM-PINN >> PINN Faster convergence; slower training B-spline
[64] PI-KAN $\gg$ PINN Faster convergence Spline
[65] DPINN $\gg$ PINN Faster convergence; slower training B-spline + Fourier
[66] PIKAN $\approx$ PINN Slower training B-spline
[42] PIKAN $\approx$ PINN slower training Various Basis
[67] EPi-cKAN $\gg$ PINN Slower training; better generalization Chebyshev
[68] Scaled-cPIKAN $\gg$ PINN Faster convergence Chebyshev
[69] PIKAN $\gg$ PINN Faster convergence Chebyshev
[70] tanh-PIKAN >> PINN Chebyshev
[71] AL-PKAN >> PINN Faster convergence B-spline decoder
[72] KANtrol >> PINN Slower training B-spline
[73] PIKAN >> PINN Slower training B-spline
[74] KINN >> PINNs Slower training B-spline
[75] PIKAN $\approx$ PINNs Slower training Chebyshev
[76] PIKAN $\approx$ PINNs Slower training Fourier-based
[77] MR-PIKAN $\gg$ PINN Slower training Chebyshev
[…]
[78] J-PIKAN $\gg$ PINN Faster convergence; slower training Jacobi (orthogonal)
[79] Legend-KINN >> MLP, KAN Faster convergence; slower training Legendre
Table 2: GitHub repositories related to KANs for regression, function approximation, or PDEs.
Repository Description Ref.
…/pykan Official PyKAN for “KAN” and “KAN 2.0”. [2]
…/Anant-Net High-dimensional PDE solver with tensor sweeps. [272]
…/RGA-KANs Deep cPIKANs with variance-preserving initialization. [278]
…/pinn_learnable_activation Compares various KAN bases vs. MLP on PDEs [44]
…/torchkan Simplified PyTorch KAN with variants [123]
…/awesome-kan Curated list of KAN resources, projects, and papers. [80]
…/Deep-KAN Spline-KAN examples and a PyPI package. [124]
…/RBF-KAN RBF-KAN examples [125]
…/KANbeFair Fair benchmarking of KANs vs MLPs. [41]
…/efficient-kan Efficient PyTorch implementation of KAN. [126]
…/spikans Separable PIKAN (SPIKAN). [199]
…/jaxKAN JAX-based KAN package with grid extension support. [275]
…/fast-kan FastKAN using RBFs. [128]
…/faster-kan Using SWitch Activation Function [129]
…/LKAN Implementations of KAN variations. [130]
…/neuromancer>->fbkans Parametric constrained optimization. [131]
…/relu_kan Minimal ReLU-KAN. [132]
…/MatrixKAN Matrix-parallelized KAN. [133]
…/PowerMLP MLP-type network with KAN-level expressiveness. [53]
…/FourierKAN Fourier-based KAN layer. [134]
…/FusedFourierKAN Optimized FourierKAN with fused GPU kernels [135]
…/fKAN Fractional KAN using Jacobi functions. [136]
…/rKAN Rational KAN (Padé/Jacobi rational designs). [137]
…/CVKAN Complex-valued KANs. [138]
…/SincKAN Sinc-based KAN with PINN applications. [54]
…/ChebyKAN Chebyshev polynomial-based KAN variant. [139]
…/OrthogPolyKANs Orthogonal polynomial-based KAN implementations. [140]
…/kaf_act PyTorch activation combining with RFF. [141]
…/KAF Kolmogorov-Arnold Fourier Networks. [76]
…/HRKAN Higher-order ReLU-KANs. [142]
…/KINN PIKAN for solid mechanics PDEs. [74]
…/PIKAN_PointNet Chebushev for solving time-independent inverse problems. [75]
…/FKAN-GCF FourierKAN-GCF for graph filtering. [143]
…/KKANs_PIML Kurkova-KANs combining MLP with basis functions. [57]
…/MLP-KAN MLP-augmented KAN activations. [144]
…/kat Kolmogorov-Arnold Transformer. [145]
…/FAN Fourier Analysis Network (FAN). [146]
…/Basis_Functions Polynomial bases for KANs (comparative study). [147]
…/Wav-KAN Wav-KAN: wavelet-based KANs. [148]
…/qkan Quantum variational activations and pruning tools. [56]
…/KAN-Converge Additive & hybrid KANs for convergence-rate experiments [149]
…/BSRBF_KAN Combines B-splines and radial basis functions. [45]
…/Bayesian-HR-KAN Introduces Bayesian higher-order ReLU-KANs. [150]
…/Legend-KINN Legendre polynomial–based KAN. [79]
…/DeepOKAN Deep Operator Network based on KAN. [59]
…/LeanKAN A memory-efficient Kolmogorov–Arnold Network. [151]
…/KANQAS_code KANQAS: KAN for Quantum Architecture Search. [273]
…/pkan Probabilistic KAN via divisive data re-sorting. [274]
…/pihnn Holomorphic KAN/PIHNN framework for solving PDEs. [271]
…/lmkan Lookup-based KAN for fast high-dim mappings. [284]
…/KAN_Initialization_Schemes Initialization schemes for spline-based KANs. [291]
…/mlp-kan KAN vs. MLP for PDEs in DeepONet/GNS frameworks. [283]
2 Bridging KANs and KST
Hilbert’s Question and Superpositions. At the 1900 International Congress of Mathematicians, Hilbert posed what later became known as Hilbert’s 13th Problem [81]. He asked whether every continuous function of three variables,
$f:[0,1]^{3}\to\mathbb{R},$
can be represented as a finite superposition of continuous functions of at most two variables. Hilbert conjectured that such a representation was impossible, suggesting that genuinely multivariate functions could not be decomposed into simpler lower-arity components.
To clarify what is meant by a superposition in this question—without addressing the existence of such representations—consider a formal expression of the form
$F(x,y,z)=f\!\bigl(g(x,y),\,h(x,\,k(y,z))\bigr),$
which illustrates how a function of three variables may be written as nested compositions of bivariate and univariate functions. Even the elementary identity
$x_{1}+\cdots+x_{n}=x_{1}+\bigl(x_{2}+\cdots+(x_{n-1}+x_{n})\cdots\bigr)$
is a simple instance of a superposition, obtained by repeated binary operations.
Contrary to Hilbert’s conjecture, the seminal results of Kolmogorov and Arnol’d showed that every continuous function of $s$ variables can indeed be represented using only sums and superpositions of univariate continuous functions. This discovery, now known as the Kolmogorov–Arnold superposition theorem, radically changed the understanding of functional representation and forms the theoretical foundation underlying modern Kolmogorov–Arnold Networks.
Kolmogorov and Arnol’d: resolving Hilbert’s 13th Problem. In a short 1956 note [82], Kolmogorov proposed the radical idea that multivariate continuous functions might be representable through superpositions of functions of fewer variables. This was not yet a theorem, but a conceptual breakthrough supported by constructive examples. In 1957, Arnol’d confirmed the conjecture for the three-variable case [83], providing an explicit representation of a general function $f(x,y,z)$ as a superposition of continuous bivariate functions. Later that same year, Kolmogorov established the first complete theorem [84], proving that any continuous function of $n>2$ variables can be expressed using only univariate functions and addition. This result definitively resolved Hilbert’s 13th Problem.
Theorem 1 (Kolmogorov Superposition Theorem (KST) [84])
Let $E_{1}=[0,1]$ and $E_{n}=[0,1]^{n}$. For every integer $n>2$, there exist continuous universal inner functions
$\phi_{q,p}:E_{1}\to\mathbb{R},\qquad p=1,\dots,n,\;\;q=0,\dots,2n,$
independent of the target function $f$, such that every continuous function $f\in C(E_{n})$ admits the representation
$f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\Phi_{q}\!\Bigg(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\Bigg),\qquad(x_{1},\ldots,x_{n})\in E_{n},$ (1)
where the outer functions $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ are continuous and depend on $f$, while the inner functions $\phi_{q,p}$ are universal.
The KST shows that multivariate continuity can be synthesized entirely from univariate nonlinearities and addition, a structural insight that later motivated two–layer interpretations and neural–network analogies.
Proof strategies and later refinements. Two major approaches emerged in the decades following Kolmogorov’s work:
(i) Baire–category proofs: Kahane [85] and Hedberg [86] showed that for generic tuples of continuous increasing functions, the superposition operator is surjective. Their arguments reveal that KST is not only true but abundant: for almost all choices of inner functions, an appropriate outer function exists.
(ii) Constructive proofs: Lorentz [87], Sprecher [88], Arnol’d [89], and others developed explicit (though highly involved) constructions. These proofs follow a common pattern: partition $E_{n}$ into subcubes, map each subcube into a disjoint real interval via the inner functions, and build the outer functions iteratively to approximate the target $f$. Kolmogorov’s original paper left several analytic details unstated; these were filled in by Arnol’d [89] and later authors, making the constructive picture complete.
This classical theory forms the starting point for the refinements described next: reductions in the number of unique functions, improvements in regularity, and the development of computationally usable versions that ultimately connect the superposition viewpoint to modern KAN architectures.
Refinements to KST. Kolmogorov’s constructive formulation (1) established the universality of the inner functions $\phi_{q,p}$, but it did so at considerable structural cost: the representation requires $2n^{2}{+}n$ distinct inner functions and 2n+12n{+}1 outer functions. While theoretically elegant, this level of redundancy makes the original form far too complex for computation or numerical approximation. The decades following Kolmogorov’s work therefore focused on two goals: (i) reducing the number of unique functions needed, and (ii) improving their regularity. These efforts produced the modern, compressed understanding of KST that underlies computational constructions and contemporary reinterpretations, including recent analyses by Ismailov and Ismailov [90].
(i) Reducing the number of functions. Sternfeld [92] showed that the number of summands $2n+1$ in the Kolmogorov superposition formula is optimal in a functional–analytic sense and cannot be reduced within the class of continuous superpositions. However, his functional–analytic framework shows that these 2n+12n{+}1 summands do not require 2n+12n{+}1 different inner or outer functions: the same small collection of functions can be reused repeatedly. Thus the number of distinct functions can be much smaller, even though the number of summands is fixed.
Lorentz’ reduction of outer functions. Lorentz’s 1962 exposition [87] clarified Kolmogorov’s argument and showed that the 2n+12n{+}1 outer functions $\Phi_{q}$ can be replaced by a single continuous function $\Phi$.
Theorem 2 (Kolmogorov superposition in two variables [87])
There exist continuous, monotone increasing functions
$\phi_{1,q},\,\phi_{2,q}:[0,1]\to[0,1],\qquad q=1,\ldots,5,$
such that every continuous function $f:[0,1]^{2}\to\mathbb{R}$ admits the representation
$f(x_{1},x_{2})=\sum_{q=1}^{5}\Phi\!\bigl(\phi_{1,q}(x_{1})+\phi_{2,q}(x_{2})\bigr),$ (2)
where the outer function $\Phi:[0,2]\to\mathbb{R}$ is continuous and depends on $f$.
In this formulation, the ten inner functions $\phi_{1,q}$ and $\phi_{2,q}$ arise because the two-variable case requires five summands, each involving one univariate function of $x_{1}$ and one of $x_{2}$. These inner functions are universal and independent of $f$, while the outer function $\Phi$ depends on the specific target.
More generally, for $s$ variables Lorentz obtained
$f(x_{1},\ldots,x_{s})=\sum_{q=1}^{2s+1}\Phi\!\left(\sum_{p=1}^{s}\phi_{p,q}(x_{p})\right),$ (3)
demonstrating that the dependence on $f$ enters only through a single continuous outer function $\Phi$.
Sprecher’s reduction of inner functions. Sprecher [88] achieved the corresponding simplification for the inner layer. He showed that the entire family $\{\phi_{q,p}\}$ can be generated from a single monotone universal function $\psi:E_{1}\to E_{1}$, together with fixed positive weights $\{\lambda_{p}\}_{p=1}^{n}$ and uniform shifts $x\mapsto x+tq$, where $t>0$ is an arbitrarily small rational parameter.
Theorem 3 (Sprecher 1965 [88])
For each integer $N\geq 2$ there exists a monotone increasing function $\psi:E_{1}\to E_{1}$ satisfying
$\psi\in\mathrm{Lip}\!\left(\frac{\ln 2}{\ln(2N-2)}\right),\qquad\psi(0)=0,\;\psi(1)=1,$
such that for every $n\leq N$, every $f\in C(E_{n})$, and every $\delta>0$, there exists a rational $t$ with $0<t\leq\delta$ for which
$f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}g\!\left(\sum_{p=1}^{n}\lambda_{p}\,\psi(x_{p}+tq)+q\right),$ (4)
where $g:\mathbb{R}\to\mathbb{R}$ is continuous, and the coefficients $\{\lambda_{p}\}_{p=1}^{n}$ together with the function $\psi$ are universal and independent of $f$.
Sprecher’s refinement thus collapses Kolmogorov’s family of inner functions to a single universal function $\psi$, and merges the family of outer functions $\{g_{q}\}$ into a single continuous $g$. In the representation (4), all significant structural complexity lies in the universal components $\psi$ and $\{\lambda_{p}\}$, while the dependence on the target function $f$ is contained solely in the choice of $g$.
By combining Lorentz’ outer-function reduction and Sprecher’s inner-function compression, one obtains the modern compressed version of the superposition theorem:
$f(x_{1},\ldots,x_{n})=\sum_{q=0}^{2n}\Phi\!\Bigg(\sum_{p=1}^{n}\lambda_{p}\,\psi(x_{p}+\varepsilon q)+\delta q\Bigg),$ (5)
where $\psi:E_{1}\to\mathbb{R}$ is a single universal inner function, $\Phi:\mathbb{R}\to\mathbb{R}$ is a continuous outer function depending on $f$, the constants $\{\lambda_{p}\}$ are positive and rationally independent, $\varepsilon>0$ is a uniform shift, and $\delta\in\mathbb{R}$ is an offset introduced for separation of summands.
(ii) Improving regularity. The inner functions in Kolmogorov’s original proof are highly irregular—Hölder continuous but nowhere differentiable. A function $u:E_{1}\to\mathbb{R}$ is Hölder continuous with exponent $\alpha\in(0,1]$ if
$|u(x)-u(y)|\leq L|x-y|^{\alpha},\qquad x,y\in E_{1},$
for some $L>0$. The case $\alpha=1$ corresponds to Lipschitz continuity.
This lack of smoothness is not accidental but a fundamental requirement. Vitushkin and Henkin [93, 94, 95] proved that no universal inner family can possess $C^{1}$ (or higher) smoothness. Although $C^{1}$ regularity implies Lipschitz continuity on compact sets, differentiability itself destroys the Kolmogorov separation property required for exact superposition. As a result, smooth inner functions are incompatible with exact KST representations.
Within the theoretical development of KST, Fridman [99] showed that the universal inner functions may be chosen Lipschitz continuous, corresponding to the maximal regularity compatible with Kolmogorov-type superpositions. Sprecher’s compressed formulation [88] can be realized within this Lipschitz framework. More recently, Actor and Knepley [101, 102] formalized this obstruction by proving that $C^{1}$ inner maps cannot satisfy the Kolmogorov separation property. These results confirm that Lipschitz continuity is optimal, and that any attempt to smooth the universal inner functions necessarily breaks exact KST.
Computational KST. Algorithmic realizations of the Kolmogorov–Arnold superposition theorem may be broadly categorized as follows:
  • Exact constructive KST, where both the universal inner function and the outer functions are generated by an explicit recursive procedure;
  • Approximate KST (outer-only), in which the inner function is fixed and only the outer functions are approximated or learned;
  • Approximate KST (inner–and–outer), where both components are approximated within a superposition-based architecture.
Exact constructions. Early digit-based constructions by Sprecher [103, 104] provided one of the first fully computable realizations of Kolmogorov’s superposition theorem. Köppen [105] later identified flaws in the original recursion, which were corrected in the rigorous formulation of Braun and Griebel [106]. Their construction guarantees continuity, monotonicity, and convergence, and shows that the universal inner function $\psi$ is not merely an abstract object, but can be generated explicitly by a finite, algorithmic recursion.
From an intuitive standpoint, the inner function $\psi$ can be viewed as a universal nonlinear reparameterization of the input. In the Kolmogorov superposition theorem, the original input lies in $0,1]^{n}$, so the parameter $n$ is the dimension of the input space. The role of $\psi$ is to encode information from this $n$-dimensional input into a single scalar variable in a universal way.
The construction of $\psi$ involves three parameters with clearly distinct roles. The dimension parameter $n$ determines how many input variables are being folded into one and therefore directly controls how complex the encoding must be. The base parameter $\gamma>1$ denotes the scaling factor used in the digit-based representation that separates successive refinement levels. The refinement level $k\in\mathbb{N}$ indicates the number of recursive construction steps used to generate the finite approximation $\psi_{k}$ of the inner function. Increasing $k$ refines the same basic geometric pattern, while changing $n$ modifies the structure of that pattern itself.
Figure 1 illustrates this distinction by plotting exact finite-level inner functions $\psi_{k}(x)$ for fixed $k$ and $\gamma$, comparing the cases $n=2$ and $n=4$. When $n=2$, the inner function displays many small visible steps, giving a more oscillatory appearance and stronger apparent nondifferentiability at the plotted scale. When $n=4$, the function appears flatter over longer intervals but exhibits fewer and larger transitions. This occurs because increasing $n$ rapidly suppresses the amplitude of fine-scale refinement steps, causing local irregularities to concentrate into larger but rarer features. As a result, varying $n$ provides clearer insight into how multivariate input dimension influences the geometry of the universal inner function than varying the refinement level alone.
Figure 1: Exact finite-level inner functions $\psi_{k}(x)$ generated by the constructive Kolmogorov–Köppen inner-function scheme for fixed refinement level $k$ and base parameter $\gamma=10$, comparing the cases $n=2$ and $n=4$. The case $n=2$ exhibits many small visible steps, leading to a more oscillatory appearance, whereas $n=4$ shows flatter regions interspersed with fewer but larger transitions due to stronger suppression of fine-scale increments. Left: global 2D view on $[0,1]$. Right: zoomed 2D view on $x\in[0.4,0.5]$, highlighting the contrast in local regularity.
Related constructive and guaranteed approaches include the work of Nakamura et al. [108], who developed algorithmic constructions of both the universal inner map $\psi$ and the outer correction map $\chi$, together with certified approximation error bounds. Together, these contributions clarify not only the theoretical existence of Kolmogorov-type representations, but also their explicit realizability and geometric structure.
Approximate KST (outer-only; inner fixed). This line of work preserves the universal inner map and reduces learning to a one-dimensional outer problem.
  • De Figueiredo (1980): fixes Sprecher’s universal inner function $\psi$ and learns the outer function $g$ using a Chebyshev basis—the first data-driven outer-fitting scheme in a KST layout [111].
  • Frisch et al. (1989): retain Lorentz’s universal inner maps and iteratively refine a single outer function $g$ [119].
  • Braun–Griebel inner + RKHS outer (2009): uses the corrected universal inner function $\psi$ and replaces constructive outer windows by spline/Fourier RKHS approximants [107].
  • Lai–Shen (2025): again fix the Lorentz–Sprecher inner maps and reconstruct the outer function using linear B–splines with $O(n^{-2})$ accuracy [281, 282].
Outer-only methods preserve inner universality and thus avoid the curse of dimensionality: all problem dependence flows through a univariate outer approximation.
Approximate KST (inner–and–outer). These approaches approximate or replace the universal inner functions of the Kolmogorov superposition theorem, thereby departing from the exact Kolmogorov formulation while enabling practical approximation and learning.
  • Kurková (1991–1992): initiated the approximation-theoretic study of the Kolmogorov superposition theorem by interpreting Kolmogorov’s structure in the sense of approximation rather than exact representation. In the Kolmogorov–Sprecher form, the universal inner functions need only satisfy prescribed values on finitely many intervals and can therefore be approximated by staircase-like combinations of a sigmoidal activation. This yields a fully sigmoidal analogue of the Kolmogorov scheme and resolves earlier concerns regarding non-smooth inner maps and non-parametric outer functions [97, 98].
    Theorem 4 (Kurková 1992 [98])
    Let $n\geq 2$, let $\sigma:\mathbb{R}\to I$ be sigmoidal, and let $f\in C(I^{n})$. For every $\varepsilon>0$ there exist $k\in\mathbb{N}$ and functions $\varphi_{i},\psi_{p,i}\in S(\sigma)$ such that
    $\Bigl|f(x_{1},\ldots,x_{n})-\sum_{i=1}^{k}\varphi_{i}\!\Bigl(\sum_{p=1}^{n}\psi_{p,i}(x_{p})\Bigr)\Bigr|<\varepsilon,\qquad\forall(x_{1},\ldots,x_{n})\in I^{n}.$
    Thus, the Kolmogorov superposition $\sum_{p}\phi_{p,q}(x_{p})$ can be replaced by smooth sigmoidal compositions, yielding learnable inner and outer functions within a two-hidden-layer architecture. Kurková’s results provided the first rigorous bridge between Kolmogorov-type superpositions and standard neural network approximation, demonstrating both theoretical universality and practical realizability.
  • Nees (1994, 1996): constructed explicit recursive piecewise-linear approximations of the inner maps $\phi_{q}^{k}$ together with data-driven outer corrections $g_{r}$, achieving geometric error decay [109, 110].
  • Spline-based architectures: The KSN model of Igelnik and Parikh (2003) [113] uses cubic splines for both inner and outer functions, while Coppejans (2005) [114] introduces monotone inner splines for statistical identifiability. Related spline-based superpositional models, including ExSpliNet [115, 116], KASAM [117], and the Urysohn-operator framework of Polar and Poluektov [118], adopt additive or hierarchical structures reminiscent of KST but do not enforce universality of the inner layer.
Caveat. Exact KST requires a universal (target-independent) inner function; approximating or smoothing this function breaks universality and may violate separation. Classical results of Vitushkin and Henkin, and the modern analysis of Actor–Knepley, show that no $C^{1}$ or overly smooth inner functions can satisfy the Kolmogorov separation property [93, 95, 94, 102]. Thus inner approximations are meaningful only in the approximation-theoretic sense, not as exact KST.
Neural interpretations. De Figueiredo (1980) [111] was the first to view the Kolmogorov–Sprecher representation as a computational architecture, expressing it in block-diagram form and proposing a data-driven procedure for learning the outer function $g$ using Chebyshev bases. Later, Hecht–Nielsen (1987) [112] interprets Sprecher’s representation as a three-stage mapping network with a fixed universal hidden layer and learned outer functions. Kurková’s approximation framework (1991–1992) [97, 98] provides compatibility with standard trainable neural networks by replacing the inner staircase with smooth sigmoidal combinations while learning the outer layer. Spline-based models such as KSN (2003) [113] and the shape-constrained variant of Coppejans (2005) [114] train both inner and outer functions within spline families, yielding practical but non-KST architectures. See Table 3 for a chronological overview of these developments, tracing the progression from the original KST theory to its computational and neural interpretations.
Table 3: Development of Kolmogorov’s Superposition Theorem towards Kolmogorov-inspired networks (1900–2024).
Year Author(s) Key Contribution Focus on
1900 Hilbert [81] Poses Hilbert’s 13th problem Theory
1956 Kolmogorov [82] Preliminary idea of superpositions; first hint toward the theorem. Theory
1957 Arnol’d [83] First explicit 3-variable construction with 9 terms; provides concrete counterexample to Hilbert 13. Theory
1957 Kolmogorov [84] Full Kolmogorov Superposition Theorem; first general $n$-D proof, Theory
1958 Arnol’d [89] Supplies missing lemmas; first to complete Kolmogorov’s proof. Theory
1962 Lorentz [87] Gives simplified canonical form with one outer function. Theory
1965 Sprecher [88] First single universal inner function; reduces many $\phi_{p,q}$ to one $\psi$ with shifts; major structural simplification over Kolmogorov. Theory
1967 Fridman [99] Shows universal inner functions may be taken $\mathrm{Lip}(1)$; first optimal smoothness result; strengthens Sprecher–Kolmogorov scheme. Theory
1980 de Figueiredo [111] First network-like interpretation; introduces block diagram + data-driven outer fitting; operationalizes Sprecher’s form. Theory
1987 Hecht–Nielsen [112] First explicit neural mapping theorem; views KST as two-hidden-layer network. Network
1989 Girosi–Poggio [96] First critical analysis of learnability; shows inner function must be non-smooth + outer non-parametric. Theory
1989 Frisch et al. [119] First computational implementation of Lorentz form; iterative learning of single outer function. Network
1991 Kurková [97] First approximation-theoretic reinterpretation; ties network size to modulus of continuity. Network
1992 Kurková [98] Two-hidden-layer sigmoidal approximants; universal inner parameters. Both
1993 Sprecher [120] Single universal $\psi$ for all $n$. Both
1993 Nakamura et al. [108] First fully constructive + guaranteed-accuracy version. Theory
1994 Nees [109] First use of spline (piecewise-linear) inner maps, introducing geometric error decay and a fully constructive, algorithmic approximation of KST with explicit error bounds; Both
1996 Sprecher [103] First executable version of $\psi$ with separation guarantees. Both
1997 Sprecher [104] Explicit algorithm for outer functions. Both
2002 Köppen [105] Corrected continuous monotone inner function of Sprecher [103]; first training-ready inner map. Network
2003 Igelnik–Parikh [113] First fully trainable Kolmogorov Spline Network; replaces fractal inner functions with smooth cubic splines and provides learnable inner and outer maps. Network
2009 Braun–Griebel [106] First correct constructive KST; fixes Sprecher’s errors [103]; modern constructive foundation. Theory
2019 Actor–Knepley [102] Proves $C^{1}$ inner functions impossible; first rigorous smoothness obstruction. Theory
2024 Liu et al. [1] First deep network architecture inspired by the Kolmogorov–Arnold representation (KAN). Network
2024: Liu et al. — Kolmogorov–Arnold Network
Liu et al. [1, 2, 31] introduce the Kolmogorov–Arnold Network, a neural architecture explicitly motivated by the structure of the KST (1), though not intended as a mathematical realization of it. The authors emphasize that KAN is not an exact implementation of the Kolmogorov Superposition Theorem, but rather a computational architecture “inspired by” the Kolmogorov–Arnold representation.
Because the Kolmogorov Superposition Theorem prescribes only a two-layer structure (universal inner functions $\phi_{q,p}$ followed by outer maps $\Phi_{q}$), Liu et al. generalize the idea by redefining the basic computational block. A KAN layer is introduced as a matrix of univariate functions,
$\bm{\Phi}_{\ell}=\{\phi_{\ell,j,i}\},\qquad i=1,\ldots,n_{\ell},\;\;j=1,\ldots,n_{\ell+1},$
where $n_{\ell}$ and $n_{\ell+1}$ denote the widths of layers $\ell$ and $\ell+1$, respectively, and where each edge $(i\!\to\!j)$ carries a trainable B-spline activation $\phi_{\ell,j,i}:\mathbb{R}\to\mathbb{R}$. Within a layer, the forward computation is purely Kolmogorov-like aggregation:
$x_{\ell+1,j}=\sum_{i=1}^{n_{\ell}}\phi_{\ell,j,i}\!\bigl(x_{\ell,i}\bigr),$ (6)
so each layer performs only “inner-type” operations: univariate nonlinearities followed by summation.
In the deep setting, the outputs of one collection of inner-type functions become the inputs to the next. No separate outer function $\Phi_{q}$ is introduced: the final layer plays the role of an outer map only in the trivial sense that it produces the scalar output. Thus a deep KAN is simply the composition of its inner-type layers,
$\mathrm{KAN}(\mathbf{x})=\bigl(\bm{\Phi}_{L-1}\circ\cdots\circ\bm{\Phi}_{0}\bigr)(\mathbf{x}),$ (7)
with “outer-like” effects distributed across all layers rather than concentrated in a single KST-style stage.
Why KANs should not be viewed as exact realizations of the Kolmogorov Superposition Theorem. Although Liu et al. describe KANs as inspired by the Kolmogorov superposition theorem, it is helpful to distinguish KST and KAN as follows:
  • Inner functions are learned, not universal. In KST the inner maps are universal and fixed for all target functions [84, 88], whereas in KANs every edge function is trained from data. This points to a structural analogy rather than an exact correspondence.
  • Regularity properties differ. Classical KST inner functions are only Hölder or Lipschitz continuous [93, 95, 99], whereas KANs use smooth B–splines. This indicates that KANs operate within an approximation framework rather than reproducing the exact KST construction.
  • Separation constraints are not enforced. Constructive proofs of KST rely on fixed shifts and integrally independent weights to separate subcube images [88, 87]. KANs do not impose these geometric constraints, so they do not replicate the mechanism underlying the theorem.
  • Width and depth do not follow KST. Sternfeld’s optimal width 2n+12n{+}1 applies to the shallow superposition setting [92]. KANs allow arbitrary widths and depths, reflecting modern deep-learning practice rather than KST optimality.
  • No final one-dimensional outer function. KST reduces approximation to a single univariate outer map [87, 88]. Deep KANs distribute nonlinearity across multiple layers and do not preserve this one-dimensional reduction.
3 Bridging KANs and Kernel Methods
Kolmogorov–Arnold networks constructed from explicit basis functions admit a direct and instructive comparison with classical kernel methods. This section establishes a concrete bridge between the two paradigms by focusing on how basis functions are combined, how coefficients are parameterized, and how dimensionality is handled. The comparison shows that a shallow, one-dimensional KAN is mathematically equivalent to a kernel regression model, while deeper or multi-dimensional KANs depart from kernel structure in systematic and consequential ways.
Kernel method perspective. A classical kernel or basis-expansion method approximates a target function $f$ by
$f(x)\;\approx\;\sum_{k}c_{k}\,\psi_{k}(x),$
where $\{\psi_{k}\}$ are fixed basis functions (polynomials, splines, Gaussian RBFs, Fourier modes, etc.) and $c_{k}\in\mathbb{R}$ are scalar expansion coefficients. Given data $\{(x_{n},f_{n})\}_{n=1}^{N}$, one forms the design matrix
$\Psi_{n,k}=\psi_{k}(x_{n}),$
and determines the coefficients $c_{k}$ by solving the linear system
$\Psi\,\mathbf{c}\;\approx\;\mathbf{f},$
where $\mathbf{c}$ and $\mathbf{f}$ denote the vectors of expansion coefficients and sampled function values, respectively. Once the basis is fixed, the problem is linear, convex, and solved directly by linear algebra.
We now compare this structure with KANs in three progressively more complex settings.
Case 1: Output-only KANs for one-dimensional problems coincide with kernel methods.
Consider a KAN applied to a one-dimensional problem, with a single scalar input, no hidden layers, and one output neuron. The network produces an approximation $g$ of the target function $f$, which can be written as
$g(x)=\sum_{k}c_{k}\,\phi_{k}(x),$
where $\{\phi_{k}\}$ denotes the chosen univariate basis functions implemented by the KAN and $\{c_{k}\}$ are trainable scalar coefficients. This representation coincides exactly with the standard form of a kernel or basis-expansion model.
The interpretation depends only on the selected basis:
  • for spline-based KANs, $\phi_{k}$ are spline basis functions and $c_{k}$ are spline coefficients;
  • for polynomial-based KANs (e.g. Chebyshev), $\phi_{k}$ are orthogonal polynomials and $c_{k}$ are polynomial coefficients;
  • for Gaussian-based KANs, $\phi_{k}$ are fixed radial basis functions and $c_{k}$ are their amplitudes.
In this output-only setting, the KAN is linear in its parameters and spans the same function class as a classical kernel or basis-expansion model. The coefficient vector $\mathbf{c}$, however, is obtained differently: KANs learn $\mathbf{c}$ via gradient-based optimization of the empirical loss, whereas classical kernel methods typically rely on direct linear-algebraic procedures such as normal equations, matrix inversion, or orthogonal projections. This difference reflects the training strategy rather than the representable function space. As shown in the spectral analysis of shallow KANs in [58], the resulting loss remains quadratic with a Gram-matrix Hessian, enabling conditioning and convergence analysis through linear algebra. Neural-network–specific expressive effects arise only once hidden layers or multi-dimensional compositions are introduced.
Therefore, in the absence of hidden layers, KANs applied to one-dimensional problems are functionally equivalent to classical kernel or basis-expansion models, with differences arising only from the training procedure rather than expressivity.
Case 2: Divergence in multiple dimensions.
The equivalence between KANs and kernel methods no longer holds once the input dimension exceeds one. Classical kernel and basis-expansion methods construct multivariate features by explicitly encoding cross-dimensional interactions. For separable bases, this is achieved through tensor products,
$\psi_{k_{1},\dots,k_{d}}(\mathbf{x})=\prod_{\ell=1}^{d}\psi_{k_{\ell}}(x_{\ell}),$
where $(k_{1},\dots,k_{d})$ is a multi-index and each $\psi_{k_{\ell}}$ is a fixed one-dimensional basis function. For radial basis functions, the dependence is instead on full $d$-dimensional distances,
$\psi_{m}(\mathbf{x})=\varphi\!\big(\|\mathbf{x}-\mathbf{x}_{m}\|\big),\qquad\mathbf{x}_{m}\in\mathbb{R}^{d},$
where $m$ indexes the kernel centers. In both cases, multivariate coupling is embedded directly in the basis functions.
These constructions are expressive but suffer from the curse of dimensionality. For polynomial or spline bases of order $p$, the number of tensor-product basis functions grows as $(p+1)^{d}$, while for radial kernels the number of required centers increases rapidly with $d$ in order to resolve multivariate structure [121, 122].
KANs adopt a different architectural principle. Given an input $\mathbf{x}=(x_{1},\dots,x_{d})$, a standard KAN layer computes the output of neuron $j$ as
$z_{j}=\sum_{\ell=1}^{d}\phi_{j,\ell}(x_{\ell}),\qquad\phi_{j,\ell}\in\mathcal{F}_{1\mathrm{D}},$
where each $\phi_{j,\ell}$ is an independently parameterized univariate function. No tensor-product basis functions $\prod_{\ell}\psi_{k_{\ell}}(x_{\ell})$ and no radial functions $\varphi(\|\mathbf{x}-\mathbf{x}_{m}\|)$ are constructed at the level of a single layer.
Multivariate interactions arise only through the presence of hidden layers. When a second KAN layer is applied to the outputs of the first, the resulting composition produces functions that depend jointly on multiple input coordinates. With increasing depth, such compositions can approximate interaction terms resembling tensor-product or radial features, but these interactions are generated through layer-wise composition rather than encoded explicitly in a single multivariate basis function.
This distinction is structural. As analyzed in [58], the expressive power of KANs in multiple dimensions is governed by depth and compositional structure, whereas classical kernel methods rely on explicit multivariate basis construction. KANs therefore avoid explicit tensorization and achieve expressivity through stacked univariate transformations.
Interpretation. For multivariate problems, KANs can be viewed as structured, low-complexity alternatives to tensor-product or radial kernel spaces. They trade explicit enumeration of all cross-dimensional interactions for a compositional representation that scales more favorably with dimension. This represents a structurally distinct approach to multivariate function approximation rather than a direct kernel approximation.
Case 3: Deep KANs and nonlinear coefficient coupling.
As established in Case 2, hidden layers are required for KANs to represent general multivariate functions. When such hidden layers are introduced, KANs depart most fundamentally from classical kernel and basis-expansion methods. In this deep setting, the coefficients of the resulting function are no longer independent parameters.
This nonlinear coupling arises from function composition across layers and has several consequences:
  • Redundancy. Multiple parameter configurations can represent the same final function, leading to flat directions and equivalent minima.
  • Loss of linear solvability. Unlike kernel methods, the coefficients cannot be recovered by solving a linear system; training requires nonlinear, iterative optimization.
  • Entangled representation. Intermediate features generated by hidden neurons are combined by subsequent layers, obscuring a direct correspondence between individual parameters and the final functional form.
Illustrative example. The effects described above already appear in the simplest one-dimensional deep setting and can be understood through two closely related aspects: growth of the effective polynomial degree and nonlinear coupling of coefficients.
  • Degree growth. Consider a KAN with two Chebyshev layers of degree $P=2$, as in the ChebyKAN construction [139] [^4] https://github.com/SynodicMonth/ChebyKAN, but without any regularization. The first (hidden) layer consists of two neurons, and the second layer produces a scalar output. Although each layer employs only quadratic basis functions, their composition across two layers yields an output polynomial of degree $4$. More generally, for a one-dimensional KAN with $L$ stacked polynomial layers of degree $P$, the effective polynomial degree satisfies
    $\deg(g)=P^{\,L},$
    where $L$ denotes the number of layers. This growth is a direct consequence of function composition and does not arise in shallow models. In this setting, high-degree terms are generated automatically, and training must adjust the parameters so that the corresponding coefficients remain close to zero when the target function does not require such a high polynomial degree.
    However, as illustrated in the item below, the same parameters that control high-degree terms also contribute to lower-degree coefficients, making it nontrivial for optimization to suppress unwanted high-degree components without affecting the lower-degree structure of the solution.
  • Nonlinear coefficient coupling. The same compositional mechanism that increases polynomial degree also changes how the polynomial coefficients depend on the underlying parameters. Let $A^{(1)}\in\mathbb{R}^{3\times 2}$ and $A^{(2)}\in\mathbb{R}^{6\times 1}$ denote the parameter matrices of the first and second layers, respectively, with entries
    $A^{(1)}=\big(a_{ij}\big),\qquad A^{(2)}=\big(b_{ij}\big).$
    For this two-layer Chebyshev–KAN with per-layer polynomial degree $P=2$, collecting powers of $x$ yields the equivalent power-form
    $g(x)=C_{0}+C_{1}x+C_{2}x^{2}+C_{3}x^{3}+C_{4}x^{4},$
    where the coefficients are
    $C_{4}$ $=8{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}^{2}b_{12}+8a_{22}^{2}b_{15},$
    $C_{3}$ $=8a_{11}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}\,b_{12}+8a_{21}a_{22}b_{15},$
    $C_{2}$ $=2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}\,b_{11}+\big(8a_{10}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}+2a_{11}^{2}-8{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}^{2}\big)b_{12}+2a_{22}b_{14}+\big(8a_{20}a_{22}+2a_{21}^{2}-8a_{22}^{2}\big)b_{15},$
    $C_{1}$ $=a_{11}b_{11}+\big(4a_{10}a_{11}-4a_{11}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}}\big)b_{12}+a_{21}b_{14}+\big(4a_{20}a_{21}-4a_{21}a_{22}\big)b_{15},$
    $C_{0}$ $=b_{10}+b_{13}-b_{12}-b_{15}+(a_{10}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}})b_{11}+2(a_{10}-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}})^{2}b_{12}+(a_{20}-a_{22})b_{14}+2(a_{20}-a_{22})^{2}b_{15}.$
    For comparison, a classical Chebyshev approximation truncated at degree $4$ has coefficients $\mathbf{c}=(c_{0},\dots,c_{4})^{\top}$ determined by a linear system of the form $\Psi\,\mathbf{c}=\mathbf{f}$ (or its least-squares variant), and admits the power-form
    $f(x)\;\approx\;(c_{0}-c_{2}+c_{4})+(c_{1}-3c_{3})x+(2c_{2}-8c_{4})x^{2}+(4c_{3})x^{3}+(8c_{4})x^{4},$
    where each polynomial coefficient depends linearly on a single entry of $\mathbf{c}$.
    The distinction is therefore explicit. In the classical case, the mapping from data to coefficients is linear and fully described by the design (kernel) matrix $\Psi$. In the deep KAN, each coefficient $C_{k}$ is a nonlinear function of parameters from multiple layers. Terms such as $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{12}$} arise from composition and appear in several coefficients simultaneously, meaning that a single parameter influences multiple polynomial orders. Consequently, increasing depth not only raises the effective polynomial degree, but also couples the coefficients nonlinearly, making their identification a fundamentally different problem from solving a linear kernel system, even in one-dimensional settings.
Summary. Case 1 establishes that shallow KANs applied to one-dimensional problems coincide with classical kernel or basis-expansion models, differing only in training procedure. Case 2 shows that for multivariate inputs, KANs abandon explicit tensor-product or radial constructions in favor of factorized, additive representations, and instead rely on hidden layers and composition to approximate multivariate interactions. Case 3 demonstrates that once hidden layers are introduced, depth induces both multiplicative growth of polynomial degree and nonlinear coupling of coefficients, fundamentally departing from classical kernel structure. Together, these mechanisms explain the expressive power of deep KANs as well as the associated challenges in optimization, conditioning, and parameter identifiability, motivating kernel-inspired regularization and stability-aware training strategies.
4 Bridging KANs and MLPs
MLPs are feedforward neural networks composed of stacked affine layers followed by element-wise nonlinear activations. Given an input vector $\mathbf{x}\in[0,1]^{n}$, an MLP with $L$ layers produces an output $\hat{f}(\mathbf{x};\bm{\theta})\in\mathbb{R}$ via
$\mathbf{z}^{(0)}$ $=\mathbf{x},$ (8)
$\mathbf{z}^{(\ell)}$ $=\sigma\!\left(\mathbf{W}^{(\ell)}\mathbf{z}^{(\ell-1)}+\mathbf{b}^{(\ell)}\right),\quad\ell=1,\dots,L-1,$ (9)
$\hat{f}(\mathbf{x};\bm{\theta})$ $=\mathbf{W}^{(L)}\mathbf{z}^{(L-1)}+\mathbf{b}^{(L)},$ (10)
where
  • $\sigma(\cdot)$ is a nonlinear activation function (e.g., ReLU, $\tanh$),
  • $\mathbf{W}^{(\ell)}\in\mathbb{R}^{n_{\ell}\times n_{\ell-1}}$ and $\mathbf{b}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ are the weights and biases at layer $\ell$,
  • $\mathbf{z}^{(\ell)}\in\mathbb{R}^{n_{\ell}}$ is the hidden representation at layer $\ell$,
  • $n_{\ell}$ is the width (number of neurons) of layer $\ell$,
  • $\bm{\theta}=\{\mathbf{W}^{(\ell)},\mathbf{b}^{(\ell)}\}_{\ell=1}^{L}$ collects all trainable parameters.
KANs and MLPs are both hierarchical function approximators, yet they differ fundamentally in where and how nonlinearities are applied. In a conventional MLP, inputs are first linearly mixed and then passed through a fixed activation function (mix $\rightarrow$ activate). In contrast, a KAN applies a (typically trainable) univariate transformation to each input coordinate before aggregation (activate $\rightarrow$ mix). This reversal of operations produces more localized, interpretable, and adaptable mappings.
This structure closely parallels Sprecher’s constructive version of the Kolmogorov superposition theorem (5), where inner univariate functions act independently on shifted coordinates:
$f(x_{1},\ldots,x_{n})=\sum_{q}\Phi\!\biggl(\smash{\underbrace{\sum_{p}\lambda_{p}\,\phi(x_{p}+\eta_{q})}_{\begin{subarray}{c}\lambda_{p}:\text{ weights}\\ \eta_{q}:\text{ shifts}\end{subarray}}}\;+\;\smash{\underbrace{c_{q}}_{\text{offset}}}\biggr),$
with positive weights $\lambda_{p}>0$, shifts $\eta_{q}$, and offsets $c_{q}$. The KAN formulation operationalizes this constructive principle within a neural architecture, turning theoretical decomposition into a learnable process.
Formal Equivalence. Wang et al. [58] establish a precise bidirectional correspondence between multilayer perceptrons (MLPs) and Kolmogorov–Arnold Networks (KANs), showing that the two architectures are formally equivalent under mild structural assumptions.
Theorem 5 (MLP $\Rightarrow$ KAN, [58])
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain, and suppose that a function $f:\mathbb{R}^{d}\to\mathbb{R}$ can be represented by an MLP of width $W$, depth $L$, and activation function
$\sigma_{k}(x)=\max(0,x)^{k},\qquad k\geq 1.$
Then there exists a KAN of width $W$, depth at most $2L$, grid size $G=2$, and $k$-th order B-spline basis functions such that
$f(\mathbf{x})=g(\mathbf{x}),\qquad\forall\mathbf{x}\in\Omega.$
Concretely, an MLP representation of the form
$f(\mathbf{x})=\sum_{i=1}^{W}\alpha_{i}\,\sigma_{k}\!\Bigl(\sum_{j=1}^{d}w_{ij}x_{j}+b_{i}\Bigr)$
admits an equivalent KAN parameterization
$f(\mathbf{x})=\sum_{i=1}^{W}\Phi_{i}\!\biggl(\sum_{j=1}^{d}\varphi_{ij}(x_{j})\biggr),$
where each univariate map $\varphi_{ij}:\mathbb{R}\to\mathbb{R}$ is a learned B-spline that approximates the composite nonlinearity $\sigma_{k}(w_{ij}x_{j}+b_{i})$. From this perspective, KANs replace fixed pointwise activations in MLPs by learnable one-dimensional functions along each edge, without increasing the architectural width.
The reverse direction holds under mild restrictions on the univariate components. Non-polynomial nonlinearities (such as SiLU) cannot be captured exactly by $\mathrm{ReLU}^{k}$ networks, but if a KAN employs only polynomial-type univariate functions—such as B-splines of order kk—then an equivalent MLP representation exists.
Theorem 6 (KAN $\Rightarrow$ MLP, [58])
Suppose that a function $f:[0,1]^{d}\to\mathbb{R}$ is represented by a KAN of width $W$, depth $L$, grid size $G$, and $k$-th order B-spline univariate functions, and assume that no non-polynomial activations are present. Then there exists an MLP with activation $\sigma_{k}(x)=\max(0,x)^{k}$, depth at most $2L$, and width $(G+2k+1)W^{2}$ that represents $f$.
The induced MLP has parameter complexity $\mathcal{O}(G^{2}W^{4}L)$, whereas the original KAN requires only $\mathcal{O}(GW^{2}L)$ parameters. This gap highlights the superior parameter efficiency of KANs, particularly for fine spline grids or higher-order bases. As a consequence, approximation and convergence results developed for $\mathrm{ReLU}^{k}$ MLPs can be transferred directly to KANs through this formal equivalence (see Section 9).
Special Case: Piecewise-Linear Functions. Schoots et al. [173] demonstrate that KANs using piecewise-linear univariate functions are functionally identical to ReLU-based MLPs. Any piecewise-linear $\varphi:\mathbb{R}\to\mathbb{R}$ with $k$ breakpoints admits the expansion
$\varphi(x)=\underbrace{a_{0}}_{\text{bias}}+\underbrace{a_{1}x}_{\text{linear term}}+\sum_{j=1}^{k}\underbrace{\alpha_{j}}_{\text{coeff.}}\,\mathrm{ReLU}\!\bigl(x-b_{j}\bigr),$
where $a_{0},a_{1},\alpha_{j},b_{j}\in\mathbb{R}$. Thus, each KAN univariate map can be embedded into a compact ReLU subnetwork, reinforcing the view of KANs as structured and interpretable MLPs.
Actor et al. [51] and Gao et al. [174] reach similar conclusions: KANs can emulate MLP behavior while introducing task-adaptive nonlinearities—such as spline, Fourier, or Chebyshev bases—that improve inductive bias and generalization for structured data.
Summary. KANs and MLPs are expressively equivalent but structurally distinct. KANs achieve localized, interpretable representations by applying learnable univariate transformations before aggregation, whereas MLPs rely on fixed activations after linear mixing. This reversal leads to substantial parameter savings and smoother function representations, while maintaining theoretical consistency with the Kolmogorov–Arnold superposition principle. In practice, KANs can thus be viewed as structured, efficient, and interpretable extensions of MLPs.
4.1 How KANs Extend the MLP Landscape
KAN layers are increasingly adopted as drop-in replacements for traditional MLP blocks across convolutional, transformer, graph, and physics-informed models. They enhance expressivity, interpretability, and parameter efficiency—often with minimal architectural or training changes.
Vision and Representation Learning. In computer vision, Convolutional KANs embed spline-based activations directly within convolutional kernels, producing more expressive yet lightweight CNNs [175] [^5] https://github.com/AntonioTepsich/Convolutional-KANs and [305]. Residual KAN modules integrate into ResNet backbones to improve gradient flow and generalization [180] [^6] https://github.com/withray/residualKAN, while U-KAN extends U-Nets for image segmentation and diffusion models [182] [^7] https://github.com/CUHK-AIM-Group/U-KAN. Similarly, the KAN-Mixer adapts MLP-Mixer architectures for image classification [195] [^8] https://github.com/engichang1467/KAN-Mixer. Additional applications include remote sensing [184], hyperspectral imaging via Wav-KAN [183], medical image classification [185], and autoencoding tasks [181] [^9] https://github.com/SekiroRong/KAN-AutoEncoder.
Sequential and Temporal Modeling. For time-series and sequential tasks, KAN-AD employs Fourier expansions for efficient anomaly detection [186] [^10] https://github.com/issaccv/KAN-AD, while T-KAN and MT-KAN enhance forecasting under concept drift and multivariate dependencies [188]. Other extensions include satellite-traffic forecasting [187], recurrent temporal KANs (TKANs[189], and the transformer-based TKAT model [190] [^11] https://github.com/remigenet/TKAT, which integrates learnable univariate mappings within self-attention layers.
Graph and Structured Data. KAN layers have also been incorporated into graph learning frameworks, where replacing MLPs in GNN message-passing blocks improves numerical stability and feature smoothness. Implementations such as GraphKAN [191] [^12] https://github.com/WillHua127/GraphKAN-Graph-Kolmogorov-Arnold-Networks and KAN4Graph [192] [^13] https://github.com/yueliu1999/KAN4Graph achieve consistent accuracy gains. GKAN integrates spline-based kernels directly into graph convolutions [193], while general-purpose variants like S-KAN and S-ConvKAN enable task-dependent activation selection across architectures [43].
Physics-Informed and Operator Learning. In scientific computing, Physics-Informed KANs preserve the original PINN objective but replace fixed activations with learnable basis functions, improving locality and interpretability. This modularity enables direct architectural transfers—e.g., DeepONet $\rightarrow$ DeepOKAN, separable PINNs $\rightarrow$ separable PIKANs, and even NTK analyses in PINNs $\rightarrow$ corresponding analyses in PIKANs. Moreover, existing advances such as variational losses, residual reweighting, and adaptive sampling naturally carry over, bridging traditional PINN frameworks with modern kernel-based representations. Representative examples are summarized in Table 4.
Table 4: Representative advanced PINN baselines [154, 155] and corresponding KAN/PIKAN counterparts.
PINN contribution KAN/PIKAN counterpart
(2017-2025) (2024-2025)
2017—[5] Foundational PINN for data-driven PDE solutions [1]
2018—[7] Uncertainty quantification for forward/inverse PINNs [150]
2019—[177] Neural ordinary differential equations (Neural ODEs) [49]
2019—[158] Neural operators for PDE solution maps (DeepONet) [59]
2020—[246] PINN for symbolic regression via recursive decomposition [247]
2020—[10] Multi-fidelity PINNs (low/high-fidelity correlation) [198]
2020—[161] Fourier-feature embeddings for multiscale structure [76]
2020—[254] Fourier Neural Operator (FNO) [253]
2021—[13] DeepXDE library (resampling strategies and benchmarks) [1]
2022—[12] NTK-based analysis explaining PINN training pathologies [174, 68]
2023—[167] Adaptive PINN (moving collocation points) [127]
2023—[227] Finite-basis PINNs with overlapping subdomains [131]
2023—[163] Separable PINN architectures [199]
2023—[176] Surrogate + PDE-constrained optimization [61]
2023—[178] Surrogate modeling in elasto-plasticity [67]
2025—[179] Multigrid and multi-resolution training strategy [77]
5 Basis Functions
This section reviews commonly used bases in KANs: B-splines, Chebyshev and Jacobi polynomials, ReLU compositions, Fourier series, Gaussian kernels, wavelets, finite-basis partitions, and Sinc functions. We begin with an overview (Table 5) to orient the reader. Subsequent subsections expand on each family, beginning with B-splines.
Table 5: Comparison of Basis Functions in KANs
Name Support Equation Form Grid Basis/Activation Ref.
Required Type
B-spline Local $\sum_{n}c_{n}B_{n}(x)$ Yes B-spline [1]
Chebyshev Global $\sum_{k}c_{k}T_{k}(\tanh x)$ No Chebyshev + Tanh [205]
Stabilized Chebyshev Global $\tanh\!\Big(\sum_{k}c_{k}\,T_{k}(\tanh x)\Big)$ No Chebyshev + linear head [70]
Chebyshev (grid) Global $\sum_{k}c_{k}T_{k}\!\Big(\tfrac{1}{m}\sum_{i}\tanh(w_{i}x+b_{i})\Big)$ Yes Chebyshev + Tanh [57]
ReLU-KAN Local $\sum_{i}w_{i}R_{i}(x)$ Yes Squared ReLU [132]
HRKAN Local $\sum_{i}w_{i}[\mathrm{ReLU}(x)]^{m}$ Yes Polynomial ReLU [142]
Adaptive ReLU-KAN Local $\sum_{i}w_{i}v_{i}(x)$ Yes Adaptive ReLU [127]
fKAN (Jacobi) Global $\sum_{n}c_{n}P_{n}(x)$ No Jacobi [136]
rKAN (Padé/Jacobi) Global $\tfrac{\sum_{i}a_{i}P_{i}(x)}{\sum_{j}b_{j}P_{j}(x)}$ No Rational + Jacobi [137]
Jacobi-KAN Global $\sum_{i}c_{i}P_{i}(\tanh x)$ No Jacobi + Tanh [75]
FourierKAN Global $\sum_{k}a_{k}\cos(kx)+b_{k}\sin(kx)$ No Fourier [143]
KAF Global $\alpha\,\mathrm{GELU}(x)+\sum_{j}\beta_{j}\psi_{j}(x)$ No RFF + GELU [76]
Gaussian Local $\sum_{i}w_{i}\exp\!\Big(-\big(\tfrac{x-g_{i}}{\varepsilon}\big)^{2}\Big)$ Yes Gaussian RBF [128]
RSWAF-KAN Local $\sum_{i}w_{i}\left(s_{i}-\tanh^{2}\!\big(\tfrac{x-c_{i}}{h_{i}}\big)\right)$ Yes Reflectional Switch [129]
CVKAN Local $\sum_{u,v}w_{uv}\exp\!\big(-|z-g_{uv}|^{2}\big)$ Yes Complex Gaussian [138]
BSRBF-KAN Local $\sum_{i}a_{i}B_{i}(x)+\sum_{j}b_{j}\exp\!\big(-\tfrac{(x-g_{j})^{2}}{\varepsilon^{2}}\big)$ Yes B-spline + Gaussian [45]
Wav-KAN Local $\sum_{j,k}c_{j,k}\,\psi\!\big(\tfrac{x-u_{j,k}}{s_{j}}\big)$ No Wavelet [148]
FBKAN Local $\sum_{j}\omega_{j}(x)K_{j}(x)$ Yes PU + B-spline [131]
SincKAN Global $\sum_{i}c_{i}\,\mathrm{Sinc}\!\big(\tfrac{\pi}{h}(x-ih)\big)$ Yes Sinc [54]
Poly-KAN Global $\sum_{i}w_{i}P_{i}(x)$ No Polynomial [147]
5.1 B-spline
B-spline bases are among the most widely adopted in KANs due to their compact support, smoothness, local control, and piecewise–polynomial structure [51, 196, 133, 62, 66, 174, 52, 64, 65, 48, 197, 47, 198, 199, 72, 200, 201, 127, 73, 74, 71, 202, 173, 58, 203, 63, 204, 61, 131, 124, 222, 260, 253, 50, 255]. Their locality and flexibility make them expressive, numerically stable, and particularly well suited for interpretable representations. The original KAN formulation [1] and its accompanying software package further contributed to their early adoption and widespread use.
Each univariate KAN map $\varphi:\mathbb{R}\to\mathbb{R}$ is represented as a linear combination of fixed–knot B-spline basis functions:
$\varphi(x)\;=\;\sum_{n=0}^{N-1}c_{n}\,B_{n}^{(k)}(x),$ (11)
where $N$ is the number of basis functions, $B_{n}^{(k)}(x)$ denotes the $n$-th B-spline of polynomial degree $k$ defined on a knot vector $\mathbf{t}=(t_{0},\ldots,t_{N+k})$, and $c_{n}\in\mathbb{R}$ are trainable coefficients. The Cox–de Boor recursion yields $C^{k-1}$ continuity, and open–uniform knot choices give well–behaved boundary conditions.
Variants of this formulation also employ fractional B-splines. Aghaei and Zaky [310] introduce a physics-informed fractional KAN in which fractional B-spline bases are combined with a tensorized operational-matrix formulation to efficiently solve distributed-order fractional differential equations. These variants extend the classical spline basis to fractional operators but retain the same underlying KAN framework.
For the standard spline construction used in most KAN implementations, the knot vector is often extended by $k$ points on each side. This padding allows the model to progressively capture finer details [58]. As illustrated in Figure 2, the number of active basis functions grows from GkG-k to $G+k$, where $G$ is the number of internal intervals in the original, non–extended grid.
Subfigures 2(a) and 2(b) visualize cubic (k=3k{=}3) B-spline bases on non–extended and extended grids, respectively. Without extension, boundary bases are truncated and lose symmetry; with extension, added knots outside the domain restore full polynomial support (gray bands). To compare expressivity, subfigures 2(c) and 2(d) synthesize univariate maps
$\varphi(x)\;=\;\sum_{n=0}^{N-1}c_{n}B_{n}^{(k)}(x),$
using the same coefficient vector across the two grids (truncated in the non–extended case). Differences in $\varphi$ thus arise solely from the basis itself: the extended grid yields richer boundary behavior and overall smoother maps, while the non–extended grid flattens near edges. This highlights the practical benefit of grid extension for boundary–sensitive tasks (e.g., PDEs).
(a) Non–extended cubic B-spline basis
(b) Extended cubic B-spline basis
(c) Learned map $\varphi(x)$ on non–extended grid
(d) Learned map $\varphi(x)$ on extended grid
Figure 2: Comparison of B-spline bases and synthesized univariate maps with/without grid extension. (a,b) Cubic (k=3k{=}3) bases on non–extended vs. extended grids. (c,d) Learned maps $\varphi(x)$ using identical random coefficients $c_{n}$. Gray regions mark padded boundary intervals.
In the original KAN [1] and KAN 2 [2] [^14] https://github.com/KindXiaoming/pykan, each univariate activation uses uniformly spaced cubic B-splines ($k=3$) together with a smooth residual term to aid gradients and enhance expressivity in flat regions:
$\tilde{\varphi}(x)\;=\;\sum_{n=0}^{N-1}c_{n}\,B_{n}^{(k)}(x)\;+\;\frac{x}{1+e^{-x}}.$
A global tanh is applied after each hidden layer (but not the output) to keep activations within the spline domain. KAN 2 further supports post–training grid refinement, which increases knot resolution without reinitializing parameters, thereby improving fidelity with minimal overhead.
Convergence implications. Recent theory gives spline-KANs some of the strongest guarantees in the KAN family. Wang [58] proves accelerated Sobolev approximation rates; Kratsios et al. [287] show optimal Besov performance on Lipschitz and even fractal domains; and Liu et al. [149] establish minimax-optimal regression rates together with optimal spline–knot scaling. These results place B-splines on the firmest theoretical footing among KAN bases. (See Sec. 9.)
Spectral bias and NTK conditioning. Spline grids control frequency reach: finer grids reduce spectral bias but may worsen conditioning [58, 44]. Initialization also matters—Rigas et al. [291] [^15] https://github.com/srigas/KAN_Initialization_Schemes show that Glorot-style and power-law initializations yield more stable NTKs and faster early convergence. (See Secs. 9.2 and 9.3.)
5.2 Chebyshev Polynomials
Chebyshev polynomials offer a clean and theoretically grounded alternative to B-splines for KAN univariate mappings—especially in PIKANs—due to their global orthogonality, excellent spectral approximation properties, and simple recursive structure suited for smooth or oscillatory targets.
Definition and Recurrence. Let $T_{k}(x)$ denote the Chebyshev polynomials of the first kind, defined recursively as
$T_{0}(x)=1,\qquad T_{1}(x)=x,\qquad T_{k}(x)=2x\,T_{k-1}(x)-T_{k-2}(x)\quad(k\geq 2),$ (12)
which form an orthogonal basis on [1,1][-1,1] with weight $w(x)=(1-x^{2})^{-1/2}$.
ChebyKAN Formulation. The ChebyKAN model [139] [^16] https://github.com/SynodicMonth/ChebyKAN represents each univariate map as
$\varphi^{(\ell)}_{q,p}(x)\;=\;\sum_{k=0}^{K}c^{(\ell)}_{q,p,k}\,T_{k}\!\big(\tilde{x}\big),\qquad\tilde{x}=\tanh(x)\ \ \text{(per-layer input normalization)},$
with trainable coefficients $c^{(\ell)}_{q,p,k}\in\mathbb{R}$. Each layer thus evaluates $T_{k}$ on normalized inputs $\tilde{x}\in[-1,1]$ for numerical stability, producing the output
[…]
consistent with the generic KAN layer formulation.
Efficient Evaluation. In practice, most implementations (e.g., [77]) compute $T_{k}$ via
$T_{k}(z)=\cos\!\left(k\,\arccos z\right),$
which vectorizes efficiently on GPUs. Alternatively, the recurrence relation (12) or Clenshaw’s algorithm provides equivalent, numerically stable computation—particularly advantageous for high-degree expansions where repeated $\arccos$ calls become costly. Mahmoud et al. [55] further employ shifted Chebyshev polynomials on $[0,1]$, defined as
$T_{n}(x)=\cos\!\big(n\,\arccos(2x-1)\big),$
to align the domain with the standard Kolmogorov–Arnold representation on non-symmetric intervals.
Normalization and Stability. Figure 3 illustrates how per-layer $\tanh$ normalization stabilizes Chebyshev activations. Panel (a) shows the standard basis $T_{k}(x)$ on [1,1][-1,1] for degree K=8K{=}8, where extrema occur at $T_{k}(\pm 1)=\pm 1$, forcing steep slopes near the boundaries. Panel (b) instead evaluates $T_{k}(\tanh x)$, compressing the effective range to $\tanh(\pm 1)\approx\pm 0.762$, which moderates endpoint slopes while preserving interior structure. This compression prevents saturation at $\pm 1$ and keeps features in a well-conditioned range, critical for deep stacks of Chebyshev layers.
The difference becomes clearer in Panel (c), comparing the deep composite maps
$\sum_{k=0}^{K}c_{k}\,T_{k}(\tanh x)\quad\text{and}\quad\sum_{k=0}^{K}c_{k}\,T_{k}(x)$
with identical coefficients $c_{k}$. Both are bounded, but by the chain rule,
$\frac{d}{dx}\!\left[\sum_{k=0}^{K}c_{k}\,T_{k}(\tanh x)\right]=\bigl(1-\tanh^{2}x\bigr)\,\sum_{k=0}^{K}c_{k}\,T_{k}^{\prime}\!\big(\tanh x\big),$
the $\tanh$-normalized version suppresses slope growth near $|x|\approx 1$, removing the synchronized oscillations visible in the unnormalized case. Even without training, this normalization yields smoother and better-conditioned responses.
(a) Chebyshev basis without $\tanh$ normalization
(b) Chebyshev basis with $\tanh$ normalization
(c) Deep Chebyshev KAN map: with vs. without $\tanh$ normalization
Figure 3: (a) Standard Chebyshev basis functions without per-layer $\tanh$ normalization. (b) Same basis with $\tanh$ normalization, showing compressed input range and moderated edge slopes. (c) Deep Chebyshev KAN map (K=8K{=}8) comparing $\sum c_{k}T_{k}(\tanh x)$ (blue) vs. $\sum c_{k}T_{k}(x)$ (orange); the normalized version exhibits smoother behavior and smaller endpoint gradients, indicating improved conditioning.
Grid-Averaged Variant. Following Toscano et al. [57], a practical enhancement introduces grid before evaluating the Chebyshev expansion. For a univariate input $x$,
$g(x)=\frac{1}{m}\sum_{i=1}^{m}\tanh\!\big(w_{i}x+b_{i}\big),\qquad\varphi^{(\ell)}_{q,p}(x)=\sum_{k=0}^{K}c^{(\ell)}_{q,p,k}\,T_{k}\!\big(g(x)\big),$
where $m$ denotes the number of centers, $b_{i}$ are uniformly spaced biases within $\beta_{\min},\beta_{\max}]$ (e.g., 0.1-0.1 to $0.1$), and $w_{i}$ control local slope. The coefficients $c^{(\ell)}_{q,p,k}$ can follow the stable initialization proposed in [205]. This averaged-$\tanh$ grid constrains $g(x)\in(-1,1)$, introduces mild spatial warping, improves numerical conditioning, and as reported in [57], supports stable training even with higher polynomial degrees.
Spectral Properties and Efficiency. The spectral parameterization of Chebyshev-based KANs was first formalized in Sidharth et al. [205], while Guo et al [206] demonstrated superior parameter efficiency and generalization in data-scarce regimes. From a theoretical perspective, Faroughi and Mostajeran [69] showed that Chebyshev PIKANs (cPIKANs) yield better-conditioned NTKs with slower spectral decay, accelerating convergence for PDEs such as diffusion and Helmholtz equations.
Stabilization and Hybrid Designs. Yu et al. [54] confirmed these advantages for function approximation and PDE learning but also highlighted failure modes of raw polynomial stacks at high depth, motivating stabilizers such as domain normalization, nested nonlinearities, and contractive mappings. Building on these insights, Daryakenari et al. [70] proposed a stabilized Chebyshev stack by inserting an additional $\tanh$ between layers and replacing the Chebyshev head with a linear readout:
$x^{(\ell+1)}_{q}\;=\;\tanh\!\Bigg(\sum_{p=1}^{P}\sum_{k=0}^{K}c^{(\ell)}_{q,p,k}\,T_{k}\!\big(\tanh(x^{(\ell)}_{p})\big)\Bigg),\qquad\ell=0,\dots,L-1,$ (13)
with the network output
$\hat{f}(\mathbf{x})\;=\;\mathbf{W}^{\text{out}}\,\mathbf{x}^{(L)}+\mathbf{b}^{\text{out}},$ (14)
where $\mathbf{x}^{(L)}=[x^{(L)}_{1},\dots,x^{(L)}_{H_{L}}$$^{\top}$. The inter-layer $\tanh$ in (13) acts as a contraction, curbing gradient growth [54], while the linear head (14) isolates the final mapping from additional polynomial expansions, improving stability in inverse and PDE learning tasks [70].
Depth Stabilization. Deep Chebyshev KANs often become unstable. Rigas et al. [278] [^17] https://github.com/srigas/RGA-KANs mitigate this with a basis-agnostic Glorot-style initialization that preserves activation variance, and a Residual–Gated Adaptive (RGA) KAN block that stabilizes deep cPIKANs on PDEs. One-liner: variance-preserving init + residual gating enable deep Chebyshev KANs to train without divergence.
Summary. Chebyshev-based KANs combine spectral approximation theory with practical stability mechanisms. Their global orthogonality and efficient recursion yield compact and interpretable representations, while nested nonlinearities and grid-averaged normalization ensure stable deep training. These features make them particularly well suited for scientific computing, operator learning, and inverse problems—complementing and often surpassing spline-based KANs within the broader KAN framework.
5.3 ReLU
ReLU–based KANs (ReLU-KANs) were introduced by Qiu et al. [132] [^18] https://github.com/quiqi/relu_kan as a hardware–efficient alternative to B-spline KANs. The key idea is to replace spline activations with compactly supported, bell–shaped functions constructed from ReLU compositions. This preserves the localized, compositional spirit of Kolmogorov–Arnold models while enabling fast, GPU–friendly primitives.
Local ReLU bases and univariate maps. In ReLU-KANs, each univariate map $\varphi:\mathbb{R}\to\mathbb{R}$ is a weighted sum of compact local bases:
$\varphi(x)\;=\;\sum_{i=0}^{G+k-1}w_{i}\,R_{i}(x),$
where $w_{i}\in\mathbb{R}$ are trainable weights, $G$ is the number of grid intervals, and $k$ controls overlap among neighboring bases. Each $R_{i}$ is supported on $s_{i},e_{i}]$ with uniformly spaced endpoints
$s_{i}\;=\;\frac{i-k}{G},\qquad e_{i}\;=\;s_{i}+\frac{k+1}{G},\quad i=0,\ldots,G+k-1.$
Inside its support, $R_{i}$ has a smooth bell shape (Figure 4):
$R_{i}(x)=\begin{cases}0,&x<s_{i},\\[2.0pt] \big((x-s_{i})(e_{i}-x)\big)^{2}\,\dfrac{16}{(e_{i}-s_{i})^{4}},&s_{i}\leq x\leq e_{i},\\[6.0pt] 0,&x>e_{i},\end{cases}$
which can be written equivalently using squared ReLUs:
$R_{i}(x)\;=\;\Big[\operatorname{ReLU}(e_{i}-x)\,\operatorname{ReLU}(x-s_{i})\Big]^{2}\cdot\frac{16}{(e_{i}-s_{i})^{4}},\qquad\operatorname{ReLU}(x)=\max(0,x).$
(a) Step-by-step construction of ReLU basis
(b) Complete set of normalized ReLU–KAN basis functions
Figure 4: (a) Step-by-step construction of the normalized ReLU-based local basis function from its constituent $\operatorname{ReLU}(e_{i}-x)$ and $\operatorname{ReLU}(x-s_{i})$ terms. (b) Complete set of normalized ReLU–KAN basis functions $R_{i}(x)$ with supports $s_{i},e_{i}] [$132].
Speed and a smoothness caveat. ReLU-KANs often train $5\text{--}20\times$ faster than spline variants in practice [132], but the squared–ReLU construction has limited smoothness, which can hinder PDE tasks requiring higher–order derivatives.
Higher–order ReLU KAN (HRKAN). So et al. [142] generalize the squared–ReLU basis by introducing local powers of ReLU. For a finite interval $s_{i},e_{i}]$, the basis function of order $m$ is
$v_{m,i}(x)\;=\;\Big[\operatorname{ReLU}(x-s_{i})\,\operatorname{ReLU}(e_{i}-x)\Big]^{\,m}\cdot\left(\frac{2}{\,e_{i}-s_{i}\,}\right)^{\!2m},$
where $\operatorname{ReLU}(x)=\max(0,x)$ and $m\in\mathbb{Z}_{+}$ controls the interior smoothness: $v_{m,i}\in C^{m-1}$. Larger $m$ produces lobes that are more peaked in the interior and decay more smoothly to zero at the boundaries, leading to higher continuity of derivatives and potentially better performance for PDEs that require smooth high-order derivatives.
Figure 5 compares individual basis functions $v_{m,i^{\star}}$ over $s_{i^{\star}},e_{i^{\star}}]$. Panel (a) shows the squared–ReLU case ($m=2$), which is $C^{1}$ but exhibits visible jumps in the second derivative at the boundaries. Panel (b) shows a higher–order ReLU with $m=4$, which is $C^{3}$ and decays smoothly to zero, eliminating derivative discontinuities. This single-basis view makes clear how increasing $m$ improves both interior smoothness and boundary regularity.
(a) Square–of–ReLU basis
(b) Higher–order ReLU basis
Figure 5: Symbolic basis on $s_{i^{\star}},e_{i^{\star}}]$: (a) $v_{2,i^{\star}}(x)$ with first and second derivatives; (b) $v_{4,i^{\star}}(x)$ with first and second derivatives. Observation: $v_{m,i}$ is globally $C^{m-1}$ (with the $m$-th derivative discontinuous at $s_{i^{\star}},e_{i^{\star}}$); hence $v_{4,i^{\star}}$ ($C^{3}$) offers smoother higher-order derivatives than $v_{2,i^{\star}}$ ($C^{1}$[142].
Figure 6 compares synthesized activations
$\varphi(x)\;=\;\sum_{i}c_{i}\,v_{m,i}(x),$
constructed from the two basis types using identical random coefficients ($G=3$, $k=1$) over support of $s_{i^{\star}},e_{i^{\star}}]=[-0.6,1]$. Because the coefficients are shared, differences between the dashed ($m=2$) and solid ($m=4$) curves arise solely from the change in $m$. The higher-order case produces narrower, more sharply peaked bumps with smoother interior derivatives, while the squared–ReLU case yields broader lobes with lower differentiability. Together with Figure 5, this demonstrates both the local and global effects of the order parameter $m$ in HRKAN.
Figure 6: Synthesized activation $\varphi(x)$ from ReLU-KAN bases using shared random coefficients. Dashed: squared–ReLU basis ($m=2$), Solid: higher–order ReLU basis ($m=4$). Higher order yields sharper, smoother lobes while preserving the same coefficient structure.
Both ReLU-KAN and HRKAN originally used fixed, uniformly spaced grids. To add adaptivity, Rigas et al. [127] [^19] https://github.com/srigas/jaxKAN define bases over a nonuniform, data–dependent grid $\mathcal{G}=\{x_{0},\ldots,x_{G}\}$. With neighborhood parameter $p$,
$s_{i}\;=\;\mathcal{G}[i]-\tfrac{1}{2}\big(\mathcal{G}[i+p]-\mathcal{G}[i-p]\big),\qquad e_{i}\;=\;2\mathcal{G}[i]-s_{i}.$
Adaptive widths and centers improve resolution in regions with singularities, steep gradients, or boundary layers; their jaxKAN implementation also uses resampling and loss reweighting for heterogeneous PDEs.
5.4 Jacobi and General Polynomials
Polynomial families constitute one of the most general and historically established classes of basis functions in KANs. Beyond splines and Chebyshev polynomials, a comprehensive benchmark by Seydi [147] [^20] https://github.com/seydi1370/Basis_Functions systematically evaluated eighteen distinct polynomial families as KAN activation functions, offering a unified comparison across orthogonal, recurrence-based, and rational constructions. Among all tested variants, the Gottlieb polynomial achieved the highest accuracy and stability metrics on the MNIST benchmark. The surveyed families can be grouped according to their mathematical origin:
  • Classical and General Orthogonal Polynomials: Charlier and Gottlieb (discrete orthogonal), Boas–Buck and Boubaker (generalized continuous families encompassing Hermite, Laguerre, and related forms).
  • Advanced Orthogonal Polynomials (Askey Scheme & Related): Askey–Wilson and Al–Salam–Carlitz ($q$-orthogonal series), Bannai–Ito (Racah generalization).
  • Recurrence-Based and Number-Theoretic Polynomials: Tribonacci, Tetranacci, Pentanacci, Hexanacci, Heptanacci, and Octanacci (generalized Fibonacci-type recurrences); Fermat, Vieta–Pell, and Narayana (number-theoretic families).
  • Rational Constructions: Padé approximants [137], explicitly adopted in the rational Jacobi network (rKAN).
  • Holomorphic Monomial Polynomials: Simple powers $(1,z,\dots,z^{P})$, where z=x+iyz=x+i\,y denotes the complex input, as used in PIHKAN [271].
Overall, Seydi’s benchmark revealed that KAN layers can flexibly host a wide variety of polynomial structures beyond splines, broadening the design space for architectures targeting spectral, combinatorial, or rational approximation behaviors. Among these, orthogonal and number-theoretic polynomials—particularly Gottlieb—exhibited the best numerical conditioning and convergence stability. These insights motivate the deeper exploration of Jacobi-type formulations discussed below, since Jacobi, Legendre, and Chebyshev families together form the classical orthogonal polynomial hierarchy within the Askey scheme.
Fractional Jacobi Basis (fKAN). To enhance smoothness, domain flexibility, and adaptivity, Aghaei introduced the Fractional KAN (fKAN) [136] [^21] https://github.com/alirezaafzalaghaei/fKAN, employing fractional-order Jacobi polynomials as trainable univariate maps:
$P_{n}^{(\alpha,\beta)}(z_{\gamma}),\qquad z_{\gamma}=\phi_{\gamma}(x)=2\,x^{\gamma}-1,\quad x\in[0,1],$
where $\alpha,\beta>-1$ are Jacobi exponents and $\gamma>0$ is a fractional warp parameter controlling the domain stretch. The basis remains orthogonal on the canonical interval [1,1][-1,1], and different $(\alpha,\beta)$ recover classical cases: Legendre for $(0,0)$, Chebyshev of the first kind for $(-\tfrac{1}{2},-\tfrac{1}{2})$, and Chebyshev of the second kind for $(\tfrac{1}{2},\tfrac{1}{2})$, as illustrated in Fig. 7.
Figure 7: Degree-$n=6$ polynomials on $z\in[-1,1]$: Chebyshev $T_{6}(z)$, Chebyshev $U_{6}(z)$, Legendre $P_{6}(z)$, and Jacobi $P_{6}^{(2.5,2.5)}(z)$. The Chebyshev–$T$ curve coincides with $T_{6}(z)$ in Fig. 3(a).
In fKAN, the univariate activation is expressed as
$\varphi(x)\;=\;P_{n}^{(\alpha,\beta)}\!\big(\,\phi_{\gamma}(\,x(r)\,)\,\big),$
where $r$ is the raw input, $x(r)\in[0,1]$ denotes a normalization map (linear, sigmoid, or $\tfrac{1}{2}(1+\tanh r)$), and $\phi_{\gamma}(x)=2x^{\gamma}-1$ applies fractional warping. The parameters $\alpha,\beta,\gamma$ are trainable, and positivity of $\alpha,\beta$ is enforced via ELU or sigmoid reparameterizations. This yields a globally smooth and tunable basis that performs effectively on regression and PDE benchmarks. Kashefi [75] implements fKAN using the $\tanh$-based normalization $x(r)=\tfrac{1}{2}(1+\tanh r)$ for stable input scaling.
Effect of Input Normalization. Implementations typically evaluate $P_{n}^{(\alpha,\beta)}(z_{\gamma})$ with $z_{\gamma}=2\,x(r)^{\gamma}-1$. Figure 8 illustrates how the normalization choice shapes the activation. Fixing $(\alpha,\beta)=(2.5,2.5)$, $n=2$, and $\gamma=0.1$, a linear mapping preserves dynamic range but produces steep edge slopes, whereas sigmoid and $\tfrac{1}{2}(1+\tanh r)$ compress the tails, improving conditioning. A smaller $\gamma<1$ increases the warp toward +1+1, further sharpening features near that end of the domain.
Figure 8: Degree-$n=2$ Jacobi polynomial with $(\alpha,\beta)=(2.5,2.5)$, $\gamma=0.1$: $\varphi(r)=P^{(\alpha,\beta)}_{2}\!\big(2\,x(r)^{\gamma}-1\big)$ under three normalizations $x(r)\in[0,1]$: linear $x=(r-r_{\min})/(r_{\max}-r_{\min})$ (dashed), sigmoid $x=\sigma(r)$, and $\tanh$-based $x=\tfrac{1}{2}(1+\tanh r)$. All curves are rescaled by $\varphi(r{=}1)$ for comparability.
Rational Extensions (rKAN). To further expand expressivity, Aghaei proposed the Rational KAN (rKAN) [137] [^22] https://github.com/alirezaafzalaghaei/rKAN, which generalizes fKAN by incorporating rational Jacobi compositions. Two representative variants are:
  • Padé–rKAN: a rational quotient of Jacobi expansions,
    $\varphi(x)=\frac{\sum_{i=0}^{p}w_{i}^{(P)}P_{i}^{(\alpha_{p},\beta_{p})}\!\big(\phi(\sigma(x))\big)}{\sum_{j=0}^{q}w_{j}^{(Q)}P_{j}^{(\alpha_{q},\beta_{q})}\!\big(\phi(\sigma(x))\big)},$
    where $w_{i}^{(P)}$ and $w_{j}^{(Q)}$ are trainable weights, and $\sigma(x)$ is an activation function.
  • Jacobi–rKAN: a rational domain warping approach,
    $\varphi(x)=P_{n}^{(\alpha,\beta)}\!\big(\phi(x;\iota)\big),$
    with $\phi(x;\iota)$ representing a rational transform parameterized by $\iota>0$ (often implemented using SoftPlus for positivity).
Both fKAN and rKAN exploit Jacobi polynomials with adaptive shape and domain parameters: fKAN emphasizes fractional-order smoothness and controlled warping, while rKAN introduces rational expressivity and sharper nonlinear transitions.
Classical Fixed-Degree Jacobi KANs. Several works retain fixed polynomial degrees ($\gamma=1$) and use only $\tanh$ input compression into [1,1][-1,1]. Each univariate activation is then represented as
$\varphi^{(l)}_{ij}(x)=\sum_{k=0}^{K}c^{(l)}_{k,ij}\,P_{k}^{(\alpha,\beta)}\!\big(\tanh(x)\big),$
with trainable coefficients $c^{(l)}_{k,ij}$ and often fixed $(\alpha,\beta)$. Kashefi [75] [^23] https://github.com/Ali-Stanford/KAN_PointNet_CFD reported that low-degree expansions ($K=2$) with Chebyshev-like parameters $\alpha=\beta=-0.5$ achieve an optimal balance between accuracy and numerical stability, suppressing Runge oscillations in inverse PDEs with sparse boundary data. Shukla et al. [42] applied the same formulation within a PIKAN framework for high–Reynolds cavity flow, using $(\alpha,\beta)=(1,1)$ and degrees K=3K=388 depending on the regime. While higher degrees increase computational cost, their Jacobi–based PIKANs matched or exceeded conventional PINNs in accuracy.
Building upon this foundation, Xiong et al. [78] proposed the Jacobian Orthogonal Polynomial–based KAN (J–PIKAN), enforcing explicit orthogonality among polynomial components to ensure stable layerwise synthesis and interpretable coefficient updates. Leveraging the three-term recurrence of Jacobi polynomials, J–PIKAN efficiently evaluates higher-degree expansions and maintains numerical stability. Using fluid-dynamics benchmarks, their study compared B-spline, Fourier, Hermite, Legendre, Chebyshev, and Jacobi bases within a unified physics-informed framework, identifying Jacobi bases with moderate parameters ($\alpha=\beta\approx 2$) as offering the best compromise between accuracy, conditioning, and convergence rate.
In parallel, Zhang et al. [79] [^24] https://github.com/zhang-zhuo001/Legend-KINN examined the Legendre specialization $(\alpha,\beta)=(0,0)$ within Jacobi-based KANs, demonstrating favorable conditioning and rapid convergence for physics-informed PDE solvers. A complementary direction is offered by the Adaptive PolyKAN [295], which treats the polynomial degree as a trainable parameter, enabling the model to adjust its functional complexity dynamically over the course of training.
Summary. Jacobi-based KANs form a versatile and mathematically rich subclass within polynomial KANs. Classical (fixed-degree) variants emphasize numerical stability and geometric fidelity, while fractional (fKAN) and rational (rKAN) extensions introduce adaptive domain warping and enhanced expressivity. Together, they provide a cohesive Jacobi toolkit—bridging spectral approximation, rational modeling, and physics-informed learning within the broader KAN framework.
5.5 Gaussian (RBF)
As researchers sought to improve the efficiency and performance of the original B–spline–based KANs, the Gaussian radial basis function (RBF) emerged as a powerful and computationally elegant alternative. Its smooth, localized, and infinitely differentiable nature makes it an excellent choice for representing univariate activation functions along the edges of a KAN.
Across recent studies [128, 59, 151, 129, 247, 125, 304, 314, 317], Gaussian–based activations have appeared in several formulations—ranging from fixed–grid hybrids to fully learnable and reflectional variants—each striking a different balance between stability, adaptivity, and efficiency.
Foundational Gaussian Layer. The fundamental univariate Gaussian basis function is defined as
$\varphi(x;c,\varepsilon)=\exp\!\left[-\Big(\tfrac{x-c}{\varepsilon}\Big)^{2}\right],$ (15)
where $c$ is the center and $\varepsilon>0$ is the width (shape parameter). A general univariate activation is then constructed as a linear combination
$\phi(x)=\sum_{i=0}^{G-1}w_{i}\,\varphi(x;g_{i},\varepsilon),$ (16)
with trainable coefficients $w_{i}$ and equispaced grid centers $g_{i}\in[a,b]$. This formulation replaces the piecewise polynomial behavior of splines with globally smooth Gaussian features while retaining locality and analytical gradients.
Hybrid Gaussian–Residual Formulation (Fixed Grid).
Gaussian RBFs are a fast, smooth alternative to B–splines for KAN activations. The key observation is that cubic B–splines can be closely approximated by scaled Gaussians. Li et al. [128] [^25] https://github.com/ZiyaoLi/fast-kan proposed FastKAN, replacing the cubic B–spline $\mathcal{B}_{3}(x-c_{i})$ with
$\mathcal{B}_{3}(x-c_{i})\;\approx\;\lambda\,\exp\!\left[-\Big(\tfrac{x-c_{i}}{\sigma}\Big)^{2}\right],$ (17)
where $\lambda\in\mathbb{R}$ is a scaling constant and $\sigma\approx h$ matches the grid spacing. As shown in Figure 9, this preserves spline-like locality and smoothness while reducing computational cost.
Figure 9: Comparison of cubic B–spline basis functions $\mathcal{B}_{3}(x)$ (non-extended) and Gaussian RBFs $\exp[-((x-c_{i})/\sigma)^{2}]$ with matched grid spacing and width.
Figure 10: Learned univariate activation $\varphi(x)$ synthesized from Gaussian RBF and cubic B–spline bases (Figure 9) using identical random coefficients.
Figure 10 demonstrates that both bases yield nearly identical activation profiles with only minor differences near peaks. Each basis decays smoothly near the boundaries, producing flat tails where support vanishes—confirming Gaussian RBFs as faithful, efficient surrogates for B–splines in KAN layers.
Building upon this foundation, Li [128] introduced the FastKAN model, where Gaussian basis functions are combined with a simple residual activation such as SiLU (Swish) to form a hybrid composite activation:
$\Phi(x)=\underbrace{\sum_{i=0}^{G-1}w_{i}\,\exp\!\Big[-\Big(\tfrac{x-g_{i}}{\varepsilon}\Big)^{2}\Big]}_{\text{RBF component}}+\,w_{b}\,\sigma(x),$ (18)
where $\sigma(x)$ is a base activation and $w_{b}\in\mathbb{R}$ modulates its contribution. The Gaussian component captures local variations, while the residual stabilizes optimization and accounts for global trends. Here, both the grid $\{g_{i}\}$ and bandwidth $\varepsilon=\tfrac{b-a}{G-1}$ are fixed, ensuring efficient and stable evaluation. This configuration, later adopted in LeanKAN [151] and RBFKAN [125], represents a strong, stable baseline for Gaussian–based KANs.
Pure RBF Layers with Learnable Parameters. A more flexible RBF formulation appears in the DeepOKAN architecture of Abueidda et al. [59] [^26] https://github.com/DiabAbu/DeepOKAN, where Gaussian bases are used in both branch and trunk networks. Each activation takes the form
$\varphi(x)=\exp\!\left[-\Big(\tfrac{x-g_{j}}{\beta}\Big)^{2}\right],$ (19)
with trainable centers $g_{j}$ and width $\beta>0$. The layer update is
$x_{l+1}=W_{l}\,\varphi(x_{l};G_{l},\beta),$ (20)
allowing the RBFs to migrate toward regions of high variation. Abueidda et al. show that RBF–KANs resolve high-frequency sinusoidal fields with markedly higher accuracy than MLPs in mechanics problems. A similar observation appears in the two-step DeepOKAN for bubble dynamics [152], where RBFs remain effective in the branch network while spline bases outperform them in the trunk for extremely high-frequency oscillations.
Beyond operator-learning settings, the Free-RBF-KAN model of Chiu et al. [317] introduces adaptive radial basis functions whose centers and widths evolve during training, improving efficiency and stability in general function-learning tasks.
Reflectional Gaussian Approximations (RSWAF). To further reduce computational cost, several works have replaced the exponential kernel with an algebraically similar Reflectional Switch Activation Function (RSWAF) proposed by Delis [129] [^27] https://github.com/AthanasiosDelis/faster-kan and extended in Bühler et al. [247]. This activation approximates the Gaussian bell using the hyperbolic–secant squared function:
$\varphi_{\text{RSWAF}}(x;c,h)=1-\tanh^{2}\!\Big(\tfrac{x-c}{h}\Big)=\mathrm{sech}^{2}\!\Big(\tfrac{x-c}{h}\Big),$ (21)
which preserves reflectional symmetry and smooth decay. Both the centers $c_{i}$ and widths $h_{i}$ may be fixed or trainable, controlled by configuration flags (e.g., train_grid, train_inv_denominator) in implementations such as FasterKAN [129]. The full activation is often expressed as
$\phi(x)=\sum_{i}w_{i}\,s_{i}\,\bigg[1-\tanh^{2}\!\Big(\tfrac{x-c_{i}}{h_{i}}\Big)\bigg],$ (22)
where $w_{i}$ are learnable amplitudes, $s_{i}$ are reference scaling factors, and both $c_{i}$ and $h_{i}$ evolve during training. This formulation yields a localized, computationally efficient surrogate that mimics Gaussian curvature while offering explicit control over both shape and location.
Hybrid and Unified Gaussian Extensions. Koenig et al. [151] [^28] https://github.com/DENG-MIT/LeanKAN proposed a hybrid activation that mixes RBFs with a base nonlinearity:
$\phi(x)\;=\;\sum_{i=1}^{N}w_{i}\,\varphi(x-c_{i})\;+\;w_{b}\,b(x),$ (23)
where $\varphi(\cdot)$ is a Gaussian RBF, $b(x)$ is a base activation (e.g., Swish or linear), and $w_{b}$ modulates its contribution. Related implementations such as RBFKAN [125] [^29] https://github.com/sidhu2690/RBF-KAN also apply simple min–max normalization,
$x_{\text{norm}}\;=\;\frac{x-x_{\min}}{x_{\max}-x_{\min}},$ (24)
to stabilize input scaling.
The BSRBF–KAN model [45] [^30] https://github.com/hoangthangta/BSRBF_KAN unifies B–splines and Gaussian RBFs within a single layer:
$\mathrm{BSRBF}(x)\;=\;w_{b}\,b(x)\;+\;w_{s}\big(\mathrm{BS}(x)+\mathrm{RBF}(x)\big),$ (25)
where $\mathrm{BS}(x)$ and $\mathrm{RBF}(x)$ are evaluated at identical grid locations, and $w_{b},w_{s}\in\mathbb{R}$. This configuration inherits the local compactness of B–splines and the global smoothness of Gaussian functions.
Complex–Valued Gaussian Extensions. In the complex domain, Wolff et al. [138] [^31] https://github.com/M-Wolff/CVKAN introduced the Complex–Valued KAN (CVKAN), extending FastKAN with concepts from complex–valued neural networks (CVNNs) [248]. The residual pathway uses a complex SiLU activation:
$b_{\mathbb{C}}(z)=\mathrm{CSiLU}(z)=\mathrm{SiLU}(\mathrm{Re}(z))+i\,\mathrm{SiLU}(\mathrm{Im}(z)),$ (26)
and the Gaussian RBF is generalized to complex inputs [249]:
$\mathrm{RBFC}(z)=\sum_{i=1}^{N}w_{i}\,\varphi(|z-c_{i}|),\qquad c_{i}\in\mathbb{C},\;w_{i}\in\mathbb{C}.$ (27)
The resulting activation combines both terms with a bias term:
$\phi(z)=w_{s}\,\mathrm{RBFC}(z)+w_{b}\,b_{\mathbb{C}}(z)+\beta,\qquad\beta\in\mathbb{C}.$ (28)
More recently, Che et al. [250] extended this formulation by introducing a ModELU–based CVKAN, replacing the split–type CSiLU with a modulus–preserving residual and learnable RBF shape parameters, supported by a formal complex–valued Kolmogorov–Arnold theorem.
Practical caveat: the shape parameter. Gaussian RBF performance is highly sensitive to the width parameter ($\varepsilon$), which governs both smoothness and conditioning [209, 210, 211, 212, 213, 214, 215, 216]. Even small changes can strongly impact accuracy and numerical stability [217, 218, 219]. From a kernel–theoretic perspective, this sensitivity reflects an uncertainty principle: larger shape (width) parameters produce flatter, globally correlated kernels that improve smoothness and accuracy but worsen conditioning, whereas smaller parameters yield well–conditioned, highly localized bases at the expense of approximation power [207, 208]. While heuristic tuning and learnable widths (as in FastKAN and DeepONet–RBF variants) offer partial remedies, robust and theoretically guided selection of this parameter remains an active research problem at the intersection of kernel methods and neural approximation.
Summary. Gaussian–based KANs form a versatile class of architectures bridging spline–based and kernel–theoretic perspectives. Starting from fixed equispaced Gaussian grids [128, 151], through adaptive RBF layers with learnable centers [59], to reflectional Gaussian and complex–valued extensions [129, 247, 138, 250], these models collectively demonstrate how Gaussian basis functions can serve as both efficient computational surrogates and expressive functional primitives within the KAN framework.
5.6 Fourier
Fourier features provide global, smooth, periodic expressivity that complements localized bases (e.g., B–splines, RBFs). Within the KAN family, Fourier–based KANs replace the univariate edge functions with harmonic expansions, retaining the KAN wiring while improving spectral resolution on tasks with strong periodic structure [143, 296, 311]. By contrast, Fourier networks such as FAN are not KANs: they embed cosine–sine transforms directly into each layer as a drop-in MLP replacement, targeting efficient periodicity modeling (often with reduced parameters/FLOPs) and stronger in-/out-of-domain generalization [146] [^32] https://github.com/YihongDong/FAN. We mention FAN here as a related spectral approach; our focus remains on Fourier features used inside KANs.
Standard FourierKAN. In the baseline FourierKAN architecture [134] [^33] https://github.com/GistNoesis/FourierKAN, each univariate mapping is expressed as a truncated Fourier series,
$\varphi(x)=\sum_{k=1}^{K}\big(a_{k}\cos(kx)+b_{k}\sin(kx)\big),$ (29)
with learnable coefficients $a_{k},b_{k}\in\mathbb{R}$ and cutoff frequency $K\in\mathbb{N}$. Cosines capture even–symmetric modes; sines capture odd–symmetric modes. The resulting $\varphi(x)$ is a flexible global harmonic mixture, well–suited to periodic or high–frequency structure. Guo et al. [169] adopt this classical truncated Fourier formulation, using fixed integer harmonics with trainable sine–cosine coefficients.
Random Fourier features. A limitation of (29) is the fixed set of frequencies $k=1,\ldots,K$. Zhang et al. [76] [^34] https://github.com/kolmogorovArnoldFourierNetwork/KAF propose the Kolmogorov–Arnold–Fourier Network (KAF), replacing fixed bases with a learnable spectral embedding via Random Fourier Features (RFF):
$\psi_{\mathrm{RFF}}(\mathbf{x})=\sqrt{\tfrac{2}{m}}\,\big[\,\cos(\mathbf{x}W+\mathbf{b}),\;\sin(\mathbf{x}W+\mathbf{b})\,\big]\in\mathbb{R}^{2m},$ (30)
where $W\in\mathbb{R}^{d_{\mathrm{in}}\times m}$ is a trainable frequency matrix and $\mathbf{b}\in\mathbb{R}^{m}$ is a learnable phase vector. KAF then blends this adaptive spectral block with a smooth baseline:
$\varphi(\mathbf{x})=\bm{\alpha}\cdot\mathrm{GELU}(\mathbf{x})\;+\;\bm{\beta}\cdot V\,\psi_{\mathrm{RFF}}(\mathbf{x}),$ (31)
with $\bm{\alpha},\bm{\beta}\in\mathbb{R}^{d_{\mathrm{in}}}$ and $V\in\mathbb{R}^{d_{\mathrm{in}}\times 2m}$. The GELU term captures low–frequency structure, while the RFF term adaptively models higher–frequency content throughout training. This continuous frequency control improves regression/classification where global and local patterns coexist [76].
Figure 11 compares a Gaussian (RBF) kernel slice with both random Fourier features (30) and a truncated Fourier expansion using the full cosine–sine basis. Settings match the code: one-dimensional grid $x\in[-3\varepsilon,3\varepsilon]$ with $\varepsilon=1.0$. The RFF map uses the unbiased scaling $1/\!\sqrt{m}$. With this setting, increasing $m$ makes the RFF curve visually coincide with the Gaussian slice—clearly showing how the two are connected in practice. The Fourier expansion (29), constructed from fixed integer harmonics $\{\cos(k\gamma x),\sin(k\gamma x)\}_{k=1}^{K}$, remains strictly periodic, so its kernel slice differs structurally from the Gaussian, in contrast to RFF where frequencies are sampled from a Gaussian distribution.
Figure 11: Gaussian (RBF) kernel slice compared with Random Fourier Features (RFF,  (30)) and a truncated Fourier expansion (29). RFF uses the unbiased scaling $1/\!\sqrt{m}$ with both $\cos$ and $\sin$ terms, so that the diagonal kernel value is close to 1. As $m$ increases, the RFF slice aligns with the Gaussian, while the Fourier expansion (fixed integer harmonics) retains its periodic structure.
In addition to classical Fourier-based designs, Jiang et al. [56] [^35] https://github.com/Jim137/qkan propose the quantum-inspired Kolmogorov–Arnold Network (QKAN), where each univariate map is realized by a data re-uploading variational quantum circuit. These circuits naturally generate Fourier-like expansions with exponentially scalable frequency modes, offering parameter-efficient alternatives to grid-based Fourier KANs while retaining the structured interpretability of classical KANs.
Summary. Fourier–based KANs—via deterministic truncations or randomized spectral embeddings—inject global harmonic structure into KAN layers. Compared with local polynomial bases such as B–splines, these global bases are particularly effective for smooth periodicities or problems with broad spectral support.
5.7 Wavelet
Wavelet-based KANs (Wav-KANs) extend the KAN framework by using localized wavelet bases, enabling multiscale representation and spatial adaptivity—especially useful for heterogeneous or hierarchical data. The architecture was introduced by Bozorgasl et al. [148] [^36] https://github.com/zavareh1/Wav-KAN and draws on both the Continuous Wavelet Transform (CWT) and Discrete Wavelet Transform (DWT).
In wavelet analysis, a signal $g(t)\in L^{2}(\mathbb{R})$ is decomposed via scaled/translated copies of a mother wavelet $\psi(t)$. The CWT is defined [224, 225] as
$C(s,\tau)\;=\;\int_{-\infty}^{\infty}g(t)\,\frac{1}{\sqrt{s}}\,\psi\!\left(\frac{t-\tau}{s}\right)\,dt,$ (32)
where $s>0$ is the scale and $\tau\in\mathbb{R}$ is the translation. The discrete analog used in signal processing is the DWT:
$a_{j}(k)\;=\;\sum_{n}g(n)\,\phi_{j,k}(n),\qquad d_{j}(k)\;=\;\sum_{n}g(n)\,\psi_{j,k}(n),$ (33)
with $\phi_{j,k}$ and $\psi_{j,k}$ the scaling and wavelet functions at resolution level $j$ and position $k$.
Wavelet activations in KAN layers. Wav-KAN embeds analytic wavelet functions directly into univariate edge functions. Each activation is parameterized by scale $s$ and translation $u$ [148, 297, 312]:
$\varphi(x)\;=\;\psi\!\left(\frac{x-u}{s}\right),$
where $\psi(t)$ is chosen from canonical families:
$\psi_{\text{mex}}(t)$ $=\tfrac{2}{\sqrt{3}\,\pi^{1/4}}(1-t^{2})\,e^{-t^{2}/2},$ (34)
$\psi_{\text{mor}}(t)$ $=\cos(\omega_{0}t)\,e^{-t^{2}/2},\quad\omega_{0}=5,$ (35)
$\psi_{\text{dog}}(t)$ $=-t\,e^{-t^{2}/2}.$ (36)
To compare expressivity across wavelet families, we form synchronized linear combinations
$\varphi(t)\;=\;\sum_{j=1}^{M}c_{j}\,\psi(t-\mu_{j}),$
using the same coefficients $\{c_{j}\}$ and centers $\{\mu_{j}\}$ for each $\psi$. Figure 12 shows three canonical wavelets (Mexican Hat, Morlet, DoG) centered at $t=0$. Using these bases, Figure 13 presents the synthesized activations $\varphi(t)$ built with identical $\{c_{j}\}$ and $\{\mu_{j}\}$. The Morlet activation exhibits higher-frequency oscillations, whereas Mexican Hat and DoG yield smoother bell-shaped profiles. Since coefficients are shared, differences arise solely from the intrinsic wavelet shapes.
Figure 12: Three wavelet bases centered at $t=0$: Mexican Hat, Morlet (with $\omega_{0}=5$), and Derivative of Gaussian (DoG). Each has localized support and distinct oscillatory behavior.
Figure 13: Synthesized activations $\varphi(t)=\sum_{j=1}^{M}c_{j}\psi(t-\mu_{j})$ using synchronized coefficients and centers across wavelet families. Differences in frequency and amplitude follow from each wavelet’s intrinsic shape.
Learning $s$ and $u$ per neuron allows Wav-KAN to adapt receptive fields and frequency resolution, aiding generalization in settings with spatially varying feature scales, such as image classification [148].
Patra et al. [200] apply wavelet activations—using a sine–Gaussian basis—to physics-informed training, tuning frequency response to obtain faster data-free convergence on coupled nonlinear PDEs. Meshir et al. [286] complement this with an NTK analysis, showing that the Morlet mother-wavelet frequency (or hidden-layer width) directly controls NTK eigenvalue decay, providing explicit spectral-bias mitigation and strong gains in physics-informed settings (Poisson/Heat/Wave/Helmholtz).
Applications further highlight the versatility of wavelet bases: Pratyush et al. [226] employ a DoG wavelet in CaLMPhosKAN for enhanced phosphorylation-site detection, while Seydi et al. [183] combine CWT and DWT for hyperspectral imaging, outperforming spline-KANs and MLPs. Wavelet–KAN blocks have also been integrated with attention mechanisms for dense prediction tasks such as medical image segmentation [313].
Summary. Wavelet-based KANs provide a flexible multiscale basis whose localized frequency control improves convergence and robustness across PDEs, bioinformatics, and high-dimensional imaging.
5.8 Finite-Basis (FBKANs)
Finite-Basis KANs, introduced by Howard et al. [131] [^37] https://github.com/pnnl/neuromancer/tree/feature/fbkans/examples/KANs, [^38] https://github.com/pnnl/neuromancer/blob/feature/fbkans/src/neuromancer/modules, extend KANs with a domain-decomposition strategy inspired by finite-basis PINNs (FBPINNs) [227, 228, 229, 230]. The key idea is to represent a global mapping as a sum of local subnetworks, each modulated by a smooth partition-of-unity (PoU) function. This yields scalability, robustness, and improved generalization for multiscale and physics-informed problems.
Let $\Omega\subset\mathbb{R}$ be covered by $L$ overlapping subdomains $\{\Omega_{j}\}_{j=1}^{L}$. Assign a smooth PoU weight $\omega_{j}(x)$ to each $\Omega_{j}$ such that [231, 232, 233]
$\mathrm{supp}\big(\omega_{j}\big)=\Omega_{j},\qquad\sum_{j=1}^{L}\omega_{j}(x)\equiv 1\ \ \forall x\in\Omega.$
Cosine windows are used to build unnormalized weights $\hat{\omega}_{j}$, which are then normalized:
$\omega_{j}(x)\;=\;\frac{\hat{\omega}_{j}(x)}{\sum_{k=1}^{L}\hat{\omega}_{k}(x)},$ (37)
with
$\hat{\omega}_{j}(x)\;=\;\begin{cases}1,&L=1,\\[4.0pt] \Big[1+\cos\!\big(\pi\tfrac{x-\mu_{j}}{\sigma_{j}}\big)\Big]^{2},&L>1,\end{cases}$ (38)
and
$\mu_{j}\;=\;\frac{l(j-1)}{L-1},\qquad\sigma_{j}\;=\;\frac{\delta\,l}{2(L-1)},\qquad l\;=\;\max(x)-\min(x),$
where $\delta>1$ is the overlap ratio controlling the window width.
Each subdomain $\Omega_{j}$ hosts a local KAN, $K_{j}(x;\bm{\theta}_{j})$, trained only on its region. The global predictor is a smooth PoU-weighted sum of local outputs:
$f(x)\;\approx\;\sum_{j=1}^{L}\omega_{j}(x)\,K_{j}(x;\bm{\theta}_{j}),$ (39)
where $\bm{\theta}_{j}$ are the local parameters. Within each local KAN, a univariate edge function $\varphi_{j}$ is expanded in cubic B-splines:
$\varphi_{j}(x)\;=\;\sum_{n=0}^{N_{j}-1}c_{j,n}\,B_{n}^{(3)}(x),$
with trainable coefficients $c_{j,n}\in\mathbb{R}$. The local spline grid is initialized from the effective support of $\omega_{j}$:
$a_{j}\;=\;\min\{x\mid\omega_{j}(x)>\varepsilon\},\quad b_{j}\;=\;\max\{x\mid\omega_{j}(x)>\varepsilon\},\quad\varepsilon\;=\;10^{-4}.$
To illustrate functional flexibility, Figure 14 shows, on [1,1][-1,1], (A) the PoU weights $\{\omega_{j}\}_{j=1}^{4}$ with $\sum_{j}\omega_{j}\equiv 1$, and (B) the PoU–weighted decomposition of a single cubic B-spline basis $B_{i}^{(3)}(x)$. The colored curves are the localized pieces $\omega_{j}(x)\,B_{i}^{(3)}(x)$; their sum coincides with the original basis up to numerical precision. Only windows overlapping the spline support contribute nonzero pieces, so the PoU localizes the representation without changing the underlying spline span.
Figure 14: Two-panel illustration of PoU–weighted B-spline decomposition on [1,1][-1,1] ($L=4$, overlap $\delta=1.5$). (A) Cosine-based PoU windows $\{\omega_{j}\}$ summing to one. (B) Decomposition of a cubic B-spline basis $B_{i}^{(3)}(x)$ into localized pieces $\omega_{j}(x)B_{i}^{(3)}(x)$; their sum (black) matches the original basis, confirming $B_{i}^{(3)}(x)=\sum_{j=1}^{L}\omega_{j}(x)\,B_{i}^{(3)}(x)$.
Summary. FBKANs are practical because they let you model different parts of the domain with specialized sub-KANs, which can be trained independently (and in parallel) to cut memory use and wall time. They work well with physics-informed losses, handle noisy data robustly, and maintain smooth transitions between subdomains without requiring hard boundary constraints. Moreover, because sub-KANs operate on smaller supports, each local kernel matrix is better conditioned—an important advantage for large-scale PDEs.
5.9 SincKAN
To model singularities, sharp gradients, and high-frequency transitions, Yu et al. [54] [^39] https://github.com/DUCH714/SincKAN introduced the Sinc-based KAN (SincKAN). The building block is the globally supported Sinc kernel
$\mathrm{Sinc}(x)=\frac{\sin x}{x},\qquad\mathrm{Sinc}(0):=1,$
a classical tool for bandlimited approximation.
A basic SincKAN activation expands a univariate function using uniformly shifted Sinc atoms:
$\varphi(x)=\sum_{i=-N}^{N}c_{i}\,\mathrm{Sinc}\!\left(\frac{\pi}{h}\,(x-ih)\right),$ (40)
where $h>0$ is the step size, $\{c_{i}\}$ are trainable c oefficients, and $N$ controls the truncation degree (total number of atoms $2N+1$).
Practical multi–step-size form. To avoid choosing a single “optimal” $h$ and to improve flexibility on finite domains, SincKAN mixes several step sizes and applies a normalized coordinate transform $\gamma^{-1}$:
$\varphi_{\mathrm{multi}}(x)=\sum_{j=1}^{M}\sum_{i=-N}^{N}c_{i,j}\,\mathrm{Sinc}\!\left(\frac{\pi}{h_{j}}\big(\gamma^{-1}(x)-ih_{j}\big)\right).$ (41)
Here $\{h_{j}\}_{j=1}^{M}$ are (learned or preset) step sizes, $M$ is the number of scales, $c_{i,j}$ are trainable coefficients, and $\gamma^{-1}$ maps the working interval to $\mathbb{R}$ while normalizing scale. Setting M=1M{=}1 and $\gamma^{-1}(x){=}x$ recovers (40).
Figures 15 and 16 show the building blocks and a synthesized activation using a fixed truncation degree $(2N+1=11)$ and spacing $h=1.0$ over $x\in[-10,10]$. Figure 15 displays the truncated Sinc dictionary $\{\mathrm{Sinc}(\tfrac{\pi}{h}(x-kh))\}_{k=-N}^{N}$, whose elements are globally supported and oscillatory with decaying sidelobes. These atoms are linearly combined to form the learned univariate map $\varphi(x)$.
Figure 15: Truncated Sinc basis functions used in SincKAN. Each kernel is centered at $x=kh$ with spacing $h=1.0$; the total number of shifts is $2N+1=11$. This dictionary underlies the learned univariate map $\varphi(x)$.
Figure 16 plots synchronized learned activations
$\varphi(x)=\sum_{i=-5}^{5}c_{i}\,\mathrm{Sinc}\!\big(\tfrac{\pi}{h}(x-ih)\big)$
with the same random coefficients $\{c_{i}\}$, while varying only the step size $h\in\{0.8,1.0,1.4\}$. The step size $h$ sets the spacing of Sinc centers and therefore the effective bandwidth: smaller $h$ $\Rightarrow$ denser centers, higher-frequency capacity, more oscillations; larger $h$ $\Rightarrow$ smoother, lower-frequency behavior. Because the coefficients are synchronized, differences across curves reflect only the choice of $h$.
Figure 16: Synchronized learned activations with a pure (non-windowed) Sinc basis (degree $2N+1=11$). The same coefficients $\{c_{i}\}$ are used for all curves; only the step size $h$ changes ($h\in\{0.8,1.0,1.4\}$). Smaller $h$ yields higher-frequency, more oscillatory structure, whereas larger $h$ produces smoother responses. Vertical dotted lines mark the Sinc centers for $h=1.0$.
For numerical stability, SincKAN applies a coordinate normalization step that maps inputs to a bounded interval (typically [1,1][-1,1]) before evaluating the Sinc expansion. This preprocessing, consistent with other KAN variants, improves conditioning and convergence across layers.
6 Accuracy Improvement
We group works by how they improve accuracy, independent of task (regression, PINN, DeepONet), basis choice, or efficiency tricks. In most cases the same mechanism transfers across tasks.
To orient the reader, we begin with a compact summary of accuracy–improving mechanisms for KANs. The table groups methods by mechanism and lists concise technique keywords with representative references. The remainder of this section expands on each group in turn, detailing when/why they help and practical choices and caveats.
6.1 Physics Constraints & Loss Design
Physics-informed learning, introduced by Raissi et al. [5], improves accuracy by embedding governing physical laws—PDE residuals and boundary/initial conditions—directly into the training objective.
A standard PINN-style loss combines three components:
[…]
where $\mathcal{L}_{\mathrm{data}}$ is the data-fidelity loss, $\mathcal{L}_{\mathrm{phys}}$ enforces PDE residuals, and $\mathcal{L}_{\mathrm{bc}}$ enforces boundary/initial conditions. The scalar weights $w_{u},w_{f},w_{b}>0$ determine the balance between data fidelity and physics consistency.
Building on this foundation, a wide range of studies have extended the basic PINN loss through specialized constraints and tailored formulations in cooperation with KAN architectures [1, 54, 60, 61, 62, 63, 57, 67, 68, 206, 201, 73, 142, 150, 64, 65, 66, 174, 251]. Among these, several works move beyond the conventional physics-based loss in (42) by embedding more advanced physics constraints and loss designs directly into the KAN architecture:
  • Structural & Conservation Laws. Enforce hard invariants such as Hamiltonian, quantum, algebraic, or geometric symmetries. Includes SPEL-KAN [257], KAN-ETS [255], DAE-KAN [270], and the symmetry-preserving GKSN architecture of Alesiani et al. [221].
  • Weak-Form & Variational Constraints. Use integral/variational formulations instead of pointwise PDE residuals. Includes WCN [251] and KANtrol [72].
  • Alternative Problem Formulations. Modify the optimization objective or physical variables. Includes AL-PKAN [71], AIVT [237], cPIKAN+RBA+EVM [42], KKAN+ssRBA [57], KAN-MHA [61], and holomorphic–potential PIHKAN [271].
  • Physics-Guided Discovery & Design. Physics-based losses used for inference, discovery, or design tasks. Includes PKAN [260] and physics-guided sensor design [252].
  • General PDE–Residual Models. Standard PINN-style residuals without special structural constraints. Includes KAN-PINN [60], Res-KAN [62], Anant-KAN [272], and PI-KAN [285].
In all these approaches, loss engineering—choosing residual forms, designing constraint terms, and applying adaptive weighting strategies—serves as a powerful inductive bias. By aligning the optimization process with the underlying physics, these designs improve accuracy, stability, and extrapolation capabilities across a broad range of PDE-driven problems. Beyond modifying the loss, several works instead adapt the sampling or grid itself, yielding comparable accuracy gains through geometric rather than variational control.
6.2 Adaptive Sampling & Grids
Adaptive sampling and grid refinement focus computational effort where the solution is most difficult to capture (steep gradients, shocks, oscillations). One simple approach is multi-resolution training, where a model is trained on alternating coarse and fine point clouds to accelerate convergence and lower computational cost [77]. More advanced methods, long used in finite elements [238, 239], adapt the model’s internal grid or sampling based on runtime dynamics for complex geometries [74, 65]. In stochastic systems, Chen et al. [251] introduced adaptive collocation sampling, selecting Gaussian test functions directly from observed data distributions.
Geometric/grid refinement with KANs. Actor et al. [51] treat spline knots as a learnable grid, which can be refined during training:
  1. 1.
    Multilevel refinement. Training begins on a coarse knot grid and periodically switches to finer grids. Since spline spaces are nested (coarse $\subset$ fine), the coarse solution transfers exactly to the refined grid. This preserves training progress, adds capacity only where needed, and achieves lower error at the same FLOPs compared with single-level training.
  2. 2.
    Free-knot adaptation. Knot locations are made trainable on $[a,b]$, with ordering enforced by a cumulative softmax:
    $t_{i}\;=\;\underbrace{a}_{\text{left endpoint}}\;+\;\underbrace{(b-a)}_{\text{interval length}}\sum_{j=1}^{i}\underbrace{\mathrm{softmax}(s)_{j}}_{\text{ordered weights}},\qquad s\in\mathbb{R}^{n},$
    ensuring that endpoints remain fixed ($t_{0}=a$, $t_{n}=b$). This naturally places knots in regions of high variation or non-smoothness. Related analyses of free-knot spline behavior in KANs have been carried out by Zheng et al. [316] [^40] https://github.com/IcurasLW/FR-KAN, who derive bounds on the number of spline knots and propose a Free-Knots KAN that improves training stability while reducing the number of trainable parameters.
On both regression and physics-informed tasks, multilevel schedules and free-knot adaptation consistently reduced errors relative to fixed grids and MLP baselines [51].
Grid–adaptive PIKANs. Rigas et al. [127, 275] extend this idea by adapting not only the spline grid but also the collocation points, so resolution follows the PDE residual. More recently, a dynamic grid-adaptation framework was proposed that treats knot allocation as a density-estimation problem governed by importance density functions, allowing curvature or training metrics to determine grid resolution during learning [315].
  1. 1.
    Grid extension with smooth state transition. Start with few spline intervals ($G$) and periodically increase $G$. At each extension, an adaptive optimizer state transition preserves momentum and avoids loss spikes, while the finer grid captures unresolved structure (Fig. 17(a)).
  2. 2.
    Residual-based adaptive resampling (RAD). Compute residuals on a dense probe set $S$ and convert them into sampling probabilities: larger residuals $\Rightarrow$ higher sampling chance. A new collocation set is then drawn accordingly (Fig. 17(b)).
Since KAN basis functions are tied to the knot grid, jointly adapting the grid and collocation points reduces projection error and concentrates model capacity in residual hotspots.
(a) Loss vs. epochs
(b) Residual-based adaptive sampling
Figure 17: Adaptive PIKAN training after [127]. (a) Loss vs. epochs: adaptive state transition smooths grid updates. (b) RAD concentrates collocation points where PDE residuals are largest.
6.3 Domain decomposition
Splitting a difficult global map into easier local ones and blending them smoothly often improves accuracy. For example, temporal subdomaining can sharpen NTK conditioning and convergence [69]. Finite-basis KANs [131], as discussed in Sec. 5.8, achieve this by decomposing $\Omega$ (or $\Omega\times[0,T]$) into $L$ overlapping subdomains $\{\Omega_{j}\}_{j=1}^{L}$ with a smooth partition of unity (PoU) $\{\omega_{j}\}_{j=1}^{L}$ such that
$\underbrace{\mathrm{supp}(\omega_{j})}_{\text{support of local weight}}=\Omega_{j},\qquad\underbrace{\sum_{j=1}^{L}\omega_{j}(x)\equiv 1}_{\text{partition of unity}}.$
A schematic is shown in Fig. 18.
A convenient 1D choice uses normalized “cosine bumps.” With domain length $l$ and overlap ratio $\delta>1$:
$\hat{\omega}_{j}(x)=\bigl[1+\cos\!\big(\pi(x-\mu_{j})/\sigma_{j}\big)\bigr]^{2},\quad\mu_{j}=\underbrace{\tfrac{l(j-1)}{L-1}}_{\text{center of bump}},\quad\sigma_{j}=\underbrace{\tfrac{\delta\,l/2}{L-1}}_{\text{width/overlap}},$
$\omega_{j}(x)=\underbrace{\frac{\hat{\omega}_{j}(x)}{\sum_{k=1}^{L}\hat{\omega}_{k}(x)}}_{\text{normalized local weight}}.$
In $d$ dimensions, the construction extends by taking products across coordinates.
The global surrogate is then the PoU blend of local KANs:
$f_{\Theta}(x)=\sum_{j=1}^{L}\underbrace{\omega_{j}(x)}_{\text{local weight}}\,\underbrace{K_{j}(x;\bm{\theta}_{j})}_{\text{local KAN}},$
where each $K_{j}$ is trained only on $\Omega_{j}$, using its own knot grid induced by $\omega_{j}$ (basis/grid choices per Sec. 5.8).
A basic error estimate illustrates the benefit:
$\|u-f_{\Theta}\|\;\leq\;\sum_{j=1}^{L}\underbrace{\|\omega_{j}\|_{\infty}}_{\text{PoU weight}}\,\underbrace{\|u-K_{j}\|_{\Omega_{j}}}_{\text{local error}},$
showing that reducing each local error reduces the global error. Increasing $L$ or the tunable overlap $\delta$ concentrates resolution where it is most needed. Each local KAN operates on a smaller, better-conditioned subproblem, allowing parallel training with improved numerical stability. In practice, per-subdomain grids adapt to local scales and consistently yield lower relative $L_{2}$ errors than single-domain KANs at the same model capacity [131].
Figure 18: Schematic of FBKAN domain decomposition and PoU blending. Overlapping subdomains $\Omega_{j}$ host lightweight local KANs $K_{j}$, smoothly combined via PoU weights $\omega_{j}(x)$. This enables parallel training and per-subdomain knot/hyperparameter choices (adapted from [131]).
6.4 Function decomposition
[…]
$K_{H}(x)=(1-\alpha)\,K_{\ell}\big(x,K_{L}(x)\big)+\alpha\,K_{nl}\big(x,K_{L}(x)\big),$
where $\alpha\in[0,1]$ is trainable, $K_{\ell}$ is the linear head, and $K_{nl}$ the nonlinear correction.
The HF loss balances accuracy, linearity preference, and regularization:
$\mathcal{L}_{HF}=\frac{1}{N_{HF}}\sum_{j=1}^{N_{HF}}\!\big(K_{H}(x_{j})-f_{H}(x_{j})\big)^{2}+\lambda_{\alpha}\,\alpha^{\,n}+w\,\|\Phi_{nl}\|_{2}^{2}.$
Here $\lambda_{\alpha}>0$ and $n\in\mathbb{N}$ control the preference for the linear block, while $w>0$ regularizes the nonlinear one.
The intuition is that the LF model captures the dominant structure, while the HF nonlinear head corrects only the residual. This reduces overfitting and stabilizes training when HF data are scarce. MFKAN achieved lower relative $L_{2}$ errors than single-fidelity KANs under the same HF budget [198].
The same fidelity split also benefits PIKANs, where $K_{L}$ serves as a coarse physics surrogate and the HF head is trained only on residual or boundary losses, making fidelity decomposition complementary to spatial domain decomposition (Sec. 6.3). Related coupled PIKAN strategies sequentially train subnetworks representing different subsystems to stabilize optimization [302], while multi-branch physics-informed KANs decompose interacting physical subsystems into parallel KAN branches with physics constraints [308].
The same fidelity split also benefits PIKANs, where $K_{L}$ serves as a coarse physics surrogate and the HF head is trained only on residual or boundary losses, making fidelity decomposition complementary to spatial domain decomposition (Sec. 6.3).
In a related direction, Jacob et al. [199] [^41] https://github.com/pnnl/spikans propose a separable PIKAN (SPIKAN), expressing the solution as a sum of products of 1D KAN factors—one per coordinate—before combining them into a global field. Bühler et al. [247] propose KAN-SR, a symbolic regression framework inspired by Udrescu et al. [246]. The method checks for separability or symmetry before modeling. If such structure exists, the function is decomposed into smaller subproblems, each fit with a compact KAN, and then recomposed into a closed-form symbolic expression. This divide–and–conquer strategy improves interpretability and robustness in noisy or high-dimensional settings.
[…]
6.5 Hybrid, Ensemble, and Data-Integrated Models
A complementary line of work boosts accuracy not by modifying KAN bases, but by combining KAN modules with other neural components—MLPs, attention blocks, or ensemble mechanisms. These hybrids exploit KAN’s structured 1D functional mappings while leveraging the expressive power of deep networks.
  • Function-combination KANs. FC-KAN combines multiple basis families (e.g., B-splines, wavelets, and radial basis functions) through element-wise operations such as sums, products, and concatenations, enabling richer functional representations within a single KAN layer [299] [^42] https://github.com/hoangthangta/FC_KAN.
  • MLP–KAN hybrids. These architectures use MLPs for broad feature extraction and KANs for structured refinement. Examples include the MLP–KAN mixture-of-experts of He et al. [46], the HPKM-PINN parallel fusion model of Xu et al. [63], and two-stage medical analysis pipelines that couple foundation-model segmentation with KAN-based classification [306].
  • Task-specific KAN replacements. Kundu et al. [273] (KANQAS) [^43] https://github.com/Aqasch/KANQAS_code replace the MLP in a Double Deep Q-Network with a KAN module for quantum architecture search.
  • Structured functional hybrids. KKAN [57] inserts small per-dimension MLPs to generate learned one-dimensional features $\Psi_{p,q}(x_{p})$, sums them across dimensions, and applies a simple outer basis expansion (Chebyshev, Legendre, sinusoidal, or RBF). This yields a pipeline
    $\text{inputs}\rightarrow\text{inner MLPs}\rightarrow\text{dimensionwise sum}\rightarrow\text{basis expansion}\rightarrow\text{output},$
    combining expressive learned features with a stable, low-bias basis. KKAN reports consistent gains over vanilla KANs and MLPs across regression, operator learning, and PDE benchmarks (Fig 20).
  • Symbolic-regression hybrids. Howard et al. [3] propose SINDy-KANs, which combine KANs with SINDy-style sparse regression to identify parsimonious symbolic representations at the level of the learned activation functions. This hybrid design targets equation discovery and dynamical-system identification, showing that KANs can be coupled not only with neural modules, but also with sparse scientific model-discovery frameworks.
Figure 20: KKAN architecture (adapted from [57]). Each input coordinate $x_{p}$ These are summed across dimensions to form $\xi_{m}$, which are expanded via basis functions in the outer block. A final summation yields the output. Bases may include Chebyshev/Legendre polynomials, sine series, or RBFs. This “MLP inside + basis outside” layout mirrors the Kolmogorov–Arnold recipe.
6.6 Sequence and Attention Hybrids
KANs are increasingly combined with sequence encoders and attention mechanisms to capture long-range temporal structure, global context, or multi-scale interactions. Most architectures fall into three categories: (i) attention-enhanced KANs, (ii) sequence-to-field encoder–decoder hybrids, and (iii) generative or data-augmentation models that stabilize training.
  • Attention-enhanced KANs. These models insert KAN blocks into attention layers to improve multi-scale representation. Includes KAN–MHA [61], FlashKAT (Group-Rational KAN in Transformers) [202], Attention-KAN-PINN for battery SOH forecasting [223], and wavelet-attention hybrids such as DWSen [312].
[…]
  • Generative / data-augmentation hybrids. These integrate KANs with GAN or adaptive-activation mechanisms to improve learning from sparse or noisy data. Includes AAKAN–WGAN [204], which uses a Wasserstein GAN to augment training data.
Across all designs, the auxiliary module supplies memory, context, or synthetic data, while the KAN block provides a compact and interpretable functional mapping at the output stage.
6.7 Discontinuities and Sharp Gradients
KAN variants approach discontinuities, shocks, and steep gradients through three main mechanisms: (i) using bases that naturally represent non-smooth features, (ii) attention or residual reweighting that concentrates learning where gradients are large, and (iii) physics-driven stabilization for shock-dominated PDEs.
  • Non-smooth or localized bases. Designed to capture kinks, singularities, or asymptotic regimes. Includes SincKAN [54], rKAN [137], and KINN for stress singularities [74].
  • Attention and adaptive residuals. Focuses computational effort on regions with steep gradients. Includes KAN–MHA [61], KKAN+ssRBA [57], and AIVT [237].
  • Physics-driven shock stabilization. Includes cPIKAN+EVM/RBA [42], Buckley–Leverett KANs [66], EPi–cKAN for elasto-plastic transitions [67],and physics-informed KAN solvers for the Vlasov–Poisson system in cosmological structure formation [301].
Overall, these approaches mitigate non-smoothness by introducing localized bases (Sinc, rational), adaptive residual focusing, or physics-aware stabilization that controls oscillations near shocks.
Edge-aware KANs. Lei et al. [65] embed an explicit edge activation inside each KAN layer, enabling sharp jumps while keeping the rest of the representation smooth:
$\psi(x)=w_{t}\,\tanh\!\big(\alpha(x-\beta)\big)+w_{s}\sum_{i}c_{i}\,S_{i}(x),$
where $\alpha,\beta,w_{t},w_{s},c_{i}\in\mathbb{R}$ and $S_{i}(\cdot)$ are B-spline basis functions.
The $\tanh$ term places and sharpens the discontinuity, while the spline component fits the surrounding smooth regions. Incorporated into a PIKAN framework, this mechanism reduces spectral bias and yields more stable representations of steep fronts, a persistent difficulty for both MLPs and classical spline-based KANs.
6.8 Optimization & Adaptive Training
Training KANs—especially in physics-informed or data-scarce settings—relies heavily on optimization choices. Two themes dominate: (i) hybrid optimization schedules and (ii) Bayesian/probabilistic methods.
[…]
Bayesian and probabilistic methods. Probabilistic formulations are increasingly used to calibrate KANs or quantify prediction uncertainty. Lin [197] employed Bayesian optimization to tune RBF–KAN hyperparameters. Giroux and Fanelli [150] [^44] https://github.com/wmdataphys/Bayesian-HR-KAN introduced variational Bayesian inference for Higher–Order ReLU–KANs, placing priors on both weights and basis parameters; KL regularization improved robustness on stochastic Poisson and Helmholtz problems. Hassan [170] proposed a spline-based Bayesian–KAN, assigning Gaussian priors $c_{n}\!\sim\!\mathcal{N}(\mu_{n},\sigma_{n}^{2})$ to the activation coefficients for uncertainty quantification across layers.
A complementary direction is the probabilistic KAN of Polar and Poluektov [274] [^45] https://github.com/andrewpolar/pkan, which replaces Bayesian priors with an ensemble built by divisive data re-sorting. This approach learns multiple KANs over recursively refined subsets of the data, capturing full input-dependent output distributions—including multi-modality—while remaining computationally lightweight and highly parallelizable.
Initialization and depth-stabilized training. Beyond optimizer choice and probabilistic tuning, Rigas et al. [278] [^46] https://github.com/srigas/RGA-KANs show that deep Chebyshev-based KANs (cPIKANs) fail primarily due to variance blow-up across layers. They introduce a basis-agnostic Glorot-like initialization that preserves forward/backward variance and a Residual–Gated Adaptive block that stabilizes deep gradient flow.
For spline KANs, Rigas et al. [291] [^47] https://github.com/srigas/KAN_Initialization_Schemes demonstrate that Glorot-style and power-law initializations substantially improve NTK conditioning and convergence across regression and PDE benchmarks, underscoring initialization as a key design lever for scalable KAN depth. Related efficiency improvements can also be achieved through feature enrichment strategies such as FEKAN [309], which augment the input representation before the KAN mapping to accelerate convergence without modifying the underlying architecture.
Advanced quasi-Newton methods. Kiyani et al. [279] show that self-scaled BFGS/Broyden optimizers markedly outperform Adam and standard BFGS on stiff PDEs. Their results reveal two effects: the PDE residual creates a highly “crinkled” loss landscape where first-order methods stall, and stronger optimizers cut this optimization error so effectively that smaller networks reach equal or better accuracy. One-liner: self-scaled quasi-Newton methods tame the PDE landscape and enable accurate, compact PIKANs. A related line of work by Poluektov and Polar  [280] [^48] https://github.com/andrewpolar/det5 replaces gradient training entirely with a Newton–Kaczmarz update on the spline/Gaussian coefficients, yielding a solver that is less sensitive to initialization and efficient for large datasets.
Variational basis-size learning. Alesiani et al. [220] propose INFINITYKAN, which treats the number of basis functions in each univariate map as a latent variable optimized via mean-field variational inference. Rather than fixing basis width, Poisson-distributed variables $\lambda_{\ell}$ determine layer-wise cardinality, allowing the representational capacity to expand or contract during training. Experiments show basis-size selection is a task-dependent and learnable component of KAN optimization.
These works highlight the role of probabilistic reasoning for calibration, interpretability, and reliable prediction in settings with randomness or limited data.
7 Efficiency Improvement
Efficiency gains in KANs arise from two complementary directions: (i) exploiting hardware parallelism and compiler support, and (ii) reducing algebraic and basis-level complexity.
7.1 Parallelism, GPU, and JAX engineering
A major line of work accelerates KANs by making their core operations GPU-friendly. Iterative B-spline evaluations are replaced by non-iterative ReLU-power activations [53] or reformulated as CUDA-optimized matrix kernels [132, 142]. Model-level restructuring further boosts throughput: dual-matrix compression with Random Fourier Features [76], custom backward kernels to reduce memory stalls [202], and GPU-accelerated B-spline evaluation [133]. Architectural parallelism also plays a role, including multi-branch KAN–MLP hybrids [63] and domain-decomposition designs for PINNs [42, 131, 199]. At the framework level, JAX implementations [70, 127] leverage jit/XLA and optimized autodiff to deliver substantial end-to-end speedups, while specialized hardware designs such as computing-in-memory architectures using Gaussian-like memory cells further improve energy-efficient KAN inference [298].
7.2 Matrix optimization and parameter-efficient bases
A second strategy reduces computational cost by simplifying basis functions and matrix operations. Replacing spline bases with faster alternatives—ReLU-power [53, 132, 142], Chebyshev/Jacobi/Legendre polynomials [42, 206, 67, 68, 74], RBFs [197, 49], or wavelets [200]—lowers FLOPs and parameter counts while maintaining approximation quality. Additional savings come from structural compression: dual-matrix fusion with RFFs [76], sparsity-promoting regularizers [62], hierarchical knot refinement [51], and topology search via Differential Evolution [48]. Operator-aware designs further reduce cost by mixing spectral and spatial representations to keep matvecs near-diagonal [253], limiting mode mixing and memory traffic on variable-coefficient PDEs. Anant-Net [272] [^49] https://github.com/ParamIntelligence/Anant-Net extends these ideas with dimension-wise tensor-product sweeps and selective differentiation, both of which reduce high-dimensional autodiff complexity and are readily transferable to KANs. In a related direction, lookup multivariate KANs (lmKANs) replace univariate inner functions with multivariate spline lookup tables and CUDA-efficient evaluation, providing a parameter-rich yet computationally light alternative to linear layers [284] [^50] https://github.com/schwallergroup/lmkan.
Together, these techniques make KANs increasingly competitive in large- and high-dimensional settings by aligning basis, algebra, and implementation with modern hardware.
8 Sparsity & Regularization
Regularization plays a central role in improving the stability and generalization of KANs, especially in noisy, high-dimensional, or ill-posed regimes. Existing approaches fall into four broad families: sparsity penalties, Lipschitz/complexity control, weight decay, and implicit or dropout-style regularizers.
L1 sparsity and entropy balancing. Liu et al. [1] introduce an activation-level $\ell_{1}$ penalty,
$\|\varphi_{ij}\|_{1}=\tfrac{1}{N_{p}}\!\sum_{s=1}^{N_{p}}\!|\varphi_{ij}(x^{(s)})|,\qquad\|\Phi\|_{1}=\sum_{i,j}\|\varphi_{ij}\|_{1},$ (43)
supplemented by an entropy balancing term,
$S(\Phi)=-\sum_{i,j}\frac{\|\varphi_{ij}\|_{1}}{\|\Phi\|_{1}}\log\!\Big(\frac{\|\varphi_{ij}\|_{1}}{\|\Phi\|_{1}}\Big),$ (44)
leading to the regularized objective
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{pred}}+\lambda\big(\mu_{1}\!\sum_{\ell}\|\Phi_{\ell}\|_{1}+\mu_{2}\!\sum_{\ell}S(\Phi_{\ell})\big).$ (45)
Variants such as EfficientKAN [126] [^51] https://github.com/Blealtan/efficient-kan and DeepKAN [124] [^52] https://github.com/sidhu2690/Deep-KAN apply $\ell_{1}$ directly to weights; the same scheme appears in symbolic regression [58] and PDE learning [62]. For Neural ODEs, Koenig et al. [49] use a layerwise $\ell_{1}$ followed by pruning,
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{fit}}+\lambda_{\text{L1}}\|\theta\|_{1},$ (46)
eliminating redundant connections without accuracy loss. Application-driven sparsity can also arise from the data representation itself; for example, SP-KAN employs sparse sinusoidal perception layers aligned with the intrinsic sparsity of infrared targets [296] [^53] https://github.com/xdFai.
Structured sparsity for symbolic regression. Bühler et al. [247] propose a composite regularizer for KAN-SR combining: (i) a magnitude penalty to suppress unused inputs, (ii) row/column entropies to localize attention, and (iii) an $\ell_{1}$ penalty on linear weights. This promotes compact, interpretable expressions while preventing the overspecification typical of symbolic regression networks.
Bayesian sparsity. Zou and Yan [259] introduce Probabilistic KANs (PKAN) and their sparse counterpart SSKAN, using Gaussian-process priors and spike–and–slab priors, respectively. These provide principled sparsity and uncertainty quantification, forming a Bayesian alternative to $\ell_{1}$-based pruning.
Lipschitz-driven regularization. Li et al. [290] show that KAN generalization is governed by a Lipschitz complexity depending on per-layer Lipschitz constants and mixed tensor norms. Their LipKAN architecture inserts 1-Lipschitz layers and replaces $\ell_{1}$ with a theoretically motivated $\ell_{1.5}$ penalty, providing tighter complexity control and improved generalization on complex distributions.
Low-rank RKHS regularization. Zhang and Zhou [289] establish generalization bounds for KANs whose activations lie in a low-rank RKHS (e.g., Wendland functions [122]), with complexity scaling in the activation rank rather than width. This motivates LoRA-style rank-constrained fine-tuning of pre-trained KAN activations: maintaining expressivity while regularizing model capacity.
L2 weight decay and extensions. Shen et al. [204] incorporate $\ell_{2}$ weight decay,
$\mathcal{L}_{\mathrm{base}}=\mathcal{L}_{\mathrm{pred}}+\lambda_{1}\mathcal{L}_{\mathrm{L2}},$ (47)
along with temporal smoothness and mutual-information terms,
$\mathcal{L}_{\mathrm{AAKAN}}=\mathcal{L}_{\mathrm{base}}+\lambda_{2}\mathcal{L}_{\mathrm{time}}+\lambda_{3}\mathcal{L}_{\mathrm{MI}},$ (48)
to stabilize dynamics. Small $\ell_{2}$ penalties are similarly used in physics-informed KANs [42, 237] to prevent overfitting while preserving optimization stability.
Implicit and dropout-style regularizers. Implicit regularization arises when activations themselves bound outputs. Daryakenari et al. [70] show that nested activations such as $\tanh(\tanh(x))$ naturally confine outputs, smoothing gradients in Chebyshev-based KANs. Altarabichi et al. [240] propose DropKAN, which masks post-activation outputs rather than inputs, injecting noise directly into spline evaluations and offering more stable regularization than standard dropout.
9 Convergence and Scaling Laws
Classical approximation theorems guarantee that neural networks can approximate any continuous function, but provide little insight into convergence rates. Recent KAN studies fill this gap by deriving Sobolev-space error bounds, analyzing gradient-flow dynamics via NTK theory [241, 12], and documenting empirical scaling laws [196].
9.1 Theoretical Approximation Rates and Empirical Scaling Laws
Sobolev and Besov rates. Wang [58] showed that deep spline-based KANs achieve a Sobolev approximation rate that doubles the classical parameter-rate $\mathcal{O}(P^{-s/d})$ of standard neural networks and kernel approximators, yielding $\mathcal{O}(P^{-2s/d})$ convergence [242, 243]. This accelerated rate reflects the interaction between the compositional structure of KANs and the expressive power of high-order spline bases.
Formally, the result is obtained by combining the exact reparametrization of $\mathrm{ReLU}^{k}$ networks by KANs (Theorem 5) with sharp approximation bounds for very deep piecewise-polynomial MLPs on Sobolev spaces [58]. Through this equivalence, established approximation theory for deep $\mathrm{ReLU}^{k}$ networks transfers directly to KAN architectures, leading to the following corollary.
Corollary 1 (Sobolev approximation rate for deep KANs [58])
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with smooth boundary and $s>0$. There exists a fixed width $W_{0}=W_{0}(d)$ such that for any $f\in W^{s}(\Omega)$ and any depth $L\geq 1$, there exists a KAN $g$ with width $W_{0}$, depth $L$, grid size $G=2$, and $k$-th order B-spline basis functions satisfying
$\|f-g\|_{L^{p}(\Omega)}\;\leq\;C\,L^{-2s/d},$
where $C$ is a constant independent of $L$.
Kratsios et al. [287] extend this result to Besov spaces ([288]), showing that spline-based Res-KANs achieve the optimal Besov approximation rate on Lipschitz and fractal sets with dimension-free sample complexity. Complementing these approximation-centric results, Liu et al. [149] establish the first minimax statistical rates for KAN regression: spline-based additive and hybrid KANs achieve the optimal
$\mathcal{O}\!\left(n^{-2r/(2r+1)}\right)$
for Sobolev-$r$ functions, together with an optimal knot-scaling rule $k_{n}\asymp n^{1/(2r+1)}$. These three perspectives jointly show that KANs match the best-known approximation and estimation rates of classical spline and sieve methods while retaining neural-network–style compositional structure.
Empirical scaling. Liu [1] observed that KAN RMSE follows a power law
$\ell\propto N^{-\alpha},$
with task-dependent $\alpha>0$. Increasing basis resolution (e.g., spline grid size $G$) consistently improves accuracy more than widening layers, echoing the theoretical depth benefit in [58] and the statistical spline-resolution rule in [149]. Together, these findings indicate that KANs achieve fast decay in both theory (Sobolev/Besov/minimax rates) and practice (empirical power laws), though [287] emphasizes that many of these gains reflect favorable optimization dynamics rather than asymptotic expressivity alone.
9.2 Spectral Bias and Frequency Learning Behavior
A model has spectral bias if low frequencies are learned before high ones [263, 264, 267]. In NTK terms, this corresponds to steep eigenvalue decay. Flatter spectra indicate more uniform frequency learning.
Reduced spectral bias in KANs. Wang [58] showed via Hessian–eigenvalue analysis that shallow KANs possess a well-conditioned landscape, enabling fast learning of oscillatory modes. Grid extension further widens the representable frequency band. Their experiments show KANs capturing high-frequency signals in a few iterations, unlike MLPs.
Farea [44] confirmed this in PDE settings: KANs and learnable-basis PINNs (e.g., Fourier/Jacobi) maintain flatter NTK eigenvalue spectra, indicating weaker spectral bias [268, 269]. However, they also showed that broader spectra can increase curvature, worsen conditioning, and destabilize training [266]. Wang [58] similarly noted that overly fine spline grids may amplify noise sensitivity, whereas modest coarsening can improve generalization.
Together, these results show that KANs learn high-frequency features more uniformly [58, 44], but at the cost of sharper loss landscapes—consistent with kernel-theoretic ill-conditioning phenomena [207, 208]. Basis choice must therefore balance spectral reach (e.g., Fourier-like) and stability (e.g., splines).
9.3 NTK-Based Convergence and Physics-Informed Extensions
Gradient-flow convergence. Gao [174] analyzed two-layer KANs under gradient flow, showing that the residual $s(t)$ evolves via
$\frac{d}{dt}s(t)=-G\,s(t),$
with $G=D^{\top}D$ the NTK Gram matrix. If $G\succ 0$, the loss decays exponentially:
$\mathcal{L}(t)\leq\Big(1-\eta\tfrac{\lambda_{\min}}{2}\Big)^{t}\mathcal{L}(0),$
giving a rigorous rate under appropriate learning rates. The same holds in expectation for SGD. Intuitively, a better-conditioned NTK (i.e., a larger and more balanced spectrum of eigenvalues) implies that gradients propagate more uniformly across training directions, leading to faster and more stable convergence in practice.
Physics-informed settings. For PDE residual losses,
$\mathcal{L}_{\mathrm{PDE}}(t)\;\leq\;\Big(1-\eta\,\tfrac{\lambda_{\min}}{2}\Big)^{t}\mathcal{L}_{\mathrm{PDE}}(0),$
but $\lambda_{\min}$ depends strongly on operator stiffness [174, 69]. Faroughi [69] showed that Chebyshev-based cKAN/cPIKAN models maintain better NTK conditioning and converge faster on diffusion, Helmholtz, Allen–Cahn, and beam PDEs. They also demonstrated that temporal decomposition and hybrid Adam–BFGS schedules improve conditioning, aligning with multi-stage results in [68, 67]. Initialization can also influence NTK conditioning: Rigas et al. [291] [^54] https://github.com/srigas/KAN_Initialization_Schemes show that variance-preserving Glorot-style and power-law schemes produce more stable initial NTKs for spline-KANs, leading to faster convergence.
9.4 Practical Trade-offs and Convergence Guarantees
KAN convergence is tightly controlled by the conditioning of the NTK Gram matrix: structured bases often enlarge the smallest eigenvalue $\lambda_{\min}$ relative to MLPs [174, 58], accelerating early optimization. Yet the same expressivity that broadens spectral coverage also increases curvature in the loss landscape, making training sensitive to basis resolution and learning-rate choices [44]. Effective practice therefore requires moderating basis richness to preserve numerical stability.
Physics-informed models highlight this balance. Chebyshev-KAN and cPIKAN architectures benefit from well-spread NTK spectra, producing faster PDE residual decay and more stable training [69, 68, 67]. But these gains diminish when grids are refined too aggressively or dimensionality grows, where operators become both costly to evaluate and harder to condition [47]. The result is a gradual trade-off between spectral reach and computational/numerical stability.
Generalization theory reaches the same conclusion from a different angle. Zhang and Zhou [289] show that KAN complexity is governed by $\ell_{1}$ coefficient norms and per-layer Lipschitz constants—precisely the terms that inflate when bases become overly sharp or high-frequency. Li et al. [290] formalize this perspective through Lipschitz complexity, deriving complexity-dependent generalization bounds and introducing LipKANs, which reduce Lipschitz growth via Lipschitz-controlled layers and $L^{1.5}$ regularization. Their results empirically demonstrate that controlling Lipschitz complexity mirrors the conditioning behavior seen in NTK dynamics: sharper bases yield higher curvature and poorer generalization.
Taken together, NTK convergence guarantees [174], approximation-rate theory [58, 289], Lipschitz-based generalization analyses [290], and empirical scaling laws [1] support a unified conclusion: KANs are expressive and parameter-efficient, but their practical performance depends on managing the expressivity–stability trade-off. The most effective models achieve wide spectral coverage while keeping curvature, conditioning, and coefficient growth under control [58, 44, 69].
10 Practical “Choose–Your–KAN” Guide
When to prefer KANs over MLPs. It is important to note that highly optimized MLP implementations remain the preferred choice in many high-throughput production settings where inference latency and hardware efficiency are the primary concerns. KANs are most advantageous when the learning problem benefits from explicit control over the underlying function space. In practice, KANs are preferable to standard MLPs when (i) data are sparse or expensive and fast convergence is required, (ii) a shallow architecture is desired for interpretability or stability, or (iii) prior knowledge about the solution structure is available. Examples include globally smooth or oscillatory functions (Chebyshev/Fourier bases), localized or multi-scale features (splines or wavelets), and nonsmooth or discontinuous behavior (Sinc- or rational-based KANs). In contrast, when large amounts of data are available and no structural prior is known, generic MLPs may remain competitive due to their simplicity and hardware efficiency.
The correct configuration of KANs depends on (i) the structure of the target function (smooth, oscillatory, discontinuous, periodic, multi-scale), (ii) computational constraints (time, memory, precision), and (iii) desired properties (accuracy, interpretability). A clear seven-step recipe addressing these aspects is given below and is also presented in algorithmic form as Algorithm 1, synthesizing Sections 59.
Step 1 — Begin with a stable default. For general regression/classification, start with a cubic B-spline KAN on a uniform grid. This setup is mature, stable, and well-conditioned [1, 2]. For PDE or boundary-sensitive tasks, enable grid extension and, if needed, apply post-training refinement (Sec. 5.1; Fig. 2; [2]).
Step 2 — Select a basis, matched to the problem structure. Choose the basis family whose functional properties align with the anticipated solution:
  • Globally smooth / oscillatory: Chebyshev (ChebyKAN) with per-layer $\tanh$ normalization [205, 139, 54, 69].
  • Controllable smoothness or boundary shaping: Jacobi (fKAN, Jacobi-KAN) with fractional/domain parameters [136, 75].
  • Discontinuities, fronts, boundary layers: SincKAN (bandlimited) [54], rational rKAN [137], or DKAN (tanh-jump + smooth background) [65].
  • Periodic structure: FourierKAN or KAF (learnable RFF) [143, 76].
  • Local, multi-scale, or bursty features: Wavelets (Wav-KAN) [148, 200].
Step 3 — If speed matters, replace heavy components. For wall-clock efficiency, modify the basis or architecture:
  • Fast local bases: ReLU-KAN/HRKAN ($\mathrm{ReLU}^{m}$[132, 142]; or Gaussian RBF variants (FastKAN) [128].
  • Parallelism and compilers: JAX jit/XLA, fused CUDA kernels, multi-branch or domain-decomposition architectures [127, 70, 131].
  • Parameter economy: Prefer orthogonal polynomials or RBFs; prune using $\ell_{1}+$entropy penalties [42, 62, 1].
Step 4 — Stabilize via normalization, grid design, and Lipschitz control. Choose conditioning mechanisms suited to the basis, and reduce Lipschitz complexity where possible:
  • Chebyshev/Jacobi → apply $\tanh$ normalization to prevent blow-up of high-order modes [54, 70].
  • RBF/Wavelet → use min–max or sigmoid/tanh scaling to keep oscillatory activations bounded [247, 225].
  • Splines → employ extended grids or free-knot refinement near sharp gradients [51, 127].
  • All bases optionally insert $1$-Lipschitz “Lip” layers (tanh/sigmoid/ReLU) between $\Sigma$ and linear maps to reduce Lipschitz complexity and improve generalization [290].
Step 5 — Add physics and adaptive sampling only when needed. For PINNs/PIKANs, form
$\mathcal{L}=w_{u}\mathcal{L}_{\mathrm{data}}+w_{f}\mathcal{L}_{\mathrm{phys}}+w_{b}\mathcal{L}_{\mathrm{bc}},$
and enhance only when necessary:
  • Augmented Lagrangian or residual-based attention for conflicting constraints [5, 71, 57].
  • Residual-adaptive sampling (RAD) for error concentration [127].
[…]
Step 6 — Use the standard optimizer progression. Warm up with Adam/RAdam, then refine with (L-)BFGS. Activate mixed precision and gradient clipping for stiff PDEs [70, 68, 52, 69].
Step 7 — Increase capacity targetedly, not uniformly. KAN error typically follows $\ell\propto P^{-\alpha}$, $\alpha>0$ (Sec. 9; [1, 58]). The most effective scaling strategies are:
  • Local refinement (free knots) [51],
  • Subdomain PoU decomposition (FBKAN) [131],
  • Selective basis enrichment (not uniform widening).
This seven-step procedure consolidates the full review into a single, consistent decision pipeline, with one basis-selection step and all relevant citations preserved.
Algorithm 1 Decision Algorithm for Choosing Between MLP and KAN and Configuring KANs
1:Input: task type, data scale, solution structure, constraints (speed, accuracy, physics)
2:Output: model class and recommended KAN configuration
3:Step 0: Identify task and constraints
4:Determine task: regression, PDE (PINN/PIKAN), operator learning, or classification
5:Assess data scale (few vs. many samples) and structural priors
6:Specify priorities: speed vs. accuracy, interpretability, physics constraints
7:Step 1: Choose MLP or KAN
8:if data are large and no structural prior is available then
9:  Use a standard MLP baseline
10:else
11:  Select a KAN architecture
12:end if
13:Step 2: Select KAN basis by solution characteristics
14:if solution is smooth or boundary-sensitive then
15:  Use B-spline KAN (cubic, extended grids)
16:else if solution is periodic or globally oscillatory then
17:  Use FourierKAN or KAF (RFF-based)
18:else if solution is polynomial-like or spectrally smooth then
19:  Use Chebyshev or Jacobi KAN with tanh normalization
20:else if solution has shocks, discontinuities, or boundary layers then
21:  Use SincKAN, rKAN, DKAN, or HRKAN
22:else if solution is multiscale or spatially heterogeneous then
23:  Use Wavelet KAN or FBKAN (PoU decomposition)
24:else if problem is high-dimensional or separable then
25:  Use KKAN, SPIKAN, or KAF (RFF)
26:end if
27:Step 3: Configure grid and basis resolution
28:Use coarse grids initially; refine adaptively (multilevel or free knots)
29:Apply residual-adaptive sampling if residuals localize
30:Tune RBF width, wavelet scale, or Sinc step sizes as needed
31:Step 4: Add physics-informed constraints (if applicable)
32:if physics constraints or PDE residuals are present then
33:  Construct PINN/PIKAN loss with data, PDE, and BC terms
34:  Apply residual reweighting, entropy viscosity, or augmented Lagrangian
35:end if
36:Step 5: Choose optimization strategy
37:Warm up with Adam/RAdam
38:Refine with (L-)BFGS or self-scaled quasi-Newton methods
39:Use variance-preserving initialization and moderate regularization
40:Step 6: Optimize for speed or accuracy
41:if speed is the primary objective then
42:  Prefer ReLU-KAN, HRKAN, FastKAN, LeanKAN
43:  Enable JAX/CUDA kernels and domain decomposition
44:else
45:  Prefer B-spline, Chebyshev/Jacobi, Wavelet KANs
46:  Use adaptive grids, physics losses, and staged optimization
47:end if
48:Step 7: Final refinement
49:Apply pruning and sparsity regularization
50:Increase capacity locally rather than uniformly
51:Validate convergence and generalization
11 Current Gaps and Path Forward
Despite rapid growth in publications and open-source implementations (Tables LABEL:KAN_vs_MLP, LABEL:githubs), the KAN ecosystem remains fragmented. Most studies show strong empirical potential in regression, operator learning, and PDE solving, yet comparisons are inconsistent and rarely grounded in theory. The field now requires principled foundations rather than isolated “KAN vs. MLP” or “basis-vs.-basis” experiments.
Limits of Current Comparative Frameworks
(1) KAN vs. MLP benchmarks.
These comparisons often match parameter counts but ignore essential differences: functional roles of parameters, basis-induced inductive bias, and normalization/grid constraints. Outcomes range from clear KAN advantages to near parity [41, 42, 43], offering little predictive insight into when KANs should win.
(2) Basis-vs.-basis races.
Comparisons of splines, Chebyshev, Gaussians, etc. [44] highlight problem dependence but mask deeper structure. Each basis involves choices of normalization, grids, and regularization—too many degrees of freedom for a single benchmark to be conclusive. For fair comparisons, future studies should standardize key variables such as grid resolution, normalization schemes, training settings, and evaluation metrics.
Overall, KANs are not a single architecture but a framework whose performance depends on principled design choices rooted in approximation theory, numerical conditioning, and optimization geometry.
Case Study: The Many Faces of a Chebyshev KAN:
A plain Chebyshev KAN can be unstable, but targeted refinements substantially improve behavior:
  • Domain stabilization: tanh normalization confines inputs to [1,1][-1,1] for numerically stable Chebyshev evaluation (Section 5.2).
  • Architectural hybrids: Linear heads reduce overfitting in inverse/PDE tasks [70].
  • Grid-informed adaptivity: Learnable tanh-based grids add local adaptivity [57].
  • Principled optimization: Hybrid Adam–L-BFGS schedules improve spectral conditioning [69, 68].
  • Adaptive polynomial degree: Making the polynomial degree itself trainable, as in Adaptive PolyKAN [295].
Similar refinements benefit other bases: adaptive centers for Gaussian KANs [59], free-knot updates for spline KANs [2, 51], and domain decomposition for finite-basis KANs [131].
A Multi-Pillar Framework for Future KAN Research
Pillar 1: Robust component library.
Canonical, stable implementations for Chebyshev, spline, Gaussian/RBF, Fourier, and wavelet bases with standardized normalization and initialization.
Pillar 2: Basis selection theory.
Principles linking problem structure to basis choice, including spectral alignment, NTK conditioning, and approximation–regularization trade-offs.
Pillar 3: Finite-width optimization theory.
Beyond NTK predictions [174, 69]; characterize finite-width loss geometry, adaptive basis movement, and optimization dynamics. Include Extreme Learning Machine–style regimes [293] as an alternative training strategy for KANs.
Pillar 4: Generalization and regularization.
Joint treatment of parameter count, basis complexity, smoothness, and explicit $\ell_{1}/\ell_{2}$ penalties for improved generalization bounds.
Pillar 5: Composition and hierarchy.
Depth in KANs currently has no formal interpretation. Open problems include understanding the functional role of layer composition, the interactions among heterogeneous bases, and the trade-off between expressivity and stability under deep compositions (see also the theoretical results of [58, 289] and the empirical observations of [283] [^55] https://github.com/geoelements-dev/mlp-kan).
Pillar 6: Interpretability and identifiability.
Uniqueness of learned univariate functions, robustness, and symbolic identifiability.
Pillar 7: RKHS-based formulation.
Develop RKHS-guided KAN theory using recent formulations [289] and earlier functional-analytic perspectives on KST [107].
Pillar 8: New basis functions.
Introduce and test new basis families in KANs, including Newton basis functions [292, 294], low-order Wendland functions, and Matérn/Sobolev kernels [122].
Outlook
A unified framework combining stable components, principled basis selection, finite-width theory, RKHS insights, and exploration of new bases can guide future KAN development.
12 Conclusion
This review has presented a systematic and unified perspective on Kolmogorov–Arnold Networks (KANs), bridging their theoretical foundations, architectural principles, and rapidly growing body of practical applications. By synthesizing insights from the literature together with our own analysis, we identify several key conclusions that characterize the current state of KAN research and point toward promising future directions.
Theoretical perspective.
Modern KANs are best understood as architectures inspired by, rather than direct realizations of, the Kolmogorov Superposition Theorem (KST). The KST provides a conceptual blueprint: its central insight—that any continuous multivariate function admits a representation as sums of univariate functions—motivates the characteristic activate–then–sum structure of KANs. However, contemporary KAN architectures depart fundamentally from the classical theorem by replacing non-smooth, universal inner functions with smooth, learnable basis functions such as splines or polynomials. This departure sacrifices a direct constructive link to KST, but enables efficient, stable, and data-driven approximation in practical settings.
Relationship to classical and modern methods.
By positioning KANs relative to kernel methods and multilayer perceptrons (MLPs), this review clarifies their functional role and architectural advantages. We showed that shallow, one-dimensional KANs are mathematically equivalent to classical kernel or basis-expansion methods, differing primarily in their training procedure. In contrast, higher-dimensional or deep KANs diverge structurally from kernel methods by constructing multivariate interactions through additive, compositional layers rather than explicit tensor-product or radial bases, thereby mitigating the curse of dimensionality.
The introduction of depth further distinguishes KANs from linear basis-expansion models. Hidden layers induce both a multiplicative growth in effective polynomial degree and a nonlinear coupling of the final function coefficients. This coefficient entanglement underlies the expressive power of deep KANs, but also explains the optimization challenges observed in practice, including non-convex loss landscapes, parameter redundancy, and sensitivity to initialization.
From the perspective of neural networks, KANs and MLPs are expressively equivalent, but differ markedly in how nonlinearity is deployed. KANs relocate nonlinearity from fixed node-wise activations to learnable edge-wise functions, replacing the MLP paradigm of sum–then–activate with an activate–then–sum architecture. This structural shift provides a strong inductive bias, improves parameter efficiency, and enhances interpretability through explicit functional components.
Practical design and applications.
A central lesson emerging from the expanding KAN ecosystem is that performance is not intrinsic to the “KAN” label itself, but instead arises from informed design choices. Among these, basis selection is the dominant design axis. B-splines naturally encode locality, Chebyshev polynomials offer spectral accuracy for smooth problems, Fourier bases are well suited for periodic structures, and wavelets enable multi-scale representations. No single basis is universally optimal; effective KAN design must align the chosen basis with the structure of the target function.
Beyond basis choice, successful KAN applications consistently employ a rich toolkit of accuracy and efficiency enhancements. These include physics-informed loss formulations, adaptive sampling and grid strategies, domain and function decomposition, and problem-specific regularization techniques. Such methods, often inherited from the broader scientific machine learning literature, are essential for achieving robust and competitive performance.
Outlook.
In summary, Kolmogorov–Arnold Networks represent a shift away from monolithic, fixed-activation models toward structured, modular, and interpretable architectures. Their strength lies not in a single rigid formulation, but in the flexibility to encode prior knowledge directly into the functional building blocks of the network. Future progress will depend on the development of a more unified theoretical framework that connects basis selection, optimization dynamics, conditioning, and generalization. Establishing such principles will be key to fully realizing the potential of KANs as a general-purpose tool for scientific computing and data-driven modeling.
13 Acknowledgments
This work was supported by the General Research Fund (GRF No. 12301824, 12300922) of Hong Kong Research Grant Council.